### v2

# Limit theorems for non-Hermitian ensembles

Abstract.
The distribution of the modulus of the extreme eigenvalues is investigated for the complex Ginibre and complex induced Ginibre ensembles in the limit of large dimensions of random matrices. The limiting distribution of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble, with a proportional rectangularity index, is the Gumbel distribution. The independence of these extrema is established, at appropriate scaling, for large matrices from the complex Ginibre ensemble as well as from the complex induced Ginibre ensemble for either fixed or proportional rectangularity indexes. The joint distribution of the eigenvalue moduli for the complex Ginibre ensemble is derived in this paper. It is equal to the joint distribution of independent random variables, each following a Gamma-Rayleigh distribution. In the limit of a large size of the complex Ginibre matrices, the left and right tail distributions of the minimum modulus are the Rayleigh and Weibull distributions, respectively. The limiting left tail distribution of minimum modulus is the same for these non-Hermitian ensembles when the rectangularity index of the complex induced Ginibre ensemble is equal to zero. This phenomenon is also verified for the right tail distribution of this minimum.
OLIVIA V. AUSTER
1. Introduction
The attractiveness of the Random Matrix Theory (RMT) lies in the possibility of using it as a means to model problems in high dimensions and perform related calculations analytically. This is mainly due to the invariance property of the probability distribution of certain matrix ensembles. As stated in [21], the statistical properties of the spectrum of random matrix ensembles are independent of the nature of the probability distribution that defines these ensembles in the limit of large sizes of these matrices. They only depend on the invariance of these distributions. Matrix ensembles that are characterised by invariant probability distribution are those from classical compact matrix groups studied in references [32] and [19]. The complex Ginibre ensemble is one of them and is a special case of the Ginibre-Girko ensemble with maximal non-Hermiticity as presented in [10] and [2].
The complex Ginibre ensemble was first defined as a mathematical concept to model phenomena from the physics of particles. Statistical properties of eigenvalues of matrices from this ensemble are studied to provide an understanding by analogy of the dynamics of nuclei. More precisely, as stated in [12], the distribution of eigenvalues of a complex Ginibre matrix is comparable to that of the distribution of the positions of charges of a two-dimensional Coulomb gas in a harmonic oscillator potential, at a specific temperature corresponding to the Dyson index $\beta=2$.
The complex Ginibre ensemble is the space of $N\times N$ complex matrices $J$ whose complex entries are the $J_{ij}=x_{ij}+\text{i}y_{ij}\in\mathbb{C}$ independent and identically distributed (i.i.d.) following a standard complex Gaussian distribution $\mathcal{N}_{\mathbb{C}}(0,1)$ with probability density
$P(J_{ij})=\frac{1}{\pi}e^{-|J_{ij}|^{2}}$
where $(x_{ij})_{1\leq i,j\leq N}\in\mathbb{R}^{N\times N}$ and $(y_{ij})_{1\leq i,j\leq N}\in\mathbb{R}^{N\times N}$. The real and imaginary parts of the entries denoted $x_{ij}$ and $y_{ij}$ are independent random variables, each following a real Gaussian distribution $\mathcal{N}_{\mathbb{R}}(0,\frac{1}{2})$. Like the real and the symplectic Ginibre ensembles, this ensemble was introduced by J. Ginibre [12]. The space of $N\times N$ complex matrices from the complex Ginibre ensemble is endowed with a probability measure here denoted $\mu(J)$, where $d\mu(J)=P(J)|D(J)|$ and $P(J)$ is the joint probability density function of the entries $J_{ij}$ defined as
$P(J)=\frac{1}{\pi^{N^{2}}}\exp\left(-\operatorname{Tr}\left(JJ^{*}\right)\right)$
and $|D(J)|=\frac{1}{2^{N^{2}}}\bigwedge_{k=1}^{N}\bigwedge_{j=1}^{N}|dJ_{kj}\wedge dJ_{kj}^{*}|$ is called the cartesian volume element as presented in [17].
The joint probability density function of the eigenvalues of complex Ginibre matrices is expressed in close form as
$P_{N}(z)=\frac{1}{\pi^{N}C_{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)|\Delta(z)|^{2}$
where $z=(z_{1},\dots,z_{N})\in\mathbb{C}^{N}$ is the $N$-tuple of eigenvalues of the $N\times N$ complex matrix $J$ and $\Delta(z)$ is the Vandermonde determinant $\Delta(z)=\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|$, with $|\Delta(z)|^{2}=\Delta(z)\overline{\Delta(z)}$. The term $C_{N}$ is the normalisation factor defined as
$C_{N}=\frac{1}{\pi^{N}}\int_{\mathbb{C}^{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}\prod_{k=1}^{N}d^{2}z_{k}=\prod_{k=1}^{N}\Gamma(k+1)$
A topic which has been widely investigated in RMT is the limiting distribution of the eigenvalues of random matrix ensembles as $N$ goes to infinity. The notion of universality in this field has found relevance in establishing analytical expressions of the limiting distribution of random matrix eigenvalues. It is a mathematical concept asserting that the limiting distribution of the eigenvalues should not depend on the particular distribution of the random matrix entries [25]. This has led to several important results among which is referenced the circular law conjecture for non-Hermitian random matrix ensembles (cf. [13], [25], [3]).
Studies of the limiting distribution of the largest and smallest eigenvalues for Gaussian ensembles are presented in [29, 30, 31]. The distributions of eigenvalues (and their spacings) have been investigated with the computation of gap probabilities with respect to radial ordering for non-Hermitian random matrices and their chiral counterparts [2]. The statistical properties of extreme eigenvalue moduli of matrices from non-Hermitian ensembles have also been studied in the literature such as the limiting distribution of the spectral radius for the complex Ginibre ensemble [22, 23] and the real Ginibre ensemble in [24].
Section 2 is devoted to the analysis of the distribution of the eigenvalue moduli for the complex Ginibre ensemble. The main results are derived for the spectral radius and the minimum modulus. The limiting left and right tail distributions of the minimum modulus is also studied for large size of matrices in Section 2.1. Pursuing the analysis in the scaling limit $\sqrt{N}$ for this random matrix ensemble as $N$ goes to infinity, the independence of the scaled minimum moduli with respect to the scaled spectral radius is established in Section 2.3. The eigenvalues of matrices from the complex Ginibre ensemble form a determinantal point process and the joint probability density of their respective radius has been investigated in the work of E. Kostlan [18] extended in [16]. This result is also discussed in the present paper in Section 2.2.
Employing the method introduced in [22] from which limit theorems are derived at the edge of the spectrum for the complex and symplectic Ginibre ensembles, results establishing the limiting distribution of the scaled spectral radius and minimum modulus for the complex induced Ginibre ensemble [7] are presented in Section 3. This random matrix ensemble might find relevance for questions raised in different fields of physics. It would also define an ideal modelling framework for the expansion of the research related to non-Hermitian matrices. The complex induced Ginibre ensemble is a special case of the Feinberg-Zee ensemble and a generalisation of the complex Ginibre ensemble. The joint probability density function of the entries of matrices $A$ from the Feinberg-Zee ensemble is
$P_{FZ}(A)\propto\exp\left(-\operatorname{Tr}V(A^{\dagger}A)\right)$
The induced Ginibre ensembles correspond to the Feinberg-Zee ensemble with potential $V(y)=-\frac{\beta}{2}(y-L\log(y))$ where $L$ is a non-negative parameter called the rectangularity index.
Let $G$ denote a $N\times N$ complex induced Ginibre matrix. The joint probability density function of its entries is
$P_{indG}^{(\beta)}(G)=\Gamma^{(\beta)}\left(\det G^{\dagger}G\right)^{\frac{\beta}{2}L}\exp\left(-\operatorname{Tr}G^{\dagger}G\right)\\$
where $\beta=2$ for the complex induced Ginibre ensemble and the determinant $\left(\det G^{\dagger}G\right)^{a}=\prod_{k=1}^{N}|\lambda_{i}|^{2a}$, $a\geq 0$. The factor $\Gamma^{(\beta)}$ is the normalisation constant
$\Gamma^{(\beta)}=\pi^{-\frac{\beta}{2}N^{2}}\left(\frac{\beta}{2}\right)^{\frac{1}{2}N(N-L)}\prod_{j=1}^{N}\frac{\Gamma(\frac{\beta}{2}j)}{\Gamma(\frac{\beta}{2}(j+L)}$
In the limit of large matrices dimensions, the eigenvalues are spread across an annulus in the complex plane which is distinguished from the phenomenon identified for complex Ginibre ensemble thus referenced as the circular law and presented in the work from V. L. Girko [13] and the paper from T. Tao, V. Vu and M. Krishnapur [25]. The joint probability density function of the eigenvalues of complex induced Ginibre matrices is
$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{\Gamma_{N}}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\\$
with normalisation constant
$\Gamma_{N}$ $=\int\cdots\int\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\prod_{j=1}^{N}d^{2}\lambda_{j}$
$=N!\pi^{N}\prod_{k=1}^{N}\Gamma(k+L)$
It is understood from the definition of the joint probability density function of the eigenvalues that there is a repulsion of the eigenvalues from the origin of the complex plane due to the term $\det(GG^{\dagger})^{L}=\prod_{j=1}^{N}|\lambda|^{2L}$ appearing in the formula. This phenomenon is explained from a mathematical perspective with the formulation of the mean eigenvalue density, i.e., $\langle\rho_{N}(\lambda)\rangle=R_{1}(\lambda)$ where $R_{1}(\lambda)$ is the one-point correlation function of the eigenvalues. More precisely taking its limit as $N$ goes to infinity and for a rectangularity index $L$, the distribution of the eigenvalues is uniform and supported by a ring centred at the origin of the complex plane with a outer circle of radius $r_{out}=\sqrt{L+N}$ and an inner circle of radius $r_{in}=\sqrt{L}$. For small rectangularity indexes $L$ (close to zero), the density of the eigenvalues is uniform over the disk of radius $\sqrt{N}$ with a mean eigenvalue density converging to the mean eigenvalue density of the complex Ginibre ensemble.
In the present paper, an investigation is performed for the complex induced Ginibre ensemble from the work of B. Rider [22]. In this reference, the author details an analytical method to determine the limiting distribution of the scaled spectral radius for matrices from the complex and symplectic Ginibre ensembles. A similar investigation is undertaken here for the scaled spectral radius and scaled minimum modulus of eigenvalues of matrices from the complex induced Ginibre ensemble. The analysis is conducted at the outer and inner edges of the ring (the eigenvalues support on the complex plane) defined with a rectangularity index assumed to be proportional to $N$, i.e., $L=\alpha N$, $\forall\alpha>0$. An exact fit between the empirical distribution and analytical formulation of these extreme moduli is acknowledged numerically. Additionally, the independence of the spectral radius and minimum modulus is studied in the limit as $N$ goes to infinity for these two non-Hermitian ensembles.
Results stating the right and left tail eigenvalues distribution functions asymptotics for the complex elliptic Ginibre ensemble in the limit of weak non-Hermiticity is presented in [4]. The limiting tail distribution functions asymptotics of the eigenvalues modulus is also of interest. The similitude of the minimum modulus limiting (left and right) tail distributions between the two investigated non-Hermitian ensembles is explored in the present paper. This is the purpose of Section 4. Conclusions are set out in Section 5.
2. Limit theorems for the complex Ginibre ensemble
Let $A$ denote a $N\times N$ complex Ginibre matrix whose symmetrised joint probability density function of its eigenvalues, $P_{N}(z_{1},\cdots,z_{N})$, is defined as in Section 1. The probability that the minimum of the eigenvalue moduli, here denoted $r_{min}^{(N)}(A)$, is greater than the radius $r\in\mathbb{R}^{+}$ is expressed as follows
$P\left(r_{min}^{(N)}(A)\geq r\right)=\int_{|z_{1}|\geq r}\cdots\int_{|z_{N}|\geq r}P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d^{2}z_{k}$
This probability corresponds to the probability that no eigenvalue lies inside the disk of radius $r$. It is a function of the radius $r$ and is initially defined analytically in reference from P. J. Forrester [8] with the following formula,
$P\left(r_{min}^{(N)}(A)\geq r\right)=\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ (1)
This result is presented in the reference from [2] whose authors applied the Gram’s formula [20] to get the probability, that no eigenvalue lies in the disk of radius $r$ centred at the origin, as a finite product of regularised upper incomplete Gamma functions. Equation (1) is derived with another approach in the present paper as follows. More precisely, the Andreief’s integration formula [9] is used to express analytically the gap probabilities $P(r_{min}^{(N)}(A)\geq a)$, $\forall a\in\mathbb{R}^{+*}$.
More precisely, from the definition of the eigenvalues joint probability density function defined in Section 1,
$P\left(r_{min}^{(N)}(A)\geq a\right)=\frac{1}{\Gamma_{N}}\int_{|z_{1}|>a}\cdots\int_{|z_{N}|>a}\det\left[z_{k}^{N-j}\right]_{j,k=1}^{N}\det\left[\bar{z}_{k}^{N-j}\right]_{j,k=1}^{N}\prod_{k=1}^{N}dm(z_{k})$
where $dm(z_{k})=e^{-|z_{k}|^{2}}d^{2}z_{k}$ and $\Gamma_{N}=N!\pi^{N}\prod_{k=0}^{N-1}k!$ is the normalisation constant.
Applying the Andreief’s integration formula,
$P\left(r_{min}^{(N)}(A)\geq a\right)=\frac{1}{\Gamma_{N}}N!\prod_{k=0}^{N-1}2\pi\int_{a}^{+\infty}e^{-r^{2}}r^{2k}rdr=\prod_{k=0}^{N-1}\frac{1}{\Gamma(k+1)}\int_{a^{2}}^{+\infty}e^{-t}t^{k}dt$
It is elementary to derive the probability density function $p_{r_{min}^{(N)}(A)}$ of the smallest modulus $r_{min}^{(N)}(A)$ from Equation (1). It is convenient to take the logarithm of the probability $P(r_{min}^{(N)}(A)\geq r)$ and find its first derivative with respect to $r$ to finally get the probability density function of the smallest modulus $r_{min}^{(N)}(A)$ as
$p_{r_{min}^{(N)}(A)}\left(r\right)$ $=2re^{-r^{2}}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]$ (2)
Using a similar approach as for the case of no scaling, the survival distribution function of scaled minimum of eigenvalue moduli $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ for $N\times N$ complex Ginibre matrices $A$ has same analytical expression as for $r_{min}^{(N)}(A)$ replacing $r$ with $\sqrt{N}r$. The corresponding probability density function is the probability density function of the minimum modulus $r_{min}^{(N)}(A)$ multiply by $\sqrt{N}$ at the point $\sqrt{N}r$, i.e.,
$p_{\frac{r_{min}^{(N)}(A)}{\sqrt{N}}}\left(r\right)=\sqrt{N}p_{r_{min}^{(N)}(A)}\left(\sqrt{N}r\right)$ (3)
The spectral radius of the real Ginibre matrices is investigated in reference [11]. It is experimentally stated that its value converges almost surely to the standard deviation of the i.i.d. entries of scaled matrices from this random matrix ensemble. The distribution of the spectral radius for a fixed size of complex Ginibre matrices is explored in the following.
Lemma 1.
$P\left(r_{max}^{(N)}(A)\leq r\right)=\prod_{k=0}^{N-1}\frac{\gamma(k+1,r^{2})}{\Gamma(k+1)}$ (4)
where $\Gamma(k)$ is the Gamma function and $\gamma(k,r)$ is the lower incomplete Gamma function.
Proof.
The cumulative distribution function of the spectral radius of eigenvalues, denoted $r_{max}^{(N)}(A)$, for the complex Ginibre ensemble is also easily retrieved with the use of the Andreief’s integration formula applied to the following multiple integrals equation
$P\left(r_{max}^{(N)}(A)\leq r\right)$ $=\int_{|z_{1}|\leq r}\cdots\int_{|z_{N}|\leq r}P_{N}(z_{1},\cdots,z_{N})\prod_{k=1}^{N}d^{2}z_{k}=\prod_{k=0}^{N-1}\frac{1}{\Gamma(k+1)}\int_{0}^{r^{2}}e^{-t}t^{k}dt$
Corollary 2.
The $N$-th gap probability, i.e., the probability that all the eigenvalues of a $N\times N$ complex Ginibre matrix lies in the disk of radius $r$, is the joint distribution of independent random variables $\gamma_{k}$ each following a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ [1], with $\alpha_{k}=k$ and $\delta_{k}=1$.
Proof.
$P(r_{1}^{(N)}(A)\leq r,r_{2}^{(N)}(A)\leq r,...,r_{N}^{(N)}(A)\leq r)=\prod_{k=1}^{N}P(\gamma_{k}<r)$ (5)
where each random variables $\gamma_{k}$ are independent and follows a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ , with $\alpha_{k}=k$ and $\delta_{k}=1$. The Gamma-Rayleigh distribution is derived in reference [1].
Corollary 3.
Let $\left((r_{i}^{(N)}(A))^{2}\right)_{i=1}^{N}$ denote the squared (unordered) moduli of the $N$ eigenvalues $\left(z_{i}^{(N)}(A)\right)_{i=1}^{N}$ of a $N\times N$ complex Ginibre matrix $A$. The distribution of the squared spectral radius $\left(r_{max}^{(N)}(A)\right)^{2}$ is the joint distribution of the elements of a finite sequence $\{\zeta_{k}\}_{k=1}^{N}$. The $\zeta_{k}$ are independent random variables. Each $\zeta_{k}$ follows a Gamma distribution with shape parameter $\alpha_{k}=k$ and scale parameter $\delta_{k}=1$, i.e., $\zeta_{k}\sim Gamma(k,1)$.
Proof.
A simple proof of this result is as follows. Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble. Let $\left(r_{i}^{(N)}(A)\right)_{1\leq i\leq N}$ denote the moduli of the eigenvalues of the matrix $A$. The joint distribution of the square moduli $\left((r_{i}^{(N)}(A))^{2}\right)_{i=1}^{N}$ is derived from Equation (4) and is expressed as
$P\left((r_{max}^{(N)}(A))^{2}\leq r\right)=\prod_{k=1}^{N}\frac{\gamma(k,r)}{\Gamma(k)}=\prod_{k=1}^{N}P(\zeta_{k}\leq r)\\$
where $P(\zeta_{k}\leq r)=\frac{\gamma(k,r)}{\Gamma(k)}$ is the cumulative distribution function of the random variable $\zeta_{k}$ following a Gamma distribution with shape parameter $\alpha_{k}=k$ and rate parameter $\delta_{k}=1$.
The corresponding probability density function of the largest modulus $r_{max}^{(N)}(A)$ is derived from Equation (4). Taking the logarithm of the cumulative distribution function $P(r_{max}^{(N)}(A)\leq r)$ and then the first derivative with respect to $r$ of $\log\left(P(r_{max}^{(N)}(A)\leq r)\right)$, the probability density funvtion of the spectral radius $r_{max}^{(N)}(A)$ is,
$p_{r_{max}^{(N)}(A)}(r)=2re^{-r^{2}}\prod_{k=0}^{N-1}\frac{\gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2j}}{\gamma(j+1,r^{2})}\right]$ (6)
The scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ is now considered in the following. The probability distribution function $P\left(\frac{r_{max}^{(N)}(A)}{\sqrt{N}}\leq r\right)$ of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ is also derived using the Andreief’s integration formula [9],
$P\left(\frac{r_{max}^{(N)}(A)}{\sqrt{N}}\leq r\right)=\prod_{k=0}^{N-1}P\left(\frac{1}{N}\sum_{j=1}^{N-k}Z_{j}\leq r^{2}\right)$ (7)
where the random variable $Z^{(k)}=\sum_{j=1}^{k+1}Z_{j}$ follows a Gamma distribution with shape parameter $k+1$ and rate parameter $1$. The random variables $Z_{j}$, with $j\in\{1,...,k+1\}$, are independent and identically distributed. Each of the $Z_{j}$ follows a standard exponential distribution.
This is a known result presented in B. Rider’s work [22] from which he established the nature of the limiting distribution of the scaled spectral radius for matrices from the complex Ginibre ensemble in the limit as $N$ goes to infinity. The distribution of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ for matrices from the complex Ginibre ensemble is the standard Gumbel distribution in the limit as $N$ goes to infinity. The limiting distribution of scaled $k$-th modulus has also been established for this random matrix ensemble as well as for the symplectic Ginibre ensemble in [23], in light of the framework presented in [22].
A similar method of derivation (as the one applied for the scaled minimum modulus) leads to the formulation of the probability density function of the scaled spectral radius for the complex Ginibre ensemble as
$p_{\frac{r_{max}^{(N)}(A)}{\sqrt{N}}}\left(r\right)=\sqrt{N}p_{r_{max}^{(N)}(A)}\left(\sqrt{N}r\right)$ (8)
2.1. Limiting left and right tail distributions of the minimum modulus as $N$ goes to infinity
Theorem 4.
Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble. The left tail distribution of minimum modulus $r_{min}^{(N)}(A)$ (i.e., the smallest values of $r_{min}^{(N)}(A)$) converges to the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$ as $N$ goes to infinity.
More precisely,
$\lim_{r\longrightarrow 0}\lim_{N\longrightarrow+\infty}P(r_{min}^{(N)}(A)<r)=1-e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (9)
Proof.
The statement of Theorem 4 is derived from the formulation of the gap probability $P(r_{min}^{(N)}(A)\geq r)$ which corresponds to the $N$-th partial product presented in Equation (1).
$\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ $=e^{-r^{2}}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=e^{-r^{2}}\prod_{k=1}^{N-1}e^{-r^{2}}e_{k}(r^{2})$
$=e^{-r^{2}}\prod_{k=1}^{N-1}\left[1-e^{-r^{2}}\sum_{t=n}^{+\infty}\frac{r^{2t}}{t!}\right]=e^{-r^{2}}H^{(N)}(r,0)$
where $e_{n}(x)=\sum_{k=0}^{n}\frac{x^{k}}{k!}$ defines the $n$-th partial sum of the exponential function and $H^{(N)}(r,0)$ is the conditional probability (cf. Appendix B) that given one eigenvalue lies at the origin of the complex plane all the others are found outside the disk centred at zero with radius $r$.
The limit
$\lim_{N\rightarrow+\infty}P(r_{min}^{(N)}(A)\geq r)=e^{-r^{2}}\lim_{N\rightarrow+\infty}H^{(N)}(r,0)$ (10)
This conditional probability $H^{(N)}(r,0)$ is a result presented in [15] and [17], where, in the limit as $N$ goes to infinity and $r$ is small (or $r$ close to zero)
$\lim_{N\rightarrow+\infty}H^{(N)}(r,0)=1-\left[\frac{r^{4}}{2}+\frac{r^{6}}{6}+\frac{r^{8}}{24}+O(r^{10})\right]=1-O\left(r^{4}\right)$ (11)
The limit presented in Equation (11) is derived as follows. For small $x$, the lower incomplete gamma function has asymptotic expansion
$\gamma(a,x)=x^{a}\sum_{n=0}^{+\infty}(-1)^{n}\frac{x^{n}}{n!(a+n)}$ (12)
Thus,
$\frac{\gamma(k+1,r^{2})}{\Gamma(k+1)}=\frac{r^{2(k+1)}}{(k+1)\Gamma(k+1)}+\frac{r^{2(k+1)}}{\Gamma(k+1)}\sum_{n=1}^{+\infty}(-1)^{n}\frac{r^{2n}}{n!(k+n+1)}$
Then, for small $r$, i.e., $0<r\ll 1$,
$\lim_{N\rightarrow+\infty}H^{(N)}(r,0)=\prod_{k=1}^{+\infty}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{\gamma(k+1,r^{2})}{\Gamma(k+1)}\right]$
$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{r^{2(k+1)}}{(k+1)\Gamma(k+1)}-\frac{r^{2(k+1)}}{\Gamma(k+1)}\sum_{n=1}^{+\infty}(-1)^{n}\frac{r^{2n}}{n!(k+n+1)}\right]$
$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{r^{2(k+1)}}{(k+1)\Gamma(k+1)}+O\left(r^{2(k+2)}\right)\right]$
Thus,
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$
$=\exp\left[\log\left(\prod_{k=1}^{+\infty}\left[1-\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right]\right)\right]=\exp\left[\sum_{k=1}^{+\infty}\log\left(1-\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)\right]$
$=\exp\left[\sum_{k=1}^{+\infty}\left[-\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]\right]=\exp\left[-\sum_{k=1}^{+\infty}\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]$
$=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]^{j}$
$=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{4(k+1)}\right)\right)\right]^{j}$
which implies for small $r$ (or $r$ close to zero),
$\lim_{N\rightarrow+\infty}H^{(N)}(0,r)$ $=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$
$=\left[\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{4(k+1)}\right)\right)\right]^{j}\right]$
$=1-\left[\frac{r^{4}}{2}+\frac{r^{6}}{6}+\frac{r^{8}}{24}+O(r^{10})\right]=1-O\left(r^{4}\right)$
This implies that
$\lim_{N\rightarrow+\infty}H^{(N)}(r,0)$ $=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{4(k+1)}\right)\right)\right]^{j}$
Consequently, as $r$ goes to zero,
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)<r\right)=1-e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$
This does correspond to the cumulative distribution of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
Remark 2.1.
Using successive integrations by parts
$\int_{r^{2}}^{+\infty}e^{-t}t^{k}dt$ $=e^{-r^{2}}\sum_{j=0}^{k}\frac{k!}{(k-j)!}r^{2(k-j)}+k!\int_{r^{2}}^{+\infty}e^{-t}dt$ (13)
Consequently,
$P(r_{min}^{(N)}(A)\geq r)$ $=\prod_{k=0}^{N-1}\left[\frac{1}{\Gamma(k+1)}e^{-r^{2}}\sum_{j=0}^{k}\frac{k!}{(k-j)!}r^{2(k-j)}+\int_{r^{2}}^{+\infty}e^{-t}dt\right]$ (14)
Theorem 5.
Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble. The right tail distribution of the minimum of moduli $r_{min}^{(N)}(A)$ converges to the Weibull distribution with shape parameter $\kappa=4$ and scale parameter $\lambda=k^{\frac{1}{k}}$ as $N$ goes to infinity, i.e., for large $r$,
$\lim_{N\longrightarrow+\infty}P\left(r_{min}^{(N)}(A)<r\right)=1-e^{-\frac{r^{4}}{4}\left(1+O\left(\frac{1}{r^{2}}\right)\right)},\text{}$ (15)
Proof.
This result is established from Equation $9$ of the reference [15].
Let $r=\sqrt{N}|\lambda|$ such that $|\lambda|\leq 1$, where $\lambda$ is a complex number lying in the unit disk centred at the origin of the complex plane.
$\lim_{N\rightarrow+\infty}P\left(\frac{r_{min}^{(N)}(A)}{\sqrt{N}}\geq|\lambda|\right)$ $=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)\geq\sqrt{N}|\lambda|\right)$
$=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,N|\lambda|^{2})}{\Gamma(k+1)}=\lim_{N\rightarrow+\infty}\left[e^{-N|\lambda|^{2}}H(\sqrt{N}|\lambda|,0)\right]$
$=\exp\left(-\frac{N^{2}|\lambda|^{4}}{4}\left(1+O\left(\frac{1}{N|\lambda|^{2}}\right)\right)\right)$
This result corresponds to the distribution of rare events, i.e., the distribution of extreme events corresponding to the scaled minimum of the moduli $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ as $N$ goes to infinity. This result refers to the large deviations theory. This limit at the logarithmic scale, as $N$ goes to infinity, is
$\lim_{N\longrightarrow+\infty}\left(-\log\left[P\left(\frac{r_{min}^{(N)}(A)}{\sqrt{N}}\geq|\lambda|\right)\right]\right)=\frac{N^{2}}{4}|\lambda|^{4}=c_{N}\inf_{\{z\geq|\lambda|\}}{I(z)}$ (16)
with speed $c_{N}=\frac{N^{2}}{4}$ and rate function $I(z)=|z|^{4}$.
2.2. The joint distribution of the eigenvalue moduli for the complex Ginibre ensemble
G. Dubach proved in [6] that the collection of squared moduli of the eigenvalues of $N\times N$ matrices from the complex Ginibre ensemble has the same distribution as the collection of independent random variables $\zeta_{k}$, each following a Gamma distribution with shape parameter $\alpha_{k}=k$ and scale parameter $\delta_{k}=1$, i.e., $\zeta_{k}\sim Gamma(k,1)$. He did provide proof of this result which involves the use of the Andreief’s integration formula [9] with radially symmetric functions defined on the complex plane.
Theorem 6.
The joint distribution of the unordered eigenvalue moduli $\{r_{k}^{(N)}(A)\}_{k=1}^{N}$ of $N\times N$ matrices $A$ from the complex Ginibre ensemble is the joint distribution of independent random variables $\zeta_{k}$ each following a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$, with $\alpha_{k}=k$ and $\delta_{k}=1$,
$\{r_{1}^{(N)}(A),...,r_{N}^{(N)}(A)\}\overset{d}{=}\{\zeta_{1},...,\zeta_{N}\}$
Proof.
Let $\left(\Omega,\mathcal{F},\mu\right)$ denote a measure space with $\Omega=\mathbb{C}$, $\mathcal{F}$ a $\sigma$-algebra on the complex plane and the measure $\mu$, a nonnegative function, defined on the measurable space $\left(\Omega,\mathcal{F}\right)$. Let $Q$ denote the set of $N$ moduli $r_{k}^{(N)}(A)$ of $N\times N$ matrices $A$ from the complex Ginibre ensemble, i.e., $Q=\left(q_{k}\right)_{k=1}^{N}=\left(r_{k}^{(N)}(A)\right)_{k=1}^{N}$. Let $\varphi$ denote a (suitable) test function with $\mu$ the reference measure on the complex plane. Using the Andreief’s integration formula
$\mathbb{E}\left[\prod_{q_{k}\in Q}\varphi(q_{k})\right]$ $=\int...\int\prod_{k=1}^{N}\varphi(|z_{k}|)P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d\mu(z_{k})$
$=\prod_{k=1}^{N}\int_{0}^{+\infty}\varphi(r)\frac{2r^{2k-1}e^{-r^{2}}}{\Gamma(k)}dr=\prod_{k=1}^{N}\int_{0}^{+\infty}\varphi(r)p_{GR(k,1)}(r)dr=\prod_{k=1}^{N}\mathbb{E}\left[\varphi(\zeta_{k})\right]$
where each $\zeta_{k}$ follows a Gamma-Rayleigh distribution with shape parameter $\alpha_{k}=k$ and scale parameter $\delta_{k}=1$. The Gamma-Rayleigh distribution is derived in the work of E. Akarawak, I. Adeleke and R. Okafor [1].
Remark 2.2.
[…]
$\left\{|z_{1}|,...,|z_{n}|\right\}\overset{d}{=}\left\{R_{(1)},...,R_{(n)}\right\}$
where $|z_{1}|\geq...\geq|z_{n}|$ are the order statistics of the moduli and where $R_{(1)}\geq...\geq R_{(n)}$ is the order statistics of independent random variables $R_{1},...,R_{n}$ with $R_{k}$ of density proportional to $t\longmapsto t^{2k-1}e^{-nV(t)}1_{[0,+\infty)}(t)$ with $V(t)=t^{2}$, for every $1\leq k\leq n$.
Remark 2.3.
The joint distribution of the moduli of $N\times N$ matrices for the complex Ginibre ensemble has been investigated in the work of E. Kostlan [18]. He stated that the collection of moduli of the eigenvalues of $N\times N$ matrices from the complex Ginibre ensemble has the same distribution as the collection of random variables $\{X_{2k}\}_{1\leq k\leq N}$ where, $\forall k\in\{1,...,N\}$, $X_{2k}$ follows a chi distribution $\chi_{2k}$. A slight extension of this result would correspond to state the next Theorem 7.
Theorem 7.
The collection of unordered eigenvalue moduli $|z_{k}|$ of $N\times N$ matrices from the complex Ginibre ensemble has the same distribution as the collection of random variables $\{\frac{X_{2k}}{\sqrt{2}}\}_{1\leq k\leq N}$ where, $\forall k\in\{1,...,N\}$, $X_{2k}$ follows a chi distribution $\chi_{2k}$.
Proof.
Let $Q$ denote the set of $N$ moduli $|z_{k}|$ of matrices from the complex Ginibre ensemble, i.e., $Q=\left(q_{k})\right)_{k=1}^{N}$. Let $f$ denote a (suitable) test function. Using the Andreief’s integration formula,
$\mathbb{E}\left[\prod_{q_{k}\in Q}f(q_{k})\right]$ $=\int...\int\prod_{k=1}^{N}f(|z_{k}|)P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d\mu(z_{k})$
$=\prod_{k=1}^{N}\int_{0}^{+\infty}f(v/\sqrt{2})\frac{v^{2k-1}e^{-v^{2}/2}}{2^{k-1}\Gamma(k)}dv=\prod_{k=1}^{N}\int_{0}^{+\infty}f(v/\sqrt{2})p_{\chi_{2k}}(v)dv=\prod_{k=1}^{N}\mathbb{E}\left[f\left(\frac{\alpha_{k}}{\sqrt{2}}\right)\right]$
Each random variable $\alpha_{k}$ follows a chi distribution with parameter $2k$, i.e., $\alpha_{k}\sim\chi_{2k}$ for $k\in\{1,...,N\}$.
2.3. Independence of the scaled spectral radius and the scaled minimum modulus for the complex Ginibre ensemble
Theorem 8.
Let A denote a $N\times N$ complex Ginibre matrix. The scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ are independent random variables.
Proof.
Applying the Andreief’s integration formula
$P\left(r_{N}\geq r\text{ and }R_{N}\leq R\right)$ $=\prod_{k=0}^{N-1}\frac{\gamma(k+1,NR^{2})}{\Gamma(k+1)}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]$
Furthermore, taking the limit as $N$ goes to infinity
$\lim_{N\rightarrow+\infty}P\left(r_{N}\geq r\text{ and }R_{N}\leq R\right)=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\gamma(k+1,NR^{2})}{\Gamma(k+1)}\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]$
Using the framework presented in [22], it is known that
$\lim_{N\rightarrow+\infty}P\left(R_{N}\leq R^{(N,x)}\right)=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\gamma(k+1,N\left(R^{(N,x)}\right)^{2})}{\Gamma(k+1)}=F_{X}(x)=e^{-e^{-x}}$
where $F_{X}(x)$ is the limiting cumulative distribution function of the scaled spectral radius $R_{N}$ defined in [22]. More precisely, it is the distribution function of the standard (maximum) Gumbel distribution. Here, $R=R^{(N,x)}=1+\sqrt{\frac{1}{2N}}\left(\log\frac{\sqrt{N/2\pi}}{\log N}+x\right)^{1/2}$ and $X$ is a random variable following a standard (maximum) Gumbel distribution. The specification of $R^{(N,x)}$ is slightly different from the result presented in [22] as a scaling $\sqrt{N}$ is used here $\left(\text{and not }2\sqrt{N}\right)$ which corresponds to the radius of the disk defined as the support of the spectrum of $N\times N$ random matrices from the complex Ginibre ensemble.
From the results presented in [2], for small $r$ of order $\frac{1}{\sqrt{N}}$ (i.e. where $Nr^{2}$ is kept fixed while $N$ goes to infinity), as $N$ goes infinity (or $N$ very large) also known as the origin limit
$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=E_{0}^{(2)}\left(\sqrt{N}r\right)$
Defined in [2], $E_{0}^{(\beta)}\left(\sqrt{N}r\right)$ is the probability that zero eigenvalue lies inside the disk of radius $\sqrt{N}r$ centred at the origin of the complex plane, and all eigenvalues lie outside. The index $\beta=2$ for the complex Ginibre ensemble. This implies that
$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}P\left(\frac{r_{min}^{(N)}(A)}{\sqrt{N}}\geq r\right)=\lim_{N\rightarrow+\infty}P\left(r_{N}\geq r\right)$
Consequently,
$\lim_{N\rightarrow+\infty}P\left(r_{N}\geq r\text{ and }R_{N}\leq R\right)$ $=\lim_{N\rightarrow+\infty}P\left(r_{N}\geq r\right)\lim_{N\rightarrow+\infty}P\left(R_{N}\leq R\right)$
The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ for the complex Ginibre ensemble are then independent random variables in the limit as $N$ goes to infinity.
3. Limit theorems for the complex induced Ginibre ensemble
The limiting distribution of the spectral radius of non-Hermitian ensembles as well as its precise localisation near the edge of the unit disk is investigated in the work of B. Rider [22]. He dedicated his studies to the complex and symplectic Ginibre ensembles introduced by J. Ginibre [12]. These statistical ensembles share similar features such as the universality conjecture known as the circular law. Their eigenvalues move towards the unit disk as the size of the matrices increases. The following results are derived for the complex induced Ginibre ensemble using the methodological approach exposed in Rider’s work. Let $G$ denote a $N\times N$ random matrix from the complex induced Ginibre ensemble with rectangularity index $L$ defined in [7]. At appropriate scalings, the scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{L+N}}$ evolves near the outer radius $r_{out}$ while the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{L}}$ is a random variable fluctuating around the inner radius $r_{in}$. The outer and inner radii are equal to $\sqrt{L+N}$ and $\sqrt{L}$, respectively.
3.1. Limit theorems of the outer and inner edges of the ring
The scaled spectral radius and scaled minimum modulus of eigenvalues for $N\times N$ matrices from the complex induced Ginibre ensemble, with proportional rectangularity index, are Gumbel distributed in the limit as $N$ goes to infinity. More precisely, the scaled spectral radius, for matrices from the complex induced Ginibre ensemble with proportional rectangularity index, i.e., $L=\alpha N$ with $\alpha>0$, follows a maximum Gumbel distribution at the edge of the outer circle of the ring as $N$ goes to infinity.
At the outer edge of the ring, setting $a=1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}$ where $f_{N}(x)$ is an increasing function in both $x$ and $N$, it is found that
$\mathcal{P}_{N}\left(a\right)=P\left(\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}\leq a\right)=\prod_{k=1}^{N}P\left(\frac{1}{(1+\alpha)N}\sum_{j=1}^{N+L-k+1}Z_{j}\leq a^{2}\right)$ (17)
where the $Z_{j}$ for $j=\{1,\cdots,N+L-k+1\}$ are independent and identically distributed random variables following a standard exponential distribution.
Theorem 9.
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}$ denote the scaled spectral radius. Then,
$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[R_{N}\leq 1+\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}+\frac{x}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}\right]=e^{-e^{-x}}$ (18)
where $\gamma_{\alpha,N}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$.
This limiting probability distribution is the standard Gumbel distribution for maxima. The corresponding cumulative density function is the function $F_{X}^{(\text{Gumbel max})}(x)=\exp(-\exp(-x))$.
The scaled spectral radius $R_{N}$ is approximated as
$R_{N}\approx 1+T_{\alpha,N}+\xi_{\alpha,N}$ (19)
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ with $X$ following a standard Gumbel (maximum) distribution.
This result is similar to the one detailed in [22] for the complex Ginibre ensemble. The exact formula of the probability density function $p_{R_{N}}\left(r\right)$ of the scaled spectral radius $R_{N}$, for a proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$, is derived from equation (17) and is defined as follows for any value of NN
$p_{R_{N}}\left(r\right)=2[(1+\alpha)N]^{\alpha N}r^{2(\alpha N-\frac{1}{2})}e^{-(1+\alpha)Nr^{2}}\prod_{k=1}^{N}\frac{\gamma\left(k+\alpha N,(1+\alpha)Nr^{2}\right)}{\Gamma(k+\alpha N)}\sum_{j=1}^{N}\left[\frac{[(1+\alpha)N]^{j}r^{2j}}{\gamma\left(j+\alpha N,(1+\alpha)Nr^{2}\right)}\right]$ (20)
with $\gamma(.,.)$ is the lower incomplete Gamma function.
The proof of Theorem 9 is presented in Appendix A.1. The analytical formulation of the probability density function $p_{R_{N}}$ of the scaled spectral radius $R_{N}$ does fit exactly its empirical distribution created from the generation of 1010 000000 complex induced Ginibre matrices with proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$. This is presented in Figure 1. As the size of these matrices increases to large numbers, the distribution of $R_{N}$ narrowing close to the outer radius $r_{out}=1$ (Figure 2).
(a) $\alpha=\frac{1}{9}$
(b) $\alpha=\frac{1}{90}(red),\alpha=\frac{1}{9}(blue)\text{ and }\alpha=\frac{4}{9}(black)$
Figure 1. (left panel) Empirical probability distribution (histogram) of the scaled spectral radius $R_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=90$, $\alpha=\frac{1}{9}$ and a proportional rectangularity index $L=\alpha N$. The corresponding exact (analytical) probability density function $p_{R_{N}}$ (solid curve). (right panel) The analytical (exact) probability density function of scaled spectral radii $R_{N}$ for $K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=90$. The rectangularity index $L$ is proportional to $N$ such that $L=\alpha N$ with $\alpha>0$. The results are presented for different values of $\alpha=\{\frac{1}{90},\frac{1}{9},\frac{4}{9}\}$. Graphs generated with MATLAB.
Copyright Olivia V. Auster for code and graphs.
Figure 2. Empirical probability distribution (histogram) of the scaled spectral radius $R_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=90$, $\alpha=\frac{1}{9}$ and a proportional rectangularity index $L=\alpha N$. The exact (analytical) probability density function $p_{R_{N}}$ (blue curve). Limiting probability distributions of the scaled spectral radius $R_{N}$ presented with the red and black curves for large $N=10e3$ and $N=2\times 10e3$, respectively. Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
The exact formula of the cumulative density function of $R_{N}$ converges towards the asymptotic distribution as $N$ goes to infinity (Figure 3).
Figure 3. (left panel) The exact cdf of the scaled spectral radius $R_{N}$ for $N=10e3$ (blue), $N=10e4$ (blue dotted) and $N=10e7$ (black) and $\alpha=1$. The asymptotic cdf of $R_{N}$ with asymptotic location and scale parameters $\mu$ and $\sigma$, respectively (red dotted curve). (right panel) The exact formulation of the scaled spectral radius $R_{N}$ cumulative distribution (black curve) for $N=10e7$. The asymptotic cumulative distribution (red curve) parametrised with asymptotic location $\mu$ and scale $\sigma$ parameters. The parameter $\alpha=1$ for the two curves. Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
The scaled minimum modulus of eigenvalues, for matrices from the complex induced Ginibre ensemble with a proportional rectangularity index, follows a Gumbel (minimum) distribution at the edge of the inner circle of the ring. At the inner edge of the ring, setting $a=1-\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ is an increasing function in both $x$ and $N$, it is found that
$\mathcal{P}_{N}\left(a\right)=P\left(\frac{r_{min}^{(N)}(A)}{\sqrt{\alpha N}}\geq a\right)=\prod_{k=1}^{N}P\left(\frac{1}{\alpha N}\sum_{j=1}^{k+L}Z_{j}\geq a^{2}\right)$ (21)
where the $Z_{j}$ for $j=\{1,\cdots,k+L\}$ are independent and identically distributed random variables following a standard exponential distribution.
Theorem 10.
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $r_{N}$ denote the scaled minimum modulus such that $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$.
$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]=e^{-e^{x}}$ (22)
in which $\gamma_{\alpha,N}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$.
This limit corresponds to the survival distribution function of the standard Gumbel distribution for minima. The scaled minimum of moduli $r_{N}$ is approximated with the following
$r_{N}\approx 1-T_{\alpha,N}+\xi_{\alpha,N}$ (23)
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ where $X$ follows a standard Gumbel (minimum) distribution.
The exact formula of the probability density function $p_{r_{N}}\left(r\right)$ of the scaled minimum modulus $r_{N}$ for proportional rectangularity index $L=\alpha N$ is derived from equation (21) and is defined as follows for any value of NN
$p_{r_{N}}\left(r\right)=2(\alpha N)^{\alpha N}r^{2(\alpha N-\frac{1}{2})}e^{-\alpha Nr^{2}}\prod_{k=1}^{N}\frac{\Gamma(k+\alpha N,\alpha Nr^{2})}{\Gamma(k+\alpha N)}\sum_{j=1}^{N}\left[\frac{(\alpha N)^{j}r^{2j}}{\Gamma(j+\alpha N,\alpha Nr^{2})}\right]$ (24)
with $\Gamma(.,.)$ the upper incomplete Gamma function.
A detailed proof of Theorem 10 is provided in Appendix A.2. Numerical results are acknowledged in Figures 4 and 5. The analytical formulation of the probability density function of the scaled minimum modulus $r_{N}$ fits exactly its empirical distribution from the generation of 1010 000000 complex induced Ginibre matrices (left panel) with proportional rectangularity index $L=\alpha N$, $\alpha>0$. As the size of these matrices increases to large numbers (e.g., $N=10e3$), the distribution of $r_{N}$ narrows around the inner radius $r_{in}=1$.
(a) $\alpha=\frac{1}{10}$
(b) $\alpha=\frac{1}{100}(red)\text{, }\alpha=\frac{1}{10}(blue)\text{ and }\alpha=\frac{4}{10}(black)$
Figure 4. (left panel) The Empirical probability distribution (histogram) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=100$, $\alpha=\frac{1}{10}$ and a proportional rectangularity index $L=\alpha N,\alpha>0$. The exact (analytical) probability density function is presented with the solid curve. (right panel) The analytical (exact) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for
$K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=100$. The rectangularity index $L$ is proportional to $N$ such that $L=\alpha N$ with $\alpha>0$. The results are presented for different values of $\alpha=\{\frac{1}{100},\frac{1}{10},\frac{4}{10}\}$. Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
Figure 5. Empirical probability distribution (histogram) and the exact (analytical) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=100$, $\alpha=\frac{1}{10}$ and a proportional rectangularity index $L=\alpha N$. Limiting probability distributions of the scaled minimum modulus $r_{N}$ presented with the red and black curves for large $N=10e4$ and $N=2\times 10e4$, respectively.
Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
The exact formula of the cumulative density function of $r_{N}$ converges towards the asymptotic distribution defined with the location parameter $\mu$ and the scale parameter $\sigma$ of the Gumbel distribution derived from Equation (23), as $N$ goes to infinity (Figure 6).
Figure 6. (left panel) The exact cdf of the scaled minimum modulus $r_{N}$ for $N=10e5$ (blue), $N=10e6$ (blue dotted) and $N=10e7$ (black) and $\alpha=1$. The asymptotic cdf of $r_{N}$ with asymptotic location and scale parameters (red dotted curve). (right panel) The exact formulation of the scaled minimum modulus $r_{N}$ cumulative distribution (black curve) for $N=10e7$ and the asymptotic cumulative distribution (red curve) parametrised with asymptotic location and scale parameters. The parameter $\alpha=1$ for each curve. Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
3.2. Independence of the spectral radius and the minimum of the eigenvalue moduli for the complex induced Ginibre ensemble
In this section, the independence of the spectral radius $r_{max}^{(N)}(G)$ and the minimum of moduli $r_{min}^{(N)}(G)$ is demonstrated as $N$ goes to infinity for the complex induced Ginibre ensemble for fixed and proportional rectangularity indexes.
Theorem 11.
Let $\mathcal{B}_{L+N}$ and $\mathcal{B}_{L}$ denote Borel sets in the neighbourhood of $\sqrt{L+N}$ and $\sqrt{L}$, respectively. The events $\{r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}\}$ and $\{r_{min}^{(N)}(G)\in\mathcal{B}_{L}\}$ are independent in the limit as $N$ goes to infinity for fixed rectangularity index $L$.
Proof.
Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of a $N\times N$ complex induced Ginibre matrix $G$, respectively. The result, presented in Theorem 11, is derived with the formulation of the probability,
$P\left(r_{min}^{(N)}(G)\geq r\text{ and }r_{max}^{(N)}(G)\leq R\right)=\int_{r\leq|\lambda_{1}|\leq R}\cdots\int_{r\leq|\lambda_{N}|\leq R}P_{N}(\lambda_{1},\cdots,\lambda_{N})\prod_{j=1}^{N}d^{2}\lambda_{j}$
where $r$ and $R$ are in $\mathcal{B}_{L}$ and $\mathcal{B}_{L+N}$, respectively. The joint probability density function of the eigenvalues $\left(\lambda_{j}\right)_{j=1}^{N}$ for the complex induced Ginibre ensemble is
$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{N!\pi^{N}}\prod_{k=1}^{N}\frac{1}{\Gamma(k+L)}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)$
Let $r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}$ and $r_{min}^{(N)}(G)\in\mathcal{B}_{L}$ where $\mathcal{B}_{L+N}$ is a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$.
Applying the Andreief’s integration formula, the probability
$P\left(r_{min}^{(N)}(G)\geq r\text{ and }r_{max}^{(N)}(G)\leq R\right)=P\left(r_{max}^{(N)}(G)\leq R\right)\prod_{k=1}^{N}\left[1-\frac{\gamma(k+L,r^{2})}{\gamma(k+L,R^{2})}\right]$
For any $R$ in the Borel set $\mathcal{B}_{L+N}$, $\lim_{N\longrightarrow+\infty}\gamma(k+L,R^{2})=\Gamma(k+L)$ for fixed rectangularity index $L$.
Thus,
$\lim_{N\longrightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+L,r^{2})}{\gamma(k+L,R^{2})}\right]=\prod_{k=1}^{\infty}\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}$
and
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\text{ and }r_{max}^{(N)}(G)\leq R\right)$ $=P\left(r_{max}^{(\infty)}(G)\leq R\right)P\left(r_{min}^{(\infty)}(G)\geq r\right)$
Theorem 12.
Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of a $N\times N$ matrices $G$ from the complex induced Ginibre ensemble, respectively. Let $\mathcal{B}_{L+N}$ denote a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$, for proportional rectangularity index $L$. The events $\{r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}\}$ and $\{r_{min}^{(N)}(G)\in\mathcal{B}_{L}\}$ are independent, where $L=\alpha N$ with $\alpha>1$, in the limit as $N$ goes to infinity.
Proof.
Setting the rectangularity index as proportional to $N$, i.e., $L=\alpha N$, $\alpha>1$, let $\mathcal{B}_{L+N}$ denote a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$. Considering $R$ in $\mathcal{B}_{L+N}$ and $r$ in $\mathcal{B}_{L}$,
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\text{ and }r_{max}^{(N)}(G)\leq R\right)$ $=P\left(r_{max}^{(\infty)}(G)\leq R\right)\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$
The asymptotics of the regularised lower incomplete Gamma functions is presented in [26]. It is, uniformly for $x\geq 0$, and $a\rightarrow+\infty$ and/or $x\rightarrow+\infty$,
$P(a,x)=\frac{\gamma(a,x)}{\Gamma(a)}=\frac{1}{\Gamma(a)}\int_{0}^{x}t^{a-1}e^{-t}dt=\frac{1}{2}\operatorname{erfc}\left(-\eta(a/2)^{1/2}\right)-\mathcal{R}_{a}(\eta)$
where $\eta=\left[2(\mu-\log(1+\mu))\right]^{1/2}$ with $\mu=\lambda-1$ and $\lambda=\frac{x}{a}$
The remainder $\mathcal{R}_{a}(\eta)$ is represented as
$\mathcal{R}_{a}(\eta)\sim(2\pi a)^{-1/2}e^{-\frac{1}{2}a\eta^{2}}\sum_{k=0}^{+\infty}c_{k}(\eta)^{-k}$
with the coefficients $c_{k}$ defined as in [26] and [27]. These results are (uniformly) valid for $\eta\in\mathbb{R}$.
Remark 3.1.
$\gamma(k+\alpha N,R^{2})=\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)$
where
$\varepsilon_{k,N}(R)=\Gamma(k+\alpha N,R^{2})$
and
$\lim_{N\longrightarrow+\infty}\varepsilon_{k,N}(R)=0$
Then,
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$
$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)}\right]=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}\right)\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$
where $E_{k,N}(R,r)=\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}(R)}{\Gamma(k+\alpha N)^{2}}+O\left(\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}^{2}(R)}{(\Gamma(k+\alpha N))^{3}}\right)$ where $\lim_{N\longrightarrow+\infty}\varepsilon_{k,N}(R)=0$.
Furthermore, using the results in [26] and [27] where $\eta=\left[2(\mu-\log(1+\mu))\right]^{1/2}$ with $\mu=\frac{r^{2}}{k+\alpha N}-1$
$\lim_{N\rightarrow+\infty}\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}$ $=\lim_{N\rightarrow+\infty}\left[\frac{1}{2}\operatorname{erfc}\left(-\eta((k+\alpha N)/2)^{1/2}\right)-\mathcal{R}_{t}(\eta)\right]$
$=1-\lim_{N\rightarrow+\infty}\mathcal{R}_{t}(\eta)$
where $t=k+\alpha N$.
The limit of the partial product $\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ is established from the limit of its lower and upper bounds.
Lower bound. The limit
$\lim_{N\rightarrow+\infty}1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}=\lim_{N\rightarrow+\infty}\mathcal{R}_{t}(\eta)$
And for an appropriate choice of $r$, i.e., $r=O\left(\sqrt{\alpha N}\right)$, the parameter $\eta$ is not infinite. This implies that the remainder $\mathcal{R}_{t}(\eta)$ is finite, i.e., $\frac{1}{\mathcal{R}_{t}(\eta)}>C$ where $C$ is a positive constant. This implies that, for $N$ large enough and finite,
$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $>\prod_{k=1}^{N}\left(1+E_{k,N}(R,r)C\right)>\prod_{k=1}^{N}\left(1+\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}(R)C}{N\Gamma(k+\alpha N)^{2}}\right)=1+O\left(\varepsilon_{k,N}(R)\right)$
Upper bound. Using the Stirling formula,
$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $<1+O\left(\frac{\varepsilon_{k,N}(R)e^{-\alpha N}}{\left(N\right)^{\alpha N-\frac{1}{2}-1}}\right)$
where $\alpha>1$.
Applying the squeeze theorem,
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)=1$
Consequently,
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$ $=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}\right]$
$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}$
$=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)=P\left(r_{min}^{(\infty)}(G)\geq r\right)$
Finally,
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\text{ and }r_{max}^{(N)}(G)\leq R\right)$ $=P\left(r_{min}^{(\infty)}(G)\geq r\right)P\left(r_{max}^{(\infty)}(G)\leq R\right)$
Theorem 13.
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with proportional rectangularity index $L=\alpha N$, $\alpha>0$. The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ are independent random variables, under scaling $\sqrt{\alpha N}$, as $N$ goes to infinity. More precisely, let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$. Setting $\rho=\sqrt{\frac{1+\alpha}{\alpha}}$,
$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\xi_{\alpha}^{(N)}(y)\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\eta_{\alpha}^{(N)}(x)\right]=e^{-e^{y}}e^{-e^{-x}}$
in which $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$ and where
$\xi_{\alpha}^{(N)}(y)=\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }\eta_{\alpha}^{(N)}(x)=\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
The scaled spectral radius $R_{N}$ and the scaled minimum of moduli $r_{N}$ are approximated as
$R_{N}$ $\simeq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }r_{N}\simeq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{Y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
where $X$ and $Y$ are standard Gumbel (maximum) and standard Gumbel (minimum) distributed random variables, respectively.
The proof of Theorem 13 is detailed in Appendix A.3.
4. Limiting distribution of the minimum modulus of eigenvalues for matrices from the complex Ginibre and complex induced ensembles. A comparison
The limiting probability distributions of the minimum modulus of eigenvalues for matrices from the complex Ginibre and complex induced Ginibre ensembles are investigated in this section. Derived in Section 2, the survival probability distribution function of the minimum modulus $r_{min}^{(N)}(A)$ of a $N\times N$ complex Ginibre matrix $A$, is used to determine the corresponding analytical probability density function, as $N$ goes infinity. As derived in Section 2,
$p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]$
It represents the exact formulation of the probability density function of the $r_{min}^{(N)}(A)$ as the product of a partial product of the regularised upper incomplete Gamma functions and a partial sum.
Theorem 14.
Let $A$ denote a $N\times N$ random matrix from the complex Ginibre ensemble. The limiting left tail probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is the probability density function of the Rayleigh distribution with parameter $\frac{1}{\sqrt{2}}$, in the neighbourhood of zero, as $N$ goes to infinity. More precisely,
$\lim_{r\longrightarrow 0}\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (25)
Proof.
The limit of partial product $\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$, as $N$ goes to infinity, has been investigated in Section 2 for the complex Ginibre ensemble, and for small $r$ in the neighbourhood of zero, it is expressed as
$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$
The upper incomplete Gamma function $\Gamma(k,x)$, for small $x$ and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
The limit of the partial sum is derived as follows. Knowing that, for $r$ in the neighbourhood of zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$,
$\lim_{N\rightarrow+\infty}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$ $=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{r^{2j}}{\Gamma(j+1,r^{2})}=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{(r^{2})^{j}}{j!}=2e^{-r^{2}}re^{r^{2}}=2r$
Thus, the limiting left tail probability density function of the minimum of moduli for matrices from the complex Ginibre ensemble is
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)$ $=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$
$=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$
This function corresponds to the probability density function of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
This result is in line with the result acknowledged with Theorem 4.
The probability density function of the minimum modulus of eigenvalues for $N\times N$ matrices $G$ from the complex induced Ginibre ensemble is presented in Section 3. Its limit is established here for the particular case where the fixed rectangularity index $L=1$.
Theorem 15.
The minimum modulus of a $N\times N$ matrix $G$ from the complex induced Ginibre ensemble, with rectangularity index $L=1$, has a left tail probability density function of the Generalised Gamma distribution with scale parameter $\alpha=1$ and other parameters $\kappa=2$ and $\gamma=4$. The corresponding limiting probability density function, for $r$ in the neighbourhood of zero, as $N$ goes to infinity, is
$\lim_{r\longrightarrow 0}\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2r^{3}e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (26)
Proof.
Let $G$ denote a $N\times N$ complex induced Ginibre matrix. The limiting probability density function of the minimum modulus of eigenvalues, denoted $r_{min}^{(N)}(G)$, as $r$ approaches zero and as $N$ goes to infinity, is
$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2re^{-r^{2}}\lim_{N\rightarrow+\infty}\left(\prod_{k=1}^{N}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=1}^{N}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]\right)$
Furthermore, as $N$ goes to infinity and $r$ approaches zero,
$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=1-O\left(r^{4}\right)$
This is established with the results presented in Section 2. The partial product is extended from N1N-1 to $N$ which does not change the limit.
Also, the upper incomplete Gamma function $\Gamma(k,x)$, for $x$ approaches zero and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
For $r$ close to zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$. This implies that the limit of the partial sum, as $r$ approaches zero and as $N$ goes to infinity, is
$\lim_{N\rightarrow+\infty}2re^{-r^{2}}\sum_{j=1}^{N}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]\\ =2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{\Gamma(j+1,r^{2})}\times\left(1-O\left(r^{4}\right)\right)=2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{j!}\times\left(1-O\left(r^{4}\right)\right)$
Finally,
$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)$ $=2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{j!}\times\left(1-O\left(r^{4}\right)\right)$
$=2r^{3}e^{-r^{2}}\times\left(1-O\left(r^{4}\right)\right)=\frac{\frac{\kappa}{\alpha^{\gamma}}}{\Gamma(\gamma/\kappa)}r^{\gamma-1}e^{-(r/\alpha)^{\kappa}}\times\left(1-O\left(r^{4}\right)\right)$
This corresponds to the probability density function of the Generalised Gamma distribution with scale parameter $\alpha=1$ and other parameters $\kappa=2$ and $\gamma=4$, where $\Gamma(.)$ is the Gamma function.
Now, considering any rectangularity index $L\geq 0$, the limiting left tail distribution of the minimum of eigenvalues moduli for $N\times N$ random matrices $G$ form the complex induced Ginibre ensemble, as $N$ goes to infinity, is presented as follows.
Theorem 16.
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L\geq 0$. The left tail distribution of minimum of the moduli $r_{min}^{(N)}(G)$ is the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$ in the limit as $N$ goes to infinity and $r$ approaches zero zero,
$\lim_{N\longrightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-\exp\left[-\frac{r^{2(L+1)}}{(L+1)!}\right]+O\left(r^{4(L+1)}\right)$ (27)
Proof.
This result is derived using the same methodological approach presented in section 2.1. For positive values of $r$ in the neighbourhood of zero,
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)$ $=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+L)}}{(k+L)!}+O\left(r^{4(k+L)}\right)\right)\right]^{j}$
Consequently, in the limit as r goes to zero
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-\exp\left[-\frac{r^{2(L+1)}}{(L+1)!}\right]+O\left(r^{4(L+1)}\right)$
This corresponds to the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$. ∎
Corollary 17.
Setting the rectangularity index $L$ to zero, the limiting distribution of the minimum modulus $r_{min}^{(N)}(G)$ for the complex induced Ginibre ensemble corresponds to the limiting distribution of the minimum modulus $r_{min}^{(N)}(A)$ for the complex Ginibre ensemble, i.e., for small $r$
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-r^{2}}+O(r^{4})=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)<r\right)$
Corollary 18.
For positive values of $r$ in the neighbourhood of zero, as $N$ goes to infinity, and for a rectangularity index $L=0$,
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)+O\left(r^{5}\right)$
with shape parameter $\kappa=2$ and scale parameter $\lambda=1$. This does correspond to the probability density function of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
Proof.
$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\frac{\kappa}{\lambda}\left(\frac{r}{\lambda}\right)^{\kappa-1}e^{-(r/\lambda)^{\kappa}}+O\left(r^{3\kappa+1}\right)=2re^{-r^{2}}+O\left(r^{7}\right)$
where $\kappa=2$ and scale parameter $\lambda=1$. ∎
The limiting distribution of the minimum modulus for the complex induced Ginibre ensemble, in the limit as $N$ goes to infinity for fixed rectangularity index $L\in\left[0,\varepsilon\right]$ (with $\varepsilon\ll 1$), is derived using the Euler-Maclaurin summation formula and the asymptotic of the upper incomplete Gamma function as $r$ goes to infinity [28].
Thus,
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}=\exp{\left[\sum_{k=1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)\right]}$
Remark 4.1.
It is notified in [27], that
$\lim_{k\longrightarrow+\infty}\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}=\frac{1}{2}\operatorname{erfc}\left(\eta\left(k+L\right)^{1/2}\right)+O\left((k+L)^{-1/2}e^{-\frac{1}{2}(k+L)\eta^{2}}\right)$ (28)
where $\operatorname{erfc}$ is the complementary error function, $\lambda=\frac{r^{2}}{k+L}$, $\mu=\lambda-1$ and $\eta=\left(2\left[\mu-\log\left(1+\mu\right)\right]\right)^{1/2}$.
From Remark 4.1,
$\lim_{k\longrightarrow+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)=0$
Now, the asymptotic expansion of the upper incomplete Gamma function (cf. [28]) as $z$ goes to infinity, is
$\lim_{z\longrightarrow+\infty}\Gamma(a,z)=z^{a-1}e^{-z}+O(z^{a-2}e^{-z})$
Thus, for large $r$,
$\frac{1}{2}\log\left(\frac{\Gamma(L+1,r^{2})}{\Gamma(L+1)}\right)=\frac{1}{2}\left[-r^{2}+\log\left(\frac{r^{2L}\left(1+O\left(\frac{1}{r^{2}}\right)\right)}{\Gamma(L+1)}\right)\right]$ (29)
A precise evaluation of the integral $\int_{1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)dk$ is $\frac{r^{4}}{4}+O(r^{2})$.
With the use of the Euler-Maclaurin summation formula, the infinite series of $\log$ is
$\sum_{k=1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)=\int_{1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)dk+\frac{1}{2}\log\left(\frac{\Gamma(L+1,r^{2})}{\Gamma(L+1)}\right)+O\left(r\right)$
Consequently,
$\sum_{k=1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)=-\frac{r^{4}}{4}+\frac{1}{2}\left[-r^{2}+\log\left(\frac{r^{2L}\left(1+O\left(\frac{1}{r^{2}}\right)\right)}{\Gamma(L+1)}\right)\right]+O\left(r^{2}\right)$
This implies that for large $r$,
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-\frac{r^{4}}{4}}\exp\left(-\frac{r^{2}}{2}+\frac{1}{2}\log\left(\frac{r^{2L}+r^{2(L-1)}O\left(1\right)}{\Gamma(L+1)}\right)+O\left(r^{2}\right)\right)$ (30)
For a rectangularity index $L=0$, for large $r$, and $A$ a $N\times N$ complex Ginibre matrix,
$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-\frac{r^{4}}{4}\left(1+O\left(\frac{1}{r^{2}}\right)\right)}=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)<r\right)$ (31)
5. Conclusions
The distribution of the minimum modulus of matrices from the complex Ginibre ensemble is analytically derived with the use of the Andreief’s integration formula. It does correspond to a $N$-th partial product of regularised upper incomplete Gamma functions. The left and right tail asymptotic distributions of this extreme modulus have also been investigated for large sizes of complex Ginibre matrices. They are the Rayleigh distribution with parameter $\sigma=1/\sqrt{2}$ and the Weibull distribution with shape parameter $\kappa=4$ and scale parameter $\lambda=k^{1/k}$, respectively. Derived for the non-Hermitian ensembles considered in this paper, the analytical formulation of the probability density functions of the extreme eigenvalue moduli (spectral radius and minimum modulus with and without scaling) exactly coincides with its empirical counterpart sampled from thousands of matrices. The independence of these random variables has also been established for these two random matrix ensembles at appropriate scalings. And it is acknowledged to hold as $N$ goes to infinity. As final new results presented here, the minimum moduli of the complex Ginibre ensemble and the complex induced ensemble, for values of rectangularity indexes close to zero, have the same limiting left and right tail distributions, the Rayleigh and the Weibull distributions, respectively. The exact limiting distribution of the minimum modulus for the complex Ginibre ensemble, as $N$ goes to infinity, is absent from the literature. This is left for forthcoming research.
Appendix A Proof of the Limit theorems for the complex induced Ginibre ensemble
A.1. Proof of Theorem 9
For the convenience of the reader, Theorem 9 is restated here.
Theorem (Limiting distribution of the scaled spectral radius $R_{N}$ for the complex induced Ginibre ensemble).
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}$ denote the scaled spectral radius. Then,
$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[R_{N}\leq 1+\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}+\frac{x}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}\right]=e^{-e^{-x}}$ (32)
where $\gamma_{\alpha,N}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$.
This limiting probability distribution is the standard Gumbel distribution for maxima. The corresponding cumulative density function is the function $F_{X}^{(\text{Gumbel max})}(x)=\exp(-\exp(-x))$.
The scaled spectral radius $R_{N}$ is approximated as
$R_{N}\simeq 1+T_{\alpha,N}+\xi_{\alpha,N}$ (33)
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ with $X$ following a standard Gumbel (maximum) distribution.
Proof.
At the outer edge of the ring, i.e., $r_{out}=1$,
$P\left(\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}\leq a\right)$ $=\frac{1}{\pi^{N}\prod_{k=1}^{N}\Gamma(k+L)}\prod_{k=0}^{N-1}\pi\int_{0}^{(1+\alpha)Na^{2}}t^{k+L}e^{-t}dt$
$=\frac{1}{\prod_{k=1}^{N}\Gamma(k+L)}\prod_{k=0}^{N-1}\int_{0}^{(1+\alpha)Na^{2}}t^{k+L}e^{-t}dt$
$=\prod_{k=1}^{N}\int_{0}^{(1+\alpha)Na^{2}}\frac{t^{(k+L-1)}e^{-t}}{\Gamma(k+L)}dt$
$=\prod_{k=1}^{N}\int_{0}^{(1+\alpha)Na^{2}}f_{Gamma(k+L,1)}(t)dt=\prod_{k=1}^{N}\frac{\gamma(k+L,(1+\alpha)Na^{2})}{\Gamma(k+L)}$
The function $f_{Gamma(k+L,1)}(t)$ is the probability density function of the Gamma distribution with shape parameter $k+L$ and rate parameter $1$.
This implies that,
$P\left(\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}\leq a\right)$ $=\prod_{k=1}^{N}P\left(\frac{1}{(1+\alpha)N}\sum_{j=1}^{N+L-k+1}Z_{j}\leq a^{2}\right)$
where the $Z_{j}$ for $j=\{1,\cdots,N-k+L+1\}$ are independent and identically distributed random variables following an exponential distribution with parameter 11 $\left(i.e.,Z_{j}\sim\exp(1)\right)$.
Near the edge of the outer circle of the ring, let $a=\frac{r_{out}}{\sqrt{(1+\alpha)N}}+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}$ denote the scaled radius close to $r_{out}$ with appropriate scaling $\sqrt{(1+\alpha)N}$. The probability
$\mathcal{P}_{N}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)$ $=\prod_{k=1}^{N}P\left(\frac{1}{(1+\alpha)N}\sum_{j=1}^{N+L-k+1}Z_{j}\leq a^{2}\right)$
$=\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{(1+\alpha)N}}\sum_{j=1}^{N+L-k+1}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k-1}{\sqrt{(1+\alpha)N}}\right)=\prod_{k=0}^{N-1}p_{k}$
where
$p_{k}=P\left(\frac{1}{\sqrt{(1+\alpha)N}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
and $\phi_{N}(x)=2f_{N}(x)+\frac{f_{N}^{2}(x)}{\sqrt{(1+\alpha)N}}=2f_{N}(x)\left(1+\frac{f_{N}(x)}{2\sqrt{(1+\alpha)N}}\right)$. The function $f_{N}(x)=o(\sqrt{(1+\alpha)N})$, meaning that $\lim_{N\rightarrow+\infty}\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=0$. The probability $\mathcal{P}_{N}(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}})$ is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf. [22]). More precisely,
$\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}p_{k}\leq\mathcal{P}_{N}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)\leq\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$
The following is derived using the same arguments as stated in [22] and applying the Markov inequality and the definition of the quantile of the exponential distribution. The function $\phi_{N}(x)$ is a positive and increasing function. This implies that
$\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}P\left(\frac{1}{\sqrt{(1+\alpha)N}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
$=\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}P\left(\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\sqrt{(1+\alpha)N}\phi_{N}(x)+k\right)$
The random variables $Z_{j}$, $i\in\{1,\cdots,(1+\alpha)N-k\}$ are independent and identically distributed. Applying the Markov inequality, with $0<\eta<1$ , as the exponential is an strictly increasing and convex function, this implies that
$\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}P\left(\frac{1}{\sqrt{(1+\alpha)N}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
$=\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}\left[1-P\left(\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)>\sqrt{(1+\alpha)N}\phi_{N}(x)+k\right)\right]$
$=\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}P\left(\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\sqrt{(1+\alpha)N}\phi_{N}(x)+k\right)$
$=\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}P\left(\sum_{j=1}^{(1+\alpha)N-k}Z_{j}\leq\sqrt{(1+\alpha)N}\phi_{N}(x)+(1+\alpha)N\right)$
$\geq\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}P\left(\sum_{j=1}^{(1+\alpha)N-k}Z_{j}\leq\sqrt{(1+\alpha)N}\phi_{N}(x)\right)$
$=\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}\left[1-P\left(\sum_{j=1}^{(1+\alpha)N-k}Z_{j}>\sqrt{(1+\alpha)N}\phi_{N}(x)\right)\right]$
$\geq\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}\left[1-e^{-\eta\sqrt{(1+\alpha)N}\phi_{N}(x)}E\left[e^{\eta Z_{1}}\right]^{(1+\alpha)N-k}\right]$
Furthermore,
$e^{-\eta(1+\alpha)N(\frac{k}{(1+\alpha)N}-1)-(1+\alpha)N\ln(1-\eta)}$ $\geq e^{((1+\alpha)N-k)(\eta-\ln(1-\eta))}$
$=e^{-\eta(1+\alpha)N(\frac{k}{(1+\alpha)N}-1)-\left((1+\alpha)N-k\right)\ln(1-\eta)}>1$
And with $0<\eta<1$ and $\forall Y\in\mathbb{R}$
$e^{Y-\eta\sqrt{(1+\alpha)N}\phi_{N}(x)}<e^{Y}\Rightarrow 1-e^{Y-\eta\sqrt{(1+\alpha)N}\phi_{N}(x)}>1-e^{Y}$
Now, setting $\eta=1-\frac{1}{\frac{k}{(1+\alpha)N}-1}$ which does maximise the remaining exponent (similarly presented in [22]) and applying the quantile formula of the exponential distribution, this implies
$\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}P\left(\frac{1}{\sqrt{(1+\alpha)N}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
$\geq\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}\left(1-e^{-(1+\alpha)N\left[\eta\left(\frac{k}{(1+\alpha)N}-1\right)+\ln(1-\eta)\right]}\right)$
$\geq\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}\left(1-e^{-(1+\alpha)N\left[\frac{k}{(1+\alpha)N}-\ln\left(\frac{k}{(1+\alpha)N}-1\right)\right]}\right)$
$\geq\left(1-e^{-(1+\alpha)N\delta_{N}^{2}}\right)^{N}$
Finally,
$\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}p_{k}\geq\left(1-e^{-(1+\alpha)N\delta_{N}^{2})}\right)^{N}$
The parameter $\delta_{N}$ is chosen such that $\left(1-e^{-(1+\alpha)N\delta_{N}^{2})}\right)^{N}=\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$ and is defined as follows
$\delta_{N}=\sqrt{\frac{2\log{N}}{(1+\alpha)N}}$
Consequently,
$\lim_{N\rightarrow+\infty}\left[\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}p_{k}\right]=\lim_{N\rightarrow+\infty}\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$
Now, applying the squeeze theorem, uniformly in $N$ and $x$, for bounded $x>$-$\infty$,
$\lim_{N\rightarrow+\infty}\log\left(\mathcal{P}_{N}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)\right)$ $=\lim_{N\rightarrow+\infty}\log\left(\prod_{k=0}^{\sqrt{2(1+\alpha)N\log(N)}}p_{k}\right)$
Also,
$\prod_{k=0}^{(1+\alpha)N\delta_{N}}P\left(\frac{1}{\sqrt{(1+\alpha)N}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
$=\prod_{k=0}^{(1+\alpha)N\delta_{N}}P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\left(\sqrt{\frac{(1+\alpha)N}{N+L-k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)\right)$
$=\prod_{k=0}^{(1+\alpha)N\delta_{N}}P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
$=\prod_{k=0}^{(1+\alpha)N\delta_{N}}P\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)<\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
$=\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$
where
$p_{k}=P\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
As in the reference [22], it is assumed that $k=o((1+\alpha)N)$. The factor $\sqrt{\frac{(1+\alpha)N}{N+L-k}}$ is assumed equal (approximatively) to one for small and moderate values of $k$. For large values of $k$, this factor is rather large and the probability
$P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\left(\sqrt{\frac{(1+\alpha)N}{N+L-k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)\right)$ is approximatively equal to one. This means that its logarithm is zero.
Finally,
$\lim_{N\rightarrow+\infty}\log\left(\mathcal{P}_{N}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)\right)$
$=\lim_{N\rightarrow+\infty}\sum_{k=0}^{\sqrt{2(1+\alpha)N\log{N}}}\log P\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
where $\delta_{N}=\sqrt{\frac{2\log{N}}{(1+\alpha)N}}$.
The Edgeworth expansion is then used to get the probability density of the following standardised random variable
$\frac{\sqrt{(1+\alpha)N-k}(\bar{Z}-\mu)}{\sigma}$ $=\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)$
where each element of sequence $Z_{1},\cdots,Z_{(1+\alpha)N-k}$ is i.i.d. exponentially distributed with parameter equal to $1$.
The random variable $\bar{Z}=\frac{1}{(1+\alpha)N-k}\sum_{j=1}^{(1+\alpha)N-k}Z_{j}$ is the empirical mean. The mean of $\bar{Z}$ is $\mu=1$ and its standard deviation is $\frac{\sigma}{\sqrt{(1+\alpha)N-k}}=\frac{1}{\sqrt{(1+\alpha)N-k}}$ with $\sigma=1$.
Applying the Edgeworth expansion, the logarithm of the probability $p_{k}$ is then
$\log\left(P\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)\right)$
$=\log\begin{pmatrix}\int_{-K_{N}}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt+O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)\\ +O\left(\frac{1}{(1+\alpha)N-k}\right)+O\left(K_{N}((1+\alpha)N-k)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)\end{pmatrix}$
$=\log\left(\int_{-K_{N}}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)+O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)$
$+O\left(\frac{1}{(1+\alpha)N-k}\right)+O\left(K_{N}((1+\alpha)N-k)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)$
where, as in [22], for $x$ restricted as in $|x|<Y$ for some large positive $Y$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq Y}\phi_{N}(x)$.
Remark A.1.
Let
$T_{N}=\int_{-K_{N}}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt$
For $z$ in the neighbourhood of zero, it is known that the Taylor expansion of $\log(1+z)=z+O(z^{2})$ and $\lim_{N\longrightarrow+\infty}T_{N}=1$.
Then, in the limit of large values of $N$,
$\log\begin{pmatrix}T_{N}+O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)+O\left(\frac{1}{(1+\alpha)N-k}\right)\\ +O\left(K_{N}((1+\alpha)N-k)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)\end{pmatrix}$
$=\log\left[T_{N}\begin{pmatrix}1+\frac{1}{T_{N}}\begin{pmatrix}O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)\\ +O\left(\frac{1}{(1+\alpha)N-k}\right)+O\left(K_{N}((1+\alpha)N-k)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)\end{pmatrix}\end{pmatrix}\right$]
$=\log\left[1+\begin{pmatrix}O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)\\ +O\left(\frac{1}{(1+\alpha)N-k}\right)+O\left(K_{N}((1+\alpha)N-k)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)\end{pmatrix}\right$]
$=\log\left[\int_{-K_{N}}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
$+\log\left[1+\begin{pmatrix}O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)\\ +O\left(\frac{1}{(1+\alpha)N-k}\right)+O\left(K_{N}((1+\alpha)N-k)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)\end{pmatrix}\right$]
$=\log\left[\int_{-K_{N}}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)$
$+O\left(\frac{1}{(1+\alpha)N-k}\right)+O\left(K_{N}((1+\alpha)N-k)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)$
As mentioned in [22], the lower limit of integration in the leading term of the logarithm of the probability $p_{k}$ is extended from $K_{N}$ down to -$\infty$. Defining the function $f(k)$ as
$f(k)=\log\left[\int_{-\infty}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
$\sum_{k=0}^{\sqrt{2(1+\alpha)N\log{N}}}f(k)$ $=\int_{0}^{\sqrt{2(1+\alpha)N\log{N}}}\log\left[\int_{-\infty}^{\phi_{N}(x)+\frac{t}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right]dt+E_{N}$
where $E_{N}$ is an error term.
Using the change of variables $u=\phi_{N}(x)+\frac{t}{\sqrt{(1+\alpha)N}}$, this implies that $du=\frac{dt}{\sqrt{(1+\alpha)N}}$ and $dt=\sqrt{(1+\alpha)N}du$
$\sum_{k=0}^{\sqrt{2(1+\alpha)N\log N}}f(k)=\sqrt{(1+\alpha)N}\int_{\phi_{N}(x)}^{\phi_{N}(x)+\sqrt{2\log{N}}}\log\left[\int_{-\infty}^{u}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right]du+E_{N}$
where
$E_{N}=O\left(\left(\frac{\sqrt{\log N}}{\sqrt{(1+\alpha)N}}\right)\vee\left(\sqrt{\log{N}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)\right)$
Remark A.2.
Identification of the error term corresponding to the sum of the error of integration at each integration point $\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}$.
With $\phi_{N}(x)=o\left(\sqrt{\log\left((1+\alpha)N\right)}\right)$ and $K_{N}=O(\log{(1+\alpha)N})$, this implies that
$\left|\sqrt{(1+\alpha)N}\sum_{k=0}^{\sqrt{2(1+\alpha)N\log N}}\int_{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}^{\phi_{N}(x)+\frac{k+1}{\sqrt{(1+\alpha)N}}}\log\left(1+\int_{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}^{s}e^{-\frac{t^{2}}{2}}dt\right)ds\right|$
$\leq C\left|\sqrt{\log{N}}e^{-\frac{\phi_{N}^{2}(x)}{2}}\right|=C\left|\sqrt{\log{N}}\left((1+\alpha)N\right)^{-1/2}\right|$
where $C$ is a positive constant.
This implies that
$\sqrt{(1+\alpha)N}\sum_{k=0}^{\sqrt{2(1+\alpha)N\log N}}\int_{\phi_{N}+\frac{k}{\sqrt{(1+\alpha)N}}}^{\phi_{N}+\frac{k+1}{\sqrt{(1+\alpha)N}}}\log\left(1+\int_{\phi_{N}+\frac{k}{\sqrt{(1+\alpha)N}}}^{s}e^{-\frac{t^{2}}{2}}dt\right)ds=O\left(\frac{\sqrt{\log(N)}}{\sqrt{(1+\alpha)N}}\right)$
This completes the remark.
The term $\frac{1}{\sqrt{(1+\alpha)N}}$ appearing in the lower limit of integration is negligible and the function $\phi_{N}(x)=o\left(\sqrt{\log\left((1+\alpha)N\right)}\right)$.
This implies the following result
$\log{\mathcal{P}_{N}}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)\approx\sqrt{(1+\alpha)N}\int_{\phi_{N}(x)}^{+\infty}\log\left[\int_{-\infty}^{t}\frac{e^{-\frac{s^{2}}{2}}}{2\pi}ds\right]dt+E_{N}$
Let $F_{\infty}(x)$ denote the limiting distribution of the scaled spectral radius of any $N\times N$ complex induced Ginibre matrix as $N$ goes to infinity. The limit of its logarithm is
$\log{F_{\infty}}(x)=\lim_{N\rightarrow+\infty}\sqrt{(1+\alpha)N}\int_{\phi_{N}(x)}^{+\infty}\log\left[1-\int_{t}^{+\infty}e^{-\frac{s^{2}}{2}}\frac{ds}{\sqrt{2\pi}}\right]dt$
where $\lim_{N\rightarrow+\infty}E_{N}=0$.
The integral $\int_{t}^{+\infty}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds$ is less than $1$ and is close to zero such that, applying the Taylor expansion of the function $\log(1-x)\approx-x+O(x^{2})$, the logarithm of the limiting distribution $F_{\infty}(x)$ is
$\log{F_{\infty}}(x)$ $=-\lim_{N\rightarrow+\infty}\sqrt{\frac{(1+\alpha)N}{2\pi}}\int_{\phi_{N}(x)}^{+\infty}\left[\int_{t}^{+\infty}e^{-\frac{s^{2}}{2}}ds+O\left(\left(\int_{t}^{+\infty}e^{-\frac{s^{2}}{2}}ds\right)^{2}\right)\right]dt$
$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{(1+\alpha)N}{2\pi}}\int_{\phi_{N}(x)}^{+\infty}\left[\frac{1}{t}e^{-t^{2}/2}+O\left(\frac{e^{-t^{2}}}{t^{2}}\right)\right]dt$
Furthermore, $\phi_{N}(x)=2f_{N}(x)+\frac{f_{N}^{2}(x)}{\sqrt{(1+\alpha)N}}=2f_{N}(x)\left(1+\frac{f_{N}(x)}{2\sqrt{(1+\alpha)N}}\right)$. The function $f_{N}(x)=o(\sqrt{(1+\alpha)N})$, meaning that $\lim_{N\rightarrow+\infty}\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=0$.
This implies that
$\lim_{N\rightarrow+\infty}\phi_{N}^{2}(x)$ $=\lim_{N\rightarrow+\infty}4f_{N}^{2}(x)\left(1+\frac{f_{N}(x)}{2\sqrt{(1+\alpha)N}}\right)^{2}$
$=\lim_{N\rightarrow+\infty}4f_{N}^{2}(x)\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}+\frac{f_{N}^{2}(x)}{4(1+\alpha)N}\right)=\lim_{N\rightarrow+\infty}4f_{N}^{2}(x)$
Also,
$\int_{\phi_{N}(x)}^{+\infty}\frac{1}{t}e^{-t^{2}/2}dt=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}-2\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt$
and in the limit as $N$ goes to infinity
$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt\approx\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times O\left(\frac{1}{\phi_{N}^{2}(x)}\right)$
This implies that
$\log{F_{\infty}}(x)$ $=-\lim_{N\rightarrow+\infty}\sqrt{\frac{(1+\alpha)N}{2\pi}}\left[\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\left(1+O\left(\frac{1}{\phi_{N}^{2}(x)}\right)\right)\right]$
$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{(1+\alpha)N}{2\pi}}\left[\frac{1}{4f_{N}^{2}(x)}e^{-2f_{N}^{2}(x)}\left(1+O\left(\frac{1}{4f_{N}^{2}(x)}\right)\right)\right]$
The function $f_{N}(x)$ is chosen with the use of the definition of a class of limiting distributions exposed in the reference [14]. For a convenient choice of the function $f_{N}(x)$, the limiting distribution $F_{\infty}(x)$ is from the class of Extreme Value Distributions composed of the three types of extreme value distributions for maxima.
$f_{N}^{2}(x)=\frac{1}{2}\log\left(\frac{e^{x}\sqrt{(1+\alpha)N/2\pi}}{\log N}\right)$
$\lim_{N\rightarrow+\infty}\log\mathcal{P}_{N}\left[R_{N}\leq 1+\sqrt{\frac{1}{2(1+\alpha)N}}\left(\log\frac{\sqrt{(1+\alpha)N/2\pi}}{\log{N}}+x\right)^{1/2}\right]$
$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{(1+\alpha)N}{2\pi}}\left[\frac{1}{4f_{N}^{2}(x)}e^{-2f_{N}^{2}(x)}\left(1+O\left(\frac{1}{4f_{N}^{2}(x)}\right)\right)\right]$
$=-\lim_{N\rightarrow+\infty}\exp(-x)\frac{\log{N}}{2\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}$
And, for any $\alpha>0$ and fixed $|x|\ll N$
$\frac{\log{N}}{2\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}$
$=\frac{1}{2}\frac{\log{N}}{\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}$
$=\frac{1}{2}\frac{\log{N}}{\left[x+\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N)\right]}$
$\sim\frac{1}{2}\frac{\log{N}}{\left[\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N\right]}$
$\sim\frac{\log{N}}{\log((1+\alpha)N/(2\pi))}$
$=\frac{1}{1+\frac{\log\left(\frac{1+\alpha}{2\pi}\right)}{\log N}}$
This implies
$\lim_{N\rightarrow+\infty}\frac{\log{N}}{2\log(e^{x}\sqrt{(1+\alpha)N/2\pi}\times\frac{1}{\log{N}})}=1$
Finally, let $R_{N}$ denote the scaled spectral radius.
$\lim_{N\rightarrow+\infty}\log\mathcal{P}_{N}\left[R_{N}\leq 1+\sqrt{\frac{1}{2(1+\alpha)N}}\left(\log\frac{\sqrt{(1+\alpha)N/2\pi}}{\log{N}}+x\right)^{1/2}\right]=-\exp(-x)$
This limit is the logarithm of the cumulative distribution function of the standard Gumbel distribution for maxima.
Furthermore, setting $\gamma_{\alpha,N}=\log\frac{\sqrt{(1+\alpha)N/2\pi}}{\log{N}}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$ and using the Taylor expansion of the square root function
$\left(\log\frac{\sqrt{(1+\alpha)N/2\pi}}{\log{N}}+x\right)^{1/2}=\gamma_{\alpha,N}^{1/2}\left(1+\frac{x}{\gamma_{\alpha,N}}\right)^{1/2}=\gamma_{\alpha,N}^{1/2}+\frac{x}{2\gamma_{\alpha,N}^{1/2}}+O\left(\left(\frac{x}{\gamma_{\alpha,N}}\right)^{2}\right)$
which implies
$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[R_{N}\leq 1+\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}+\frac{x}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}\right]=\exp(-\exp(-x))$
This limiting distribution is the standard Gumbel distribution of maxima. The corresponding cumulative density function is the function $F_{Gumbel}^{(max)}(x)=\exp(-\exp(-x))$.
The scaled spectral radius $R_{N}$ is a random variable approximated as
$R_{N}$ $\simeq 1+\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}+\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$
$=1+\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}-\frac{\log(Z)}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$
$=1+T_{\alpha,N}+\xi_{\alpha,N}$
where the term $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$ with $\gamma_{\alpha N}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$. The random variable $Z$ denote a random variable following a standard exponential distribution which implies that $X=-\log(Z)$ is a standard Gumbel(maximum)-distributed random variable. The random variable $\xi_{\alpha,N}=-\frac{\log(Z)}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ is Gumbel-distributed. This completes the proof of Theorem 9.
A.2. Proof of Theorem 10
It is demonstrated in Section 2 that the scaled minimum radius $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ of eigenvalues for matrices from the complex Ginibre ensemble, in the neighbourhood of zero, follows a Rayleigh distribution. The complex Ginibre ensemble does correspond to the complex induced Ginibre ensemble when the rectangularity index $L$ is equal to zero. Similarly stated in Section 3, Theorem 10 is as follows.
Theorem (Limiting distribution of the scaled minimum modulus $r_{N}$ for the complex induced Ginibre ensemble).
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $r_{N}$ denote the scaled minimum modulus such that $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$. Then,
$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]=e^{-e^{x}}$ (34)
where $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$.
This limit corresponds to the survival distribution function of the standard Gumbel distribution for minima. The scaled minimum modulus $r_{N}$ is approximated with the following
$r_{N}\simeq 1-T_{\alpha,N}+\xi_{\alpha,N}$ (35)
where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$. The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ where $X$ follows a standard Gumbel (minimum) distribution.
Proof.
The limiting distribution of the scaled minimum radius at the inner edge of the ring, in the limit as $N$ goes to infinity and with the rectangularity index $L$ proportional to $N$, i.e., $L=\alpha N$, $\alpha>0$, is studied in this section.
At the inner edge of the ring, i.e., for the inner radius $r_{in}=1$, the chosen scaling is $\sqrt{L}$. The survival probability of the scaled minimum of eigenvalues moduli, denoted $\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$, is
$P\left(\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}\geq a\right)$ $=\frac{1}{\prod_{k=1}^{N}\Gamma(k+\alpha N)}\prod_{k=0}^{N-1}\int_{\alpha Na^{2}}^{+\infty}t^{k+L}e^{-t}dt$
$=\prod_{k=1}^{N}\int_{\alpha Na^{2}}^{+\infty}\frac{t^{k+L-1}e^{-t}}{\Gamma(k+L)}dt=\prod_{k=1}^{N}\int_{\alpha Na^{2}}^{+\infty}f_{Gamma(k+L,1)}dt$
$=\prod_{k=1}^{N}\frac{\Gamma\left(k+\alpha N,\alpha Na^{2}\right)}{\Gamma(k+\alpha N)}$
The function $f_{Gamma(k+L,1)}(t)$ is the probability density function of the Gamma distribution with shape parameter $k+L$ and rate parameter $1$.
which implies,
$P\left(\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}\geq a\right)=\prod_{k=1}^{N}P\left(\sum_{j=1}^{k+L}Z_{j}\geq\alpha Na^{2}\right)=\prod_{k=1}^{N}P\left(\frac{1}{\alpha N}\sum_{j=1}^{k+L}Z_{j}\geq a^{2}\right)$
where the $Z_{j}$ for $j=\{1,\cdots,k+L\}$ are independent and identically distributed random variables following an exponential distribution with parameter 11 $\left(\text{i.e., }Z_{j}\sim\exp(1)\right)$.
Let $a=1-\frac{f_{N}(x)}{\sqrt{L}}=1-\frac{f_{N}(x)}{\sqrt{\alpha N}}$, with $\alpha>0$, this implies
$\mathcal{P}_{N}\left(1-\frac{f_{N}(x)}{\sqrt{\alpha N}}\right)$ $=\mathcal{P}_{N}\left(\frac{\min_{1\leq k\leq N}|z_{k}|}{\sqrt{\alpha N}}\geq 1-\frac{f_{N}(x)}{\sqrt{\alpha N}}\right)=\prod_{k=1}^{N}p_{k}$
where
$p_{k}=P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
and $\phi_{N}(x)=2f_{N}(x)\left(1-\frac{f_{N}(x)}{2\sqrt{\alpha N}}\right)$. The function $f_{N}(x)=o\left(\sqrt{\alpha N}\right)$, meaning that $\lim_{N\rightarrow+\infty}\frac{f_{N}(x)}{\sqrt{\alpha N}}=0$.
The probability $\mathcal{P}_{N}(1-\frac{f_{N}(x)}{\sqrt{\alpha N}})$ is bounded by any partial product of $p_{k}$, i.e., bounded by the partial product $\prod_{k=0}^{\alpha N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf. [22]), such that
$\delta_{N}=\sqrt{\frac{2\log N}{\alpha N}}$
Also,
$\prod_{k=1}^{k=\alpha N\delta_{N}}p_{k}\prod_{\alpha N\delta_{N}}^{N}p_{k}\leq\mathcal{P}_{N}\left(1-\frac{f_{N}(x)}{\sqrt{\alpha N}}\right)\leq\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$
And the random variable $\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)$ converges in distribution to a standard normal variable as $N$ goes to infinity (The Central Limit Theorem).
Thus,
$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
$\approx\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-\frac{1}{2}\left(1+\operatorname{erf}\left(-\frac{\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)}{\sqrt{2}}\right)\right)\right]$
$>\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\frac{k^{2}}{\alpha N}}\right)>\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\alpha N\delta_{N}^{2}}\right)>\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$
Consequently,
$\lim_{N\rightarrow+\infty}\left[\prod_{k=1}^{\alpha N\delta_{N}}p_{k}\prod_{k=\alpha N\delta_{N}}^{N}p_{k}\right]=\lim_{N\rightarrow+\infty}\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$
Now, applying the squeeze theorem, uniformly in $N$ and $x$, for bounded $x>$-$\infty$,
$\lim_{N\rightarrow+\infty}\log\left(\mathcal{P}_{N}\left(1-\frac{f_{N}(x)}{\sqrt{\alpha N}}\right)\right)=\lim_{N\rightarrow+\infty}\log\left(\prod_{k=1}^{\alpha N\delta_{N}}p_{k}\right)=\lim_{N\rightarrow+\infty}\log\left(\prod_{k=1}^{\sqrt{2\alpha N\log(N)}}p_{k}\right)$
Also,
$\prod_{k=1}^{\alpha N\delta_{N}}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=1}^{\alpha N\delta_{N}}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\left(\sqrt{\frac{\alpha N}{\alpha N+k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)\right)$
$=\prod_{k=1}^{\alpha N\delta_{N}}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=1}^{\alpha N\delta_{N}}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$
where
$p_{k}=P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
As in the reference [22], it is assumed that $k=o(\alpha N)$. The factor $\sqrt{\frac{\alpha N}{\alpha N+k}}$ is assumed equal (approximatively) to one for small and moderate values of $k$. For large values of $k$, this factor is rather large and the probability
$P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\left(\sqrt{\frac{\alpha N}{\alpha N+k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)\right)$ is approximatively equal to one. This means that its logarithm is zero.
Finally,
$\lim_{N\rightarrow+\infty}\log\left(\mathcal{P}_{N}\left(1-\frac{f_{N}(x)}{\sqrt{\alpha N}}\right)\right)$
$=\lim_{N\rightarrow+\infty}\sum_{k=1}^{\sqrt{2\alpha N\log{N}}}\log P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
where $\delta_{N}=\sqrt{\frac{2\log{N}}{\alpha N}}$.
The Edgeworth expansion is then used to get the probability density of the standardised random variable denoted
$\frac{\sqrt{\alpha N+k}(\bar{Z}-\mu)}{\sigma}$ $=\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)$
where each element of sequence $Z_{1},\cdots,Z_{\alpha N+k}$ is i.i.d. exponentially distributed with parameter equal to $1$.
The variable $\bar{Z}=\frac{1}{\alpha N+k}\sum_{j=1}^{\alpha N+k}Z_{j}$ is the empirical mean. The mean of $\bar{Z}$ is $\mu=1$ and its standard deviation is
$\frac{\sigma}{\sqrt{\alpha N+k}}=\frac{1}{\sqrt{\alpha N+k}}$ with $\sigma=1$.
Applying the Edgeworth expansion, the logarithm of the probability $p_{k}$ is then
$\log\left[P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)\right]$
$=\log\left[\int_{-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}}^{+K_{N}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N+k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)+O\left(\frac{1}{\alpha N+k}\right)$
$+O\left(K_{N}(\alpha N+k)^{-3/2}\right)+O\left(e^{-\frac{1}{2}K_{N}^{2}}\right)$
where, similarly stated in [22], for $x$ restricted as in $|x|<Y$ for some large positive $Y$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq Y}\phi_{N}(x)$.
Defining the function $f(k)$ as
$f(k)=\log\left[\int_{-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right]$
Then,
$\sum_{k=1}^{\sqrt{2\alpha N\log N}}f(k)=\int_{1}^{\sqrt{2\alpha N\log N}}\log\left[\int_{-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}\right]dt+E_{N}$
Let $u=\phi_{N}(x)+\frac{t}{\sqrt{\alpha N}}$ which implies that $dt=\sqrt{\alpha N}du$.
$\sum_{k=1}^{\sqrt{2\alpha N\log N}}f(k)=\sqrt{\alpha N}\int_{\phi_{N}(x)+\frac{1}{\sqrt{\alpha N}}}^{\phi_{N}(x)+\sqrt{2\log N}}\log\left[\int_{-u}^{+\infty}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right]du+E_{N}$
where $E_{N}$ is the error term
$E_{N}=O\left(\left(\frac{\sqrt{\log N}}{\sqrt{\alpha N}}\right)\vee\left(\sqrt{\log{N}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)\right)$
The term $\frac{1}{\sqrt{\alpha N}}$ appearing in the lower limit of integration is negligible and the function $\phi_{N}(x)=o\left(\sqrt{\log\left(\alpha N\right)}\right)$. This implies the following result
$\lim_{N\rightarrow+\infty}\log\mathcal{P}_{N}\left(1-\frac{f_{N}(x)}{\sqrt{\alpha N}}\right)=\sqrt{\alpha N}\int_{\phi_{N}(x)}^{+\infty}\log\left[\int_{-t}^{+\infty}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right]dt+E_{N}$
Let $F_{\infty}(x)$ denote the limiting distribution of the scaled minimum modulus of any complex induced Ginibre matrix as $N$ goes to infinity. The logarithm
$\log\left[\int_{-t}^{+\infty}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right]=\log\left[1-\int_{-\infty}^{-t}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right]$
and $\lim_{N\rightarrow+\infty}E_{N}=0$.
The integral $\int_{-\infty}^{-t}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds$ is less than $1$ and is close to zero such that, applying the Taylor expansion of the function $\log(1-x)\approx-x+O(x^{2})$.
Thus, the logarithm of the limiting survival distribution function $1-F_{\infty}(x)$ is
$\log\left(1-F_{\infty}(x)\right)$ $=-\lim_{N\rightarrow+\infty}\sqrt{\frac{\alpha N}{2\pi}}\int_{\phi_{N}(x)}^{+\infty}\left[\int_{-\infty}^{-t}e^{-\frac{s^{2}}{2}}ds+O\left(\left(\int_{-\infty}^{-t}e^{-\frac{s^{2}}{2}}ds\right)^{2}\right)\right]dt$
$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{\alpha N}{2\pi}}\int_{\phi_{N}(x)}^{+\infty}\left[\int_{t}^{+\infty}e^{-\frac{s^{2}}{2}}ds+O\left(\left(\int_{t}^{+\infty}e^{-\frac{s^{2}}{2}}ds\right)^{2}\right)\right]dt$
$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{\alpha N}{2\pi}}\int_{\phi_{N}(x)}^{+\infty}\left[\frac{e^{-t^{2}/2}}{t}+O\left(\frac{e^{-t^{2}}}{t^{2}}\right)\right]dt$
Furthermore, $\phi_{N}(x)=2f_{N}(x)-\frac{f_{N}^{2}(x)}{\sqrt{\alpha N}}=2f_{N}(x)\left(1-\frac{f_{N}(x)}{2\sqrt{\alpha N}}\right)$.
The function $f_{N}(x)=o(\sqrt{\alpha N})$, meaning that $\lim_{N\rightarrow+\infty}\frac{f_{N}(x)}{\sqrt{\alpha N}}=0$.
This implies that
$\lim_{N\rightarrow+\infty}\phi_{N}^{2}(x)$ $=\lim_{N\rightarrow+\infty}4f_{N}^{2}(x)\left(1-\frac{f_{N}(x)}{2\sqrt{\alpha N}}\right)^{2}$
$=\lim_{N\rightarrow+\infty}4f_{N}^{2}(x)\left(1-\frac{f_{N}(x)}{\sqrt{\alpha N}}+\frac{f_{N}^{2}(x)}{4\alpha N}\right)=\lim_{N\rightarrow+\infty}4f_{N}^{2}(x)$
Also,
$\int_{\phi_{N}(x)}^{+\infty}\frac{1}{t}e^{-t^{2}/2}dt=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}-\int_{\phi_{N}(x)}^{+\infty}\frac{2e^{-t^{2}/2}}{t^{3}}dt$
and in the limit as $N$ goes to infinity
$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt\approx\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times O\left(\frac{1}{\phi_{N}^{2}(x)}\right)$
This implies that
$\lim_{N\rightarrow+\infty}\log\left(\mathcal{P}_{N}\left(1-\frac{f_{N}(x)}{\sqrt{\alpha N}}\right)\right)=\log\left(1-F_{\infty}(x)\right)$
$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{\alpha N}{2\pi}}\left[\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\left(1+O\left(\frac{1}{\phi_{N}^{2}(x)}\right)\right)\right]$
$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{\alpha N}{2\pi}}\left[\frac{1}{4f_{N}^{2}(x)}e^{-2f_{N}^{2}(x)}\left(1+O\left(\frac{1}{4f_{N}^{2}(x)}\right)\right)\right]$
For a convenient choice of the function $f_{N}(x)$, the limiting distribution $F_{\infty}(x)$ is from the class of Extreme Value Distributions composed of the three types of extreme value distributions for minima [14].
$f_{N}^{2}(x)=\frac{1}{2}\log\left(\frac{e^{-x}\sqrt{\alpha N/2\pi}}{\log N}\right)$
which implies
$\lim_{N\rightarrow+\infty}\log\mathcal{P}_{N}\left[r_{N}\geq 1-\sqrt{\frac{1}{2\alpha N}}\left(\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}-x\right)^{1/2}\right]$
$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{\alpha N}{2\pi}}\left[\frac{1}{4f_{N}^{2}(x)}e^{-2f_{N}^{2}(x)}\left(1+O\left(\frac{1}{4f_{N}^{2}(x)}\right)\right)\right]$
$=-\lim_{N\rightarrow+\infty}\exp(x)\frac{\log{N}}{2\log(e^{-x}\sqrt{\alpha N/2\pi}\times\frac{1}{\log{N}})}$
Furthermore, for $|x|\ll N$ and using the same reasoning applied for the derivation of the limiting distribution of the scaled spectral radius for matrices from the complex induced Ginibre ensemble
$\lim_{N\rightarrow+\infty}\frac{\log{N}}{2\log(e^{-x}\sqrt{\alpha N/2\pi}\times\frac{1}{\log{N}})}=1$
Finally,
$\lim_{N\rightarrow+\infty}\log\mathcal{P}_{N}\left[r_{N}\geq 1-\sqrt{\frac{1}{2\alpha N}}\left(\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}-x\right)^{1/2}\right]=-\exp(x)$
which is the logarithm of the survival probability function of the standard Gumbel distribution for minima.
Furthermore, setting $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$ and using the Taylor expansion of the square root function
$\left(\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}-x\right)^{1/2}=\gamma_{\alpha,N}^{1/2}\left(1-\frac{x}{\gamma_{\alpha,N}}\right)^{1/2}=\gamma_{\alpha,N}^{1/2}-\frac{x}{2\gamma_{\alpha,N}^{1/2}}+O\left(\left(\frac{x}{\gamma_{\alpha,N}}\right)^{2}\right)$
which implies that
$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\left(\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}-\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right)\right]$
$=\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]$
$=\exp(-\exp(x))$
The limiting distribution function of the scaled minimum modulus of eigenvalues for matrices from the complex induced Ginibre ensemble is
$F_{\infty}(x)=1-\exp\left(-\exp\left(x\right)\right)$
and corresponds to distribution function of the standard Gumbel distribution for minima.
The scaled minimum modulus denoted $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$ is a random variable well approximated as
$r_{N}$ $\simeq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
$=1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}-\frac{\log(Z)}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
$=1-T_{\alpha,N}+\xi_{\alpha,N}$
where the term $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$ with $\gamma_{\alpha N}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$. The random variable $Z$ denotes a random variable following a standard exponential distribution which implies that $X=-\log(Z)$ is a standard Gumbel(minimum)-distributed random variable. The random variable $\xi_{\alpha,N}=-\frac{\log(Z)}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ is Gumbel-distributed. This completes the proof of Theorem 10.
A.3. Proof of Theorem 13
The independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble is established from Theorem 13 presented here again for more convenience.
Theorem (Independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble).
Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with proportional rectangularity index $L=\alpha N$, $\alpha>0$. The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ are independent random variables, under scaling $\sqrt{\alpha N}$, as $N$ goes to infinity. More precisely, let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$.
Setting $\rho=\sqrt{\frac{1+\alpha}{\alpha}}$,
$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\xi_{\alpha}^{(N)}(y)\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\eta_{\alpha}^{(N)}(x)\right]=e^{-e^{y}}e^{-e^{-x}}$
with $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$ and where
$\xi_{\alpha}^{(N)}(y)=\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }\eta_{\alpha}^{(N)}(x)=\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
The scaled spectral radius $R_{N}$ and the scaled minimum of moduli $r_{N}$ are approximated as
$R_{N}$ $\simeq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }r_{N}\simeq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{Y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
where $X$ and $Y$ are standard Gumbel (maximum) and standard Gumbel (minimum) distributed random variables, respectively.
Proof.
The independence of the scaled spectral radius and the scaled minimum modulus is derived from their joint cumulative distribution function
$P\left(r_{N}\geq r\text{}\text{ and }\text{}R_{N}\leq R\right)$ $=P\left(r_{min}^{(N)}(G)\geq\sqrt{\alpha N}r\text{}\text{ and }\text{}r_{max}^{(N)}(G)\leq\sqrt{\alpha N}R\right)$
$=\frac{1}{\prod_{k=1}^{N}\Gamma(k+\alpha N)}\prod_{k=0}^{N-1}\int_{\alpha Nr^{2}}^{\alpha NR^{2}}t^{k+L}e^{-t}dt$
$=\prod_{k=1}^{N}\int_{\alpha Nr^{2}}^{\alpha NR^{2}}f_{Gamma(k+L,1)}dt$
The function $f_{Gamma(k+L,1)}(t)$ is the probability density function of the Gamma distribution with shape parameter $k+L$ and rate parameter $1$.
Finally,
$P\left(r_{min}^{(N)}(G)\geq\sqrt{\alpha N}r\text{}\text{ and }\text{}r_{max}^{(N)}(G)\geq\sqrt{\alpha N}R\right)$ $=\prod_{k=1}^{N}P\left(r^{2}\leq\frac{1}{\alpha N}\sum_{j=1}^{\alpha N+k}Z_{j}\leq R^{2}\right)$
where the $Z_{j}$ for $j=\{1,\cdots,k+L\}$ are independent and identically distributed random variables following a standard exponential distribution.
Let $r_{in}=\sqrt{L}$ denote the inner radius and $r_{out}=\sqrt{L+N}$ the outer radius defining the edge of the eigenvalues support for matrices from the complex induced Ginibre ensemble. The rectangularity index $L$ is assumed proportional to $N$, i.e., $L=\alpha N$, $\alpha>0$.
Considering the scaling $\sqrt{\alpha N}$, the inner radius $r_{in}=1$ and the outer radius $r_{out}=\sqrt{\frac{1+\alpha}{\alpha}}=\gamma$. Let $r=1-\frac{f_{N}(y)}{\sqrt{\alpha N}}$ and $R=\sqrt{\frac{1+\alpha}{\alpha}}+\frac{f_{N}(x)}{\sqrt{\alpha N}}=\rho+\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ and $f_{N}(y)$ are $o\left(\alpha N\right)$.
This implies
$P\left(r_{min}^{(N)}(G)\geq\sqrt{\alpha N}r\text{}\text{ and }\text{}r_{max}^{(N)}(G)\leq\sqrt{\alpha N}R\right)$
$=\prod_{k=1}^{N}P\left(r^{2}\leq\frac{1}{\alpha N}\sum_{j=1}^{k+L}Z_{j}\leq R^{2}\right)$
$=\prod_{k=1}^{N}\left[P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\right]$
where $C=\rho^{2}-1>0$. The functions $\phi_{N}^{(r)}(y)=2f_{N}(y)\left(1-\frac{f_{N}(y)}{2\sqrt{\alpha N}}\right)$ and
$\phi_{N}^{(R)}(x)=2\rho f_{N}(x)\left(1+\frac{f_{N}(x)}{2\rho\sqrt{\alpha N}}\right)$.
Furthermore, with proportional rectangularity index $L=\alpha N$, $\alpha>0$,
$\prod_{k=1}^{N}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\times$
$\prod_{k=1}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$
$=\prod_{k=1}^{N}p_{k}\prod_{k=1}^{N}h_{k}$
where
$p_{k}=P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)<\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
and
$h_{k}=\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$
The partial product $\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$ is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf. [22]), such as
$\prod_{k=1}^{\alpha N\delta_{N}}p_{k}\prod_{k=\alpha N\delta_{N}}^{N}p_{k}\leq\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\leq\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$
The following is derived using the same arguments as stated in [22] and using the Markov inequality and the definition of the quantile of the exponential distribution.
The function $\phi_{N}^{(R)}(x)$ is a positive and increasing function.
$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\sqrt{\alpha N}\left(\phi_{N}^{(R)}(x)+C\right)-k\right)$
The random variables $Z_{j}$, $i\in\{1,\cdots,\alpha N+k\}$ are independent and identically distributed.
Applying the Markov inequality, with $0<\eta<1$ , as the exponential is an strictly increasing and convex function, this implies that
$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{(\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\sqrt{\alpha N}\left(\phi_{N}^{(R)}(x)+C\right)-k\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}Z_{j}\leq\sqrt{\alpha N}\left(\phi_{N}^{(R)}(x)+C\right)+\alpha N\right)$
$\geq\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}Z_{j}\leq\sqrt{\alpha N}\phi_{N}^{(R)}(x)\right)$
$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{\alpha N+k}Z_{j}>\sqrt{\alpha N}\phi_{N}^{(R)}(x)\right)\right]$
$\geq\prod_{k=\alpha N\delta_{N}}^{N}\left[1-e^{-\eta\sqrt{\alpha N}\phi_{N}(x)}E\left[e^{\eta Z_{1}}\right]^{\alpha N+k}\right]$
Furthermore,
$e^{-\eta\alpha N\left(\frac{k}{\alpha N}-1\right)-\alpha N\ln(1-\eta)}$ $\geq e^{(\alpha N+k)(\eta-\ln(1-\eta))}$
$=e^{-\eta\alpha N(\frac{k}{\alpha N}-1)-\left(\alpha N+k\right)\ln(1-\eta)}>1$
And with $0<\eta<1$ and $\forall Y\in\mathbb{R}$
$e^{Y-\eta\sqrt{\alpha N}\phi_{N}^{(R)}(x)}<e^{Y}\Rightarrow 1-e^{Y-\eta\sqrt{\alpha N}\phi_{N}^{(R)}(x)}>1-e^{Y}$
Now, setting $\eta=1-\frac{1}{\frac{k}{\alpha N}-1}$ which does maximise the remaining exponent (similarly presented in [22]) and applying the the quantile formula of the exponential distribution, this implies
$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
$\geq\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\alpha N\left[\eta\left(\frac{k}{\alpha N}-1\right)+\ln(1-\eta)\right]}\right)$
$\geq\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\alpha N\left[\frac{k}{\alpha N}-\ln\left(\frac{k}{\alpha N}-1\right)\right]}\right)$
$\geq\left(1-e^{-\alpha N\delta_{N}^{2}}\right)^{N}$
Finally,
$\prod_{k=\alpha N\delta_{N}}^{N}p_{k}\geq\left(1-e^{-\alpha N\delta_{N}^{2})}\right)^{N}$
The parameter $\delta_{N}$ is chosen such that $\left(1-e^{-\alpha N\delta_{N}^{2})}\right)^{N}=\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$, i.e., $\delta_{N}=\sqrt{\frac{2\log{N}}{\alpha N}}$.
The partial product $\prod_{k=1}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$ is bounded. It is established from the following that the limit of the partial product $\prod_{k=1}^{\alpha N\delta_{N}}h_{k}$ is equal to one as $N$ goes infinity.
More precisely,
$\prod_{k=1}^{\alpha N\delta_{N}}h_{k}\prod_{k=\alpha N\delta_{N}}^{N}h_{k}\leq\prod_{k=1}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]\leq\prod_{k=1}^{\alpha N\delta_{N}}h_{k}$
Lower bound of the partial product $\prod_{k=\alpha N\delta_{N}}^{N}h_{k}$.
For large $N$, as the random variable $\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)$ follows a standard Gaussian distribution, and taking into account a similar remark as in the paper of B. Rider [22] saying that the upper limit of integration can be extended to $\infty$, the probability
$P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\approx 1$
Now, considering the probability
$P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)$
$=1-P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)>-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)$
$=1-P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)<\phi_{N}^{(r)}(y)+\frac{k}{\sqrt{\alpha N}}\right)<\frac{1}{N^{2}}$
where the random variable $\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)$ follows a standard Gaussian distribution.
This implies that
$P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)>-\frac{1}{N^{2}}$
Consequently,
$1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}>1-\frac{1}{N^{2}}$
Thus,
$\prod_{k=\alpha N\delta_{N}}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$
$>\prod_{k=\alpha N\delta_{N}}^{N}\left(1-\frac{1}{N^{2}}\right)>\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$
The lower bound is $1-O\left(\frac{1}{N}\right)$ and goes to $1$ as $N$ goes to infinity.
Now,
$\log\prod_{k=1}^{\alpha N\delta_{N}}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
Furthermore, using the classical Edgeworth expansion
$\log P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
$=\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N-k}}\sup_{|c|\leq Y}(\phi_{N}^{(R)}(c))^{2}e^{-\frac{(\phi_{N}^{(R)}(c))^{2}}{2}}\right)$
$+O\left(\frac{1}{\alpha N-k}\right)+O\left(\phi_{N}^{(r)}(y)(\alpha N-k)^{-3/2}\right)+O\left(e^{-\frac{(\phi_{N}^{(r)}(y))^{2}}{2}}\right)$
where, as in [22], for $x$ restricted as in $|x|\leq Y$ for some large positive $Y$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq Y}\phi_{N}(x)$.
The leading order sum as the Riemann integral
$\sum_{k=1}^{\sqrt{2N\log N}}\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$ $\simeq\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\log\left[\int_{-\infty}^{u}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du$
$+\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{+\infty}\log\left[\int_{-u}^{+\infty}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du+E_{N}$
where
$E_{N}=O\left(\left(\sqrt{\frac{\log N}{\alpha N}}\right)\vee\left(\sqrt{\log{N}}\sup_{|c|\leq Y}(\phi_{N}^{(R)}(c))^{2}e^{-\frac{(\phi_{N}^{(R)}(c))^{2}}{2}}\right)\right)$
Following the framework presented in [22] and detailed in Appendices A.1 and A.2,
$\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\log\left[\int_{-\infty}^{u}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du=\log F_{X}^{(\infty)}(x)=-\exp(-x)$
and
$\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{+\infty}\log\left[\int_{-u}^{+\infty}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du=\log\left(1-F_{Y}^{(\infty)}(y)\right)=-\exp(y)$
which implies
$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]$
$=F_{X}^{(\infty)}(x)\left(1-F_{Y}^{(\infty)}(y)\right)$
with $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$.
These limiting probability distributions are the standard Gumbel distribution for maxima and the standard Gumbel distribution for minima with cumulative density function $F_{X}(x)=\exp(-\exp(-x))$ and $F_{Y}(y)=1-\exp(-\exp(y))$, respectively.
This proves the independence of the scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$ for $N\times N$ matrices from the complex induced Ginibre ensemble with proportional rectangularity index as $N$ goes to infinity.
Appendix B Conditional probability for the complex Ginibre ensemble
Let $\mathcal{D}(z_{0},s)$ denote the disk of radius $s$ centred at $z_{0}$ and $H^{(N)}(s,z_{0})$ the conditional probability that given one eigenvalue lies at the point $z_{0}$, the others are found outside $\mathcal{D}(z_{0},s)$, i.e.,
$H^{(N)}(s,z_{0})$ $=P(\left\{z_{2},\cdots,z_{N}\right\}\notin\mathcal{D}(z_{0},s)|z_{1}=z_{0})$
$=\frac{P(\left\{z_{2},\cdots,z_{N}\right\}\notin\mathcal{D}(z_{0},s)\cap z_{1}=z_{0})}{P(z_{1}=z_{0})}$
$=\frac{1}{P(z_{1}=z_{0})}\int_{\mathbb{C}}d^{2}z_{2}\cdots\int_{\mathbb{C}}d^{2}z_{N}P(z_{0},z_{2},\cdots,z_{N})\prod_{k=2}^{N}[1_{\left\{z_{k}\notin\mathcal{D}(z_{0},s)\right\}}]$
The probability $P(z_{1}=z_{0})$ is the probability density at the point $z_{0}$. It is equal to the one-point correlation function in the vicinity of the point $z_{0}$ (i.e., the density in the vicinity of the point $z_{0}$) divided by the total number of eigenvalues which is equal to $N$.
More precisely,
$P(z_{1}=z_{0})=\frac{R_{1}^{(N)}(z_{0})}{N}$
This implies that
$H^{(N)}(s,z_{0})=\frac{N}{R_{1}^{(N)}(z_{0})}\int_{\mathbb{C}}d^{2}z_{2}\cdots\int_{\mathbb{C}}d^{2}z_{N}P(z_{0},z_{2},\cdots,z_{N})\prod_{k=2}^{N}[1_{\left\{z_{k}\notin\mathcal{D}(z_{0},s)\right\}}]$
The joint probability density function given that one eigenvalue is at origin $z_{0}=0$ is
$P(0,z_{2},\cdots,z_{N})=\frac{1}{N!\pi^{N}\prod_{j=0}^{N-1}j!}e^{-\sum_{j=2}^{N}|z_{j}|^{2}}\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}$
Let $z_{1}=z_{0}=0$, the Vandermonde determinant
$\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}=\prod_{j=2}^{N}|z_{j}|^{2}\prod_{2\leq i<j\leq N}|z_{i}-z_{j}|^{2}$
The term
$\prod_{2\leq i<j\leq N}|z_{i}-z_{j}|^{2}$ $=|\Delta(z_{2},\cdots,z_{N})|^{2}$
$=\Delta(z_{2},\cdots,z_{N})\overline{\Delta}(z_{2},\cdots,z_{N})=\det\left(z_{k}^{N-j}\right)_{k,j=2}^{N}\det\left(\bar{z}_{k}^{N-j}\right)_{k,j=2}^{N}$
which implies
$\prod_{j=2}^{N}|z_{j}|^{2}\prod_{2\leq i<j\leq N}|z_{i}-z_{j}|^{2}$ $=\prod_{j=2}^{N}|z_{j}|^{2}\det\left(z_{k}^{N-j}\right)_{k,j=2}^{N}\det\left(\bar{z}_{k}^{N-j}\right)_{k,j=2}^{N}$
$=\prod_{j=2}^{N}z_{j}\det\left(z_{k}^{N-j}\right)_{k,j=2}^{N}\prod_{j=2}^{N}\bar{z}_{j}\det\left(\bar{z}_{k}^{N-j}\right)_{k,j=2}^{N}$
$=\det\left(z_{k}^{N+1-j}\right)_{k,j=2}^{N}\det\left(\bar{z}_{k}^{N+1-j}\right)_{k,j=2}^{N}$
Finally,
$P(0,z_{2},\cdots,z_{N})$ $=\frac{1}{N!\pi^{N}\prod_{j=0}^{N-1}j!}e^{-\sum_{j=2}^{N}|z_{j}|^{2}}\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}$
$=\frac{1}{N!\pi^{N}\prod_{j=0}^{N-1}j!}e^{-\sum_{j=2}^{N}|z_{j}|^{2}}\det\left(z_{k}^{N+1-j}\right)_{k,j=2}^{N}\det\left(\bar{z}_{k}^{N+1-j}\right)_{k,j=2}^{N}$
Furthermore,
$R_{1}^{(N)}(0)=\frac{1}{\pi}\frac{\Gamma(N,0)}{\Gamma(N)}=\frac{1}{\pi}$
Applying the Andreief’s integration formula
$H^{(N)}(s,0)$ $=\frac{N(N-1)!}{N!\pi^{N-1}\prod_{j=0}^{N-1}j!}\det\left[\int_{|z|>s}\left(|z|^{2}\right)^{N+1-j}e^{-|z|^{2}}d^{2}z\right]_{j=2}^{N}$
$=\frac{1}{\prod_{j=0}^{N-1}j!}\prod_{k=1}^{N-1}\int_{s^{2}}^{+\infty}e^{-t}t^{k}dt=\prod_{k=1}^{N-1}\frac{\Gamma(k+1,s^{2})}{\Gamma(k+1)}$
Also,
$e^{x}=\sum_{k=0}^{+\infty}\frac{x^{k}}{k!}=\sum_{k=0}^{n}\frac{x^{k}}{k!}+\sum_{k=n+1}^{+\infty}\frac{x^{k}}{k!}$
and then,
$e^{-x}\sum_{k=0}^{n}\frac{x^{k}}{k!}=1-e^{-x}\sum_{k=n+1}^{+\infty}\frac{x^{k}}{k!}$
which implies that
$H^{(N)}(s,0)=\prod_{k=1}^{N-1}\frac{\Gamma(k+1,s^{2})}{\Gamma(k+1)}=\prod_{k=1}^{N-1}e^{-s^{2}}\sum_{n=0}^{k}\frac{s^{2n}}{n!}=\prod_{k=1}^{N-1}\left[1-e^{-s^{2}}\sum_{n=k+1}^{+\infty}\frac{s^{2n}}{n!}\right]$
as stated in reference [17].

### v3

@@
 The limiting distribution of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble, with a proportional rectangularity index, is the Gumbel distribution.
-The independence of these extrema is established, at appropriate scaling, for large matrices from the complex Ginibre ensemble as well as from the complex induced Ginibre ensemble for either fixed or proportional rectangularity indexes.
-The joint distribution of the eigenvalue moduli for the complex Ginibre ensemble is derived in this paper.
-It is equal to the joint distribution of independent random variables, each following a Gamma-Rayleigh distribution.
+The independence of these extrema is established, at appropriate scaling, for large matrices from the complex Ginibre ensemble as well as from the complex induced Ginibre ensemble for fixed and proportional rectangularity indexes.
 In the limit of a large size of the complex Ginibre matrices, the left and right tail distributions of the minimum modulus are the Rayleigh and Weibull distributions, respectively.
-The limiting left tail distribution of minimum modulus is the same for these non-Hermitian ensembles when the rectangularity index of the complex induced Ginibre ensemble is equal to zero.
+The limiting left tail distribution of the minimum modulus is the same for these non-Hermitian ensembles when the rectangularity index of the complex induced Ginibre ensemble is equal to zero.
 This phenomenon is also verified for the right tail distribution of this minimum.
@@
 The complex Ginibre ensemble is the space of $N\times N$ complex matrices $J$ whose complex entries are the $J_{ij}=x_{ij}+\text{i}y_{ij}\in\mathbb{C}$ independent and identically distributed (i.i.d.) following a standard complex Gaussian distribution $\mathcal{N}_{\mathbb{C}}(0,1)$ with probability density
-$P(J_{ij})=\frac{1}{\pi}e^{-|J_{ij}|^{2}}$
+$P(J_{ij})=\frac{1}{\pi}e^{-|J_{ij}|^{2}}$ (1)
 where $(x_{ij})_{1\leq i,j\leq N}\in\mathbb{R}^{N\times N}$ and $(y_{ij})_{1\leq i,j\leq N}\in\mathbb{R}^{N\times N}$.
@@
 The space of $N\times N$ complex matrices from the complex Ginibre ensemble is endowed with a probability measure here denoted $\mu(J)$, where $d\mu(J)=P(J)|D(J)|$ and $P(J)$ is the joint probability density function of the entries $J_{ij}$ defined as
-$P(J)=\frac{1}{\pi^{N^{2}}}\exp\left(-\operatorname{Tr}\left(JJ^{*}\right)\right)$
+$P(J)=\frac{1}{\pi^{N^{2}}}\exp\left(-\operatorname{Tr}\left(JJ^{*}\right)\right)$ (2)
 and $|D(J)|=\frac{1}{2^{N^{2}}}\bigwedge_{k=1}^{N}\bigwedge_{j=1}^{N}|dJ_{kj}\wedge dJ_{kj}^{*}|$ is called the cartesian volume element as presented in [15].
 The joint probability density function of the eigenvalues of complex Ginibre matrices is expressed in close form as
-$P_{N}(z)=\frac{1}{\pi^{N}C_{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)|\Delta(z)|^{2}$
+$P_{N}(z)=\frac{1}{\pi^{N}C_{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)|\Delta(z)|^{2}$ (3)
 where $z=(z_{1},\dots,z_{N})\in\mathbb{C}^{N}$ is the $N$-tuple of eigenvalues of the $N\times N$ complex matrix $J$ and $\Delta(z)$ is the Vandermonde determinant $\Delta(z)=\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|$, with $|\Delta(z)|^{2}=\Delta(z)\overline{\Delta(z)}$.
 The term $C_{N}$ is the normalisation factor defined as
-$C_{N}=\frac{1}{\pi^{N}}\int_{\mathbb{C}^{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}\prod_{k=1}^{N}d^{2}z_{k}=\prod_{k=1}^{N}\Gamma(k+1)$
+$C_{N}=\frac{1}{\pi^{N}}\int_{\mathbb{C}^{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}\prod_{k=1}^{N}d^{2}z_{k}=\prod_{k=1}^{N}\Gamma(k+1)$ (4)
 A topic which has been widely investigated in RMT is the limiting distribution of the eigenvalues of random matrix ensembles as $N$ goes to infinity.
@@
 Pursuing the analysis in the scaling limit $\sqrt{N}$ for this random matrix ensemble as $N$ goes to infinity, the independence of the scaled minimum moduli with respect to the scaled spectral radius is established in Section 2.2.
-The eigenvalues of matrices from the complex Ginibre ensemble form a determinantal point process and the joint probability density of their respective radius has been investigated in the work of E.
-Kostlan [18] extended in [16].
-This result is also discussed in the present paper in Section 2.2.
+The eigenvalues of matrices from the complex Ginibre ensemble form a determinantal point process and the joint probability density of their respective radius has been investigated in [16], and is extended in [14].
 Employing the method introduced in [20] from which limit theorems are derived at the edge of the spectrum for the complex and symplectic Ginibre ensembles, results establishing the limiting distribution of the scaled spectral radius and minimum modulus for the complex induced Ginibre ensemble [5] are presented in Section 3.
@@
 The joint probability density function of the entries of matrices $A$ from the Feinberg-Zee ensemble is
-$P_{FZ}(A)\propto\exp\left(-\operatorname{Tr}V(A^{\dagger}A)\right)$
+$P_{FZ}(A)\propto\exp\left(-\operatorname{Tr}V(A^{\dagger}A)\right)$ (5)
 The induced Ginibre ensembles correspond to the Feinberg-Zee ensemble with potential $V(y)=-\frac{\beta}{2}(y-L\log(y))$ where $L$ is a non-negative parameter called the rectangularity index.
 Let $G$ denote a $N\times N$ complex induced Ginibre matrix. The joint probability density function of its entries is
-$P_{indG}^{(\beta)}(G)=\Gamma^{(\beta)}\left(\det G^{\dagger}G\right)^{\frac{\beta}{2}L}\exp\left(-\operatorname{Tr}G^{\dagger}G\right)\\$
+$P_{indG}^{(\beta)}(G)=\Gamma^{(\beta)}\left(\det G^{\dagger}G\right)^{\frac{\beta}{2}L}\exp\left(-\operatorname{Tr}G^{\dagger}G\right)\\$ (6)
 where $\beta=2$ for the complex induced Ginibre ensemble and the determinant $\left(\det G^{\dagger}G\right)^{a}=\prod_{k=1}^{N}|\lambda_{i}|^{2a}$, $a\geq 0$.
 The factor $\Gamma^{(\beta)}$ is the normalisation constant
-$\Gamma^{(\beta)}=\pi^{-\frac{\beta}{2}N^{2}}\left(\frac{\beta}{2}\right)^{\frac{1}{2}N(N-L)}\prod_{j=1}^{N}\frac{\Gamma(\frac{\beta}{2}j)}{\Gamma(\frac{\beta}{2}(j+L)}$
+$\Gamma^{(\beta)}=\pi^{-\frac{\beta}{2}N^{2}}\left(\frac{\beta}{2}\right)^{\frac{1}{2}N(N-L)}\prod_{j=1}^{N}\frac{\Gamma(\frac{\beta}{2}j)}{\Gamma(\frac{\beta}{2}(j+L)}$ (7)
 In the limit of large matrices dimensions, the eigenvalues are spread across an annulus in the complex plane which is distinguished from the phenomenon identified for complex Ginibre ensemble thus referenced as the circular law and presented in the work from V.
@@
 The joint probability density function of the eigenvalues of complex induced Ginibre matrices is
-$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{\Gamma_{N}}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\\$
+$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{\Gamma_{N}}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\\$ (8)
 with normalisation constant
-$\Gamma_{N}$ $=\int\cdots\int\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\prod_{j=1}^{N}d^{2}\lambda_{j}$
-$=N!\pi^{N}\prod_{k=1}^{N}\Gamma(k+L)$
+$\Gamma_{N}=\int\cdots\int\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\prod_{j=1}^{N}d^{2}\lambda_{j}=N!\pi^{N}\prod_{k=1}^{N}\Gamma(k+L)$ (9)
 It is understood from the definition of the joint probability density function of the eigenvalues that there is a repulsion of the eigenvalues from the origin of the complex plane due to the term $\det(GG^{\dagger})^{L}=\prod_{j=1}^{N}|\lambda|^{2L}$ appearing in the formula.
@@
 where each random variables $\gamma_{k}$ are independent and follows a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ , with $\alpha_{k}=k$ and $\delta_{k}=1$.
-The Gamma-Rayleigh distribution is derived in reference [1].
-Corollary 3.
-Let $\left((r_{i}^{(N)}(A))^{2}\right)_{i=1}^{N}$ denote the squared (unordered) moduli of the $N$ eigenvalues $\left(z_{i}^{(N)}(A)\right)_{i=1}^{N}$ of a $N\times N$ complex Ginibre matrix $A$.
-The distribution of the squared spectral radius $\left(r_{max}^{(N)}(A)\right)^{2}$ is the joint distribution of the elements of a finite sequence $\{\zeta_{k}\}_{k=1}^{N}$.
-The $\zeta_{k}$ are independent random variables.
-Each $\zeta_{k}$ follows a Gamma distribution with shape parameter $\alpha_{k}=k$ and scale parameter $\delta_{k}=1$, i.e., $\zeta_{k}\sim Gamma(k,1)$.
-Proof.
-A simple proof of this result is as follows.
-Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble.
-Let $\left(r_{i}^{(N)}(A)\right)_{1\leq i\leq N}$ denote the moduli of the eigenvalues of the matrix $A$.
-The joint distribution of the square moduli $\left((r_{i}^{(N)}(A))^{2}\right)_{i=1}^{N}$ is derived from Equation (4) and is expressed as
-$P\left((r_{max}^{(N)}(A))^{2}\leq r\right)=\prod_{k=1}^{N}\frac{\gamma(k,r)}{\Gamma(k)}=\prod_{k=1}^{N}P(\zeta_{k}\leq r)\\$
-where $P(\zeta_{k}\leq r)=\frac{\gamma(k,r)}{\Gamma(k)}$ is the cumulative distribution function of the random variable $\zeta_{k}$ following a Gamma distribution with shape parameter $\alpha_{k}=k$ and rate parameter $\delta_{k}=1$.
+The Gamma-Rayleigh distribution is derived in the work of E.
+Akarawak, I.
+Adeleke and R.
+Okafor [1].
 The corresponding probability density function of the largest modulus $r_{max}^{(N)}(A)$ is derived from Equation (13).
@@
 Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble.
-The left tail distribution of minimum modulus $r_{min}^{(N)}(A)$ (i.e., the smallest values of $r_{min}^{(N)}(A)$) converges to the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$ as $N$ goes to infinity.
-More precisely,
-$\lim_{r\longrightarrow 0}\lim_{N\longrightarrow+\infty}P(r_{min}^{(N)}(A)<r)=1-e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (9)
+The left tail distribution of minimum modulus $r_{min}^{(N)}(A)$ (i.e., the smallest values of $r_{min}^{(N)}(A)$) converges to the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$ as $N$ goes to infinity
+More precisely, for $0<r\ll 1$,
+$\lim_{N\longrightarrow+\infty}P(r_{min}^{(N)}(A)<r)=1-e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (18)
 Proof.
@@
 $\lim_{N\rightarrow+\infty}H^{(N)}(r,0)$ $=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{4(k+1)}\right)\right)\right]^{j}$
-Consequently, as $r$ goes to zero,
+Consequently, for $0<r\ll 1$,
 $\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)<r\right)=1-e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$
@@
 with speed $c_{N}=\frac{N^{2}}{4}$ and rate function $I(z)=|z|^{4}$.
-2.2. The joint distribution of the eigenvalue moduli for the complex Ginibre ensemble
-G.
-Dubach proved in [6] that the collection of squared moduli of the eigenvalues of $N\times N$ matrices from the complex Ginibre ensemble has the same distribution as the collection of independent random variables $\zeta_{k}$, each following a Gamma distribution with shape parameter $\alpha_{k}=k$ and scale parameter $\delta_{k}=1$, i.e., $\zeta_{k}\sim Gamma(k,1)$.
-He did provide proof of this result which involves the use of the Andreief’s integration formula [9] with radially symmetric functions defined on the complex plane.
-Theorem 6.
-The joint distribution of the unordered eigenvalue moduli $\{r_{k}^{(N)}(A)\}_{k=1}^{N}$ of $N\times N$ matrices $A$ from the complex Ginibre ensemble is the joint distribution of independent random variables $\zeta_{k}$ each following a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$, with $\alpha_{k}=k$ and $\delta_{k}=1$,
-$\{r_{1}^{(N)}(A),...,r_{N}^{(N)}(A)\}\overset{d}{=}\{\zeta_{1},...,\zeta_{N}\}$
-Proof.
-Let $\left(\Omega,\mathcal{F},\mu\right)$ denote a measure space with $\Omega=\mathbb{C}$, $\mathcal{F}$ a $\sigma$-algebra on the complex plane and the measure $\mu$, a nonnegative function, defined on the measurable space $\left(\Omega,\mathcal{F}\right)$.
-Let $Q$ denote the set of $N$ moduli $r_{k}^{(N)}(A)$ of $N\times N$ matrices $A$ from the complex Ginibre ensemble, i.e., $Q=\left(q_{k}\right)_{k=1}^{N}=\left(r_{k}^{(N)}(A)\right)_{k=1}^{N}$.
-Let $\varphi$ denote a (suitable) test function with $\mu$ the reference measure on the complex plane.
-Using the Andreief’s integration formula
-$\mathbb{E}\left[\prod_{q_{k}\in Q}\varphi(q_{k})\right]$ $=\int...\int\prod_{k=1}^{N}\varphi(|z_{k}|)P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d\mu(z_{k})$
-$=\prod_{k=1}^{N}\int_{0}^{+\infty}\varphi(r)\frac{2r^{2k-1}e^{-r^{2}}}{\Gamma(k)}dr=\prod_{k=1}^{N}\int_{0}^{+\infty}\varphi(r)p_{GR(k,1)}(r)dr=\prod_{k=1}^{N}\mathbb{E}\left[\varphi(\zeta_{k})\right]$
-where each $\zeta_{k}$ follows a Gamma-Rayleigh distribution with shape parameter $\alpha_{k}=k$ and scale parameter $\delta_{k}=1$.
-The Gamma-Rayleigh distribution is derived in the work of E.
-Akarawak, I.
-Adeleke and R.
-Okafor [1].
-Remark 2.2.
-[…]
-$\left\{|z_{1}|,...,|z_{n}|\right\}\overset{d}{=}\left\{R_{(1)},...,R_{(n)}\right\}$
-where $|z_{1}|\geq...\geq|z_{n}|$ are the order statistics of the moduli and where $R_{(1)}\geq...\geq R_{(n)}$ is the order statistics of independent random variables $R_{1},...,R_{n}$ with $R_{k}$ of density proportional to $t\longmapsto t^{2k-1}e^{-nV(t)}1_{[0,+\infty)}(t)$ with $V(t)=t^{2}$, for every $1\leq k\leq n$.
-Remark 2.3.
-The joint distribution of the moduli of $N\times N$ matrices for the complex Ginibre ensemble has been investigated in the work of E.
-Kostlan [18].
-He stated that the collection of moduli of the eigenvalues of $N\times N$ matrices from the complex Ginibre ensemble has the same distribution as the collection of random variables $\{X_{2k}\}_{1\leq k\leq N}$ where, $\forall k\in\{1,...,N\}$, $X_{2k}$ follows a chi distribution $\chi_{2k}$.
-A slight extension of this result would correspond to state the next Theorem 7.
-Theorem 7.
-The collection of unordered eigenvalue moduli $|z_{k}|$ of $N\times N$ matrices from the complex Ginibre ensemble has the same distribution as the collection of random variables $\{\frac{X_{2k}}{\sqrt{2}}\}_{1\leq k\leq N}$ where, $\forall k\in\{1,...,N\}$, $X_{2k}$ follows a chi distribution $\chi_{2k}$.
-Proof.
-Let $Q$ denote the set of $N$ moduli $|z_{k}|$ of matrices from the complex Ginibre ensemble, i.e., $Q=\left(q_{k})\right)_{k=1}^{N}$.
-Let $f$ denote a (suitable) test function.
-Using the Andreief’s integration formula,
-$\mathbb{E}\left[\prod_{q_{k}\in Q}f(q_{k})\right]$ $=\int...\int\prod_{k=1}^{N}f(|z_{k}|)P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d\mu(z_{k})$
-$=\prod_{k=1}^{N}\int_{0}^{+\infty}f(v/\sqrt{2})\frac{v^{2k-1}e^{-v^{2}/2}}{2^{k-1}\Gamma(k)}dv=\prod_{k=1}^{N}\int_{0}^{+\infty}f(v/\sqrt{2})p_{\chi_{2k}}(v)dv=\prod_{k=1}^{N}\mathbb{E}\left[f\left(\frac{\alpha_{k}}{\sqrt{2}}\right)\right]$
-Each random variable $\alpha_{k}$ follows a chi distribution with parameter $2k$, i.e., $\alpha_{k}\sim\chi_{2k}$ for $k\in\{1,...,N\}$.
 2.2. Independence of the scaled spectral radius and the scaled minimum modulus for the complex Ginibre ensemble
@@
 Using the framework presented in [20], it is known that
-$\lim_{N\rightarrow+\infty}P\left(R_{N}\leq R^{(N,x)}\right)=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\gamma(k+1,N\left(R^{(N,x)}\right)^{2})}{\Gamma(k+1)}=F_{X}(x)=e^{-e^{-x}}$
+$\lim_{N\rightarrow+\infty}P\left(R_{N}\leq R^{(N,x)}\right)=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\gamma\left(k+1,N\left(R^{(N,x)}\right)^{2}\right)}{\Gamma(k+1)}=F_{X}(x)=e^{-e^{-x}}$
 where $F_{X}(x)$ is the limiting cumulative distribution function of the scaled spectral radius $R_{N}$ defined in [20].
@@
 where $Nr^{2}$ is kept fixed while $N$ goes to infinity), as $N$ goes infinity (or $N$ very large) also known as the origin limit
-$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=E_{0}^{(2)}\left(\sqrt{N}r\right)$
+$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}E_{0}^{(2)}\left(\sqrt{N}r\right)$
 Defined in [2], $E_{0}^{(\beta)}\left(\sqrt{N}r\right)$ is the probability that zero eigenvalue lies inside the disk of radius $\sqrt{N}r$ centred at the origin of the complex plane, and all eigenvalues lie outside.
@@
 Let $A$ denote a $N\times N$ random matrix from the complex Ginibre ensemble.
-The limiting left tail probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is the probability density function of the Rayleigh distribution with parameter $\frac{1}{\sqrt{2}}$, in the neighbourhood of zero, as $N$ goes to infinity.
-More precisely,
-$\lim_{r\longrightarrow 0}\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (25)
+The limiting left tail probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is the probability density function of the Rayleigh distribution with parameter $\frac{1}{\sqrt{2}}$, as $r$ goes to zero and $N$ goes to infinity.
+More precisely, for $0<r\ll 1$,
+$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (34)
 Proof.
-The limit of partial product $\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$, as $N$ goes to infinity, has been investigated in Section 2 for the complex Ginibre ensemble, and for small $r$ in the neighbourhood of zero, it is expressed as
+The limit of partial product $\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$, as $N$ goes to infinity, has been investigated in Section 2 for the complex Ginibre ensemble, and for $0<r\ll 1$, it is expressed as
 $\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$
 The upper incomplete Gamma function $\Gamma(k,x)$, for small $x$ and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
-The limit of the partial sum is derived as follows. Knowing that, for $r$ in the neighbourhood of zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$,
+The limit of the partial sum is derived as follows. Knowing that, for $0<r\ll 1$,
+$\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$
 $\lim_{N\rightarrow+\infty}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$ $=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{r^{2j}}{\Gamma(j+1,r^{2})}=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{(r^{2})^{j}}{j!}=2e^{-r^{2}}re^{r^{2}}=2r$
-Thus, the limiting left tail probability density function of the minimum of moduli for matrices from the complex Ginibre ensemble is
+Thus, the limiting left tail probability density function of the minimum modulus for matrices from the complex Ginibre ensemble is
 $\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)$ $=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$
@@
 The minimum modulus of a $N\times N$ matrix $G$ from the complex induced Ginibre ensemble, with rectangularity index $L=1$, has a left tail probability density function of the Generalised Gamma distribution with scale parameter $\alpha=1$ and other parameters $\kappa=2$ and $\gamma=4$.
-The corresponding limiting probability density function, for $r$ in the neighbourhood of zero, as $N$ goes to infinity, is
-$\lim_{r\longrightarrow 0}\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2r^{3}e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (26)
+The corresponding limiting probability density function, for $0<r\ll 1$, as $N$ goes to infinity, is
+$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2r^{3}e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (35)
 Proof.
@@
 This is established with the results presented in Section 2. The partial product is extended from N1N-1 to $N$ which does not change the limit.
-Also, the upper incomplete Gamma function $\Gamma(k,x)$, for $x$ approaches zero and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
-For $r$ close to zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$. This implies that the limit of the partial sum, as $r$ approaches zero and as $N$ goes to infinity, is
+Also, the upper incomplete Gamma function $\Gamma(k,x)$, as $x$ approaches zero and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
+For $r$ close to zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$. This implies that the limit of the partial sum, as $r$ approaches zero and $N$ goes to infinity, is
 $\lim_{N\rightarrow+\infty}2re^{-r^{2}}\sum_{j=1}^{N}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]\\ =2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{\Gamma(j+1,r^{2})}\times\left(1-O\left(r^{4}\right)\right)=2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{j!}\times\left(1-O\left(r^{4}\right)\right)$
@@
 Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L\geq 0$.
-The left tail distribution of minimum of the moduli $r_{min}^{(N)}(G)$ is the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$ in the limit as $N$ goes to infinity and $r$ approaches zero zero,
+The left tail distribution of minimum of the moduli $r_{min}^{(N)}(G)$ is the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$ in the limit as $N$ goes to infinity and $r$ approaches zero,
 $\lim_{N\longrightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-\exp\left[-\frac{r^{2(L+1)}}{(L+1)!}\right]+O\left(r^{4(L+1)}\right)$ (36)
 Proof.
-This result is derived using the same methodological approach presented in section 2.1. For positive values of $r$ in the neighbourhood of zero,
+This result is derived using the same methodological approach presented in section 2.1.
+For $0<r\ll 1$,
 $\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)$ $=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+L)}}{(k+L)!}+O\left(r^{4(k+L)}\right)\right)\right]^{j}$
@@
 Corollary 14.
-Setting the rectangularity index $L$ to zero, the limiting distribution of the minimum modulus $r_{min}^{(N)}(G)$ for the complex induced Ginibre ensemble corresponds to the limiting distribution of the minimum modulus $r_{min}^{(N)}(A)$ for the complex Ginibre ensemble, i.e., for small $r$
+Setting the rectangularity index $L$ to zero, the limiting distribution of the minimum modulus $r_{min}^{(N)}(G)$ for the complex induced Ginibre ensemble corresponds to the limiting distribution of the minimum modulus $r_{min}^{(N)}(A)$ for the complex Ginibre ensemble, i.e., for $0<r\ll 1$,
 $\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-r^{2}}+O(r^{4})=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)<r\right)$
 Corollary 15.
-For positive values of $r$ in the neighbourhood of zero, as $N$ goes to infinity, and for a rectangularity index $L=0$,
-$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)+O\left(r^{5}\right)$
+For $0<r\ll 1$, as $N$ goes to infinity, and for a rectangularity index $L=0$,
+$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)-O\left(r^{5}\right)$
 with shape parameter $\kappa=2$ and scale parameter $\lambda=1$. This does correspond to the probability density function of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
 Proof.
+For $0<r\ll 1$,
 $\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\frac{\kappa}{\lambda}\left(\frac{r}{\lambda}\right)^{\kappa-1}e^{-(r/\lambda)^{\kappa}}+O\left(r^{3\kappa+1}\right)=2re^{-r^{2}}+O\left(r^{7}\right)$
@@
 A.2. Proof of Theorem 7
-It is demonstrated in Section 2 that the scaled minimum radius $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ of eigenvalues for matrices from the complex Ginibre ensemble, in the neighbourhood of zero, follows a Rayleigh distribution.
+It is demonstrated in Section 2 that the scaled minimum radius $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ of eigenvalues for matrices from the complex Ginibre ensemble, close to zero, follows a Rayleigh distribution.
 The complex Ginibre ensemble does correspond to the complex induced Ginibre ensemble when the rectangularity index $L$ is equal to zero.

### v4

@@
 Abstract.
-The distribution of the modulus of the extreme eigenvalues is investigated for the complex Ginibre and complex induced Ginibre ensembles in the limit of large dimensions of random matrices.
+The complex Ginibre ensemble and its generalisation, the complex induced Ginibre ensemble, have been well studied in the field of Random Matrix Theory.
+However, not all questions have formal responses.
+In the present work, asymptotic distributions and the independence of the modulus of the extreme eigenvalues are studied in the limit of large dimensions of random matrices for these non-Hermitian ensembles.
+They are derived with the use of Andreief’s integration formula and the known methodological approach defined for the study of the limiting distribution of the scaled spectral radius at the edge of the complex Ginibre ensemble.
 The limiting distribution of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble, with a proportional rectangularity index, is the Gumbel distribution.
-The independence of these extrema is established, at appropriate scaling, for large matrices from the complex Ginibre ensemble as well as from the complex induced Ginibre ensemble for fixed and proportional rectangularity indexes.
 In the limit of a large size of the complex Ginibre matrices, the left and right tail distributions of the minimum modulus are the Rayleigh and Weibull distributions, respectively.
-The limiting left tail distribution of the minimum modulus is the same for these non-Hermitian ensembles when the rectangularity index of the complex induced Ginibre ensemble is equal to zero.
-This phenomenon is also verified for the right tail distribution of this minimum.
+The limiting left tail distribution of the minimum modulus is the same for these matrix ensembles, with a rectangularity index of the complex induced Ginibre ensemble equal to zero.
+This phenomenon is also verified for the right tail distribution of this random minimum.
+The independence of the extreme moduli is formally established, at appropriate scaling, for large matrices from the complex Ginibre ensemble as well as from the complex induced Ginibre ensemble for fixed and proportional rectangularity indexes.
+This study extends knowledge in the field of Random Matrix Theory for random variables, like the minimum modulus of matrices from non-Hermitian ensembles, whose limiting stochastic dynamic has never been explored.
 OLIVIA V. AUSTER
@@
 The complex Ginibre ensemble is one of them and is a special case of the Ginibre-Girko ensemble with maximal non-Hermiticity as presented in [9] and [2].
-The complex Ginibre ensemble was first defined as a mathematical concept to model phenomena from the physics of particles.
+Like the real and the symplectic Ginibre ensembles, this ensemble was introduced by J.
+Ginibre [11].
+It was first defined as a mathematical concept to model phenomena from the physics of particles.
 Statistical properties of eigenvalues of matrices from this ensemble are studied to provide an understanding by analogy of the dynamics of nuclei.
-More precisely, as stated in [10], the distribution of eigenvalues of a complex Ginibre matrix is comparable to that of the distribution of the positions of charges of a two-dimensional Coulomb gas in a harmonic oscillator potential, at a specific temperature corresponding to the Dyson index $\beta=2$.
-The complex Ginibre ensemble is the space of $N\times N$ complex matrices $J$ whose complex entries are the $J_{ij}=x_{ij}+\text{i}y_{ij}\in\mathbb{C}$ independent and identically distributed (i.i.d.) following a standard complex Gaussian distribution $\mathcal{N}_{\mathbb{C}}(0,1)$ with probability density
+The distribution of eigenvalues of a complex Ginibre matrix is comparable to that of the distribution of the position of charges of a two-dimensional Coulomb gas in a harmonic oscillator potential, at a specific temperature corresponding to the Dyson index $\beta=2$.
+The complex Ginibre ensemble is the space of $N\times N$ complex matrices $J$, whose complex entries are the $J_{ij}=x_{ij}+\text{i}y_{ij}\in\mathbb{C}$ independent and identically distributed (i.i.d.) following a standard complex Gaussian distribution $\mathcal{N}_{\mathbb{C}}(0,1)$ with probability density function
 $P(J_{ij})=\frac{1}{\pi}e^{-|J_{ij}|^{2}}$ (1)
@@
 The real and imaginary parts of the entries denoted $x_{ij}$ and $y_{ij}$ are independent random variables, each following a real Gaussian distribution $\mathcal{N}_{\mathbb{R}}(0,\frac{1}{2})$.
-Like the real and the symplectic Ginibre ensembles, this ensemble was introduced by J.
-Ginibre [10].
-The space of $N\times N$ complex matrices from the complex Ginibre ensemble is endowed with a probability measure here denoted $\mu(J)$, where $d\mu(J)=P(J)|D(J)|$ and $P(J)$ is the joint probability density function of the entries $J_{ij}$ defined as
+The space of $N\times N$ complex matrices from the complex Ginibre ensemble is endowed with a probability measure here denoted $\mu(J)$, where $d\mu(J)=P(J)|D(J)|$ and $P(J)$ the joint probability density function of the entries $J_{ij}$ is
 $P(J)=\frac{1}{\pi^{N^{2}}}\exp\left(-\operatorname{Tr}\left(JJ^{*}\right)\right)$ (2)
-and $|D(J)|=\frac{1}{2^{N^{2}}}\bigwedge_{k=1}^{N}\bigwedge_{j=1}^{N}|dJ_{kj}\wedge dJ_{kj}^{*}|$ is called the cartesian volume element as presented in [15].
-The joint probability density function of the eigenvalues of complex Ginibre matrices is expressed in close form as
+The term $|D(J)|=\frac{1}{2^{N^{2}}}\bigwedge_{k=1}^{N}\bigwedge_{j=1}^{N}|dJ_{kj}\wedge dJ_{kj}^{*}|$ is called the cartesian volume element as presented in [17].
+The joint probability density function of the eigenvalues of complex Ginibre matrices is expressed in closed form as
 $P_{N}(z)=\frac{1}{\pi^{N}C_{N}}\exp\left(-\sum_{k=1}^{N}|z_{k}|^{2}\right)|\Delta(z)|^{2}$ (3)
-where $z=(z_{1},\dots,z_{N})\in\mathbb{C}^{N}$ is the $N$-tuple of eigenvalues of the $N\times N$ complex matrix $J$ and $\Delta(z)$ is the Vandermonde determinant $\Delta(z)=\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|$, with $|\Delta(z)|^{2}=\Delta(z)\overline{\Delta(z)}$.
+where $z=(z_{1},\dots,z_{N})\in\mathbb{C}^{N}$ is the $N$-tuple of the eigenvalues of the $N\times N$ complex matrix $J$ and $\Delta(z)$ is the Vandermonde determinant $\Delta(z)=\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|$, with $|\Delta(z)|^{2}=\Delta(z)\overline{\Delta(z)}$.
 The term $C_{N}$ is the normalisation factor defined as
@@
 It is a mathematical concept asserting that the limiting distribution of the eigenvalues should not depend on the particular distribution of the random matrix entries [25].
-This has led to several important results among which is referenced the circular law conjecture for non-Hermitian random matrix ensembles (cf.
-[11], [23], [3]).
+This has led to several important results among which is referenced the circular law conjecture for non-Hermitian ensembles (cf.
+[3, 12, 25]).
 Studies of the limiting distribution of the largest and smallest eigenvalues for Gaussian ensembles are presented in [27, 28, 29].
-The distributions of eigenvalues (and their spacings) have been investigated with the computation of gap probabilities with respect to radial ordering for non-Hermitian random matrices and their chiral counterparts [2].
-The statistical properties of extreme eigenvalue moduli of matrices from non-Hermitian ensembles have also been studied in the literature such as the limiting distribution of the spectral radius for the complex Ginibre ensemble [20, 21] and the real Ginibre ensemble in [22].
-Section 2 is devoted to the analysis of the distribution of the eigenvalue moduli for the complex Ginibre ensemble.
-The main results are derived for the spectral radius and the minimum modulus.
-The limiting left and right tail distributions of the minimum modulus is also studied for large size of matrices in Section 2.1.
-Pursuing the analysis in the scaling limit $\sqrt{N}$ for this random matrix ensemble as $N$ goes to infinity, the independence of the scaled minimum moduli with respect to the scaled spectral radius is established in Section 2.2.
-The eigenvalues of matrices from the complex Ginibre ensemble form a determinantal point process and the joint probability density of their respective radius has been investigated in [16], and is extended in [14].
-Employing the method introduced in [20] from which limit theorems are derived at the edge of the spectrum for the complex and symplectic Ginibre ensembles, results establishing the limiting distribution of the scaled spectral radius and minimum modulus for the complex induced Ginibre ensemble [5] are presented in Section 3.
+The distributions of eigenvalues (and their spacings) have been investigated with the computation of gap probabilities with respect to radial ordering for non-Hermitian random matrices and their chiral counterparts in [2].
+The statistical properties of the spectral radius of matrices from non-Hermitian ensembles have also been studied in the literature, as it is for the complex Ginibre ensemble [22, 23] and the real Ginibre ensemble [24].
+The present paper is structured as follows.
+Section 2 detailed the methodology used.
+The latter encompasses the application of Andreief’s integration formula [8] and the framework established by B.
+Rider in [22].
+The results are presented in Section 3.
+Section 3.1 is devoted to the analysis of the distribution of the eigenvalue moduli for the complex Ginibre ensemble.
+It is especially performed for the spectral radius and the minimum modulus.
+The limiting left and right tail distributions of the minimum modulus are studied for large matrices in Section 3.1.1.
+Pursuing the analysis for this random matrix ensemble in the scaling limit $\sqrt{N}$ as $N$ goes to infinity, the independence of the scaled minimum moduli with respect to the scaled spectral radius is established in Section 3.1.2.
+The eigenvalues of matrices from the complex Ginibre ensemble form a determinantal point process.
+The joint probability density of their respective modulus has been investigated in [18].
+This result is extended in [16].
+Employing the method introduced in [22] from which limit theorems are derived at the edge of the spectrum of the complex and symplectic Ginibre ensembles, results establishing the limiting distribution of the scaled spectral radius and minimum modulus for the complex induced Ginibre ensemble [6] are presented in Section 3.2.
 This random matrix ensemble might find relevance for questions raised in different fields of physics.
@@
 The complex induced Ginibre ensemble is a special case of the Feinberg-Zee ensemble and a generalisation of the complex Ginibre ensemble.
-The joint probability density function of the entries of matrices $A$ from the Feinberg-Zee ensemble is
+The joint probability density function of the entries of matrices $A$ from the Feinberg-Zee ensemble is of the form
 $P_{FZ}(A)\propto\exp\left(-\operatorname{Tr}V(A^{\dagger}A)\right)$ (5)
-The induced Ginibre ensembles correspond to the Feinberg-Zee ensemble with potential $V(y)=-\frac{\beta}{2}(y-L\log(y))$ where $L$ is a non-negative parameter called the rectangularity index.
+The induced Ginibre ensembles correspond to the Feinberg-Zee ensemble with potential $V$ expressed as $V(y)=-\frac{\beta}{2}(y-L\log(y))$, where $L$ is a non-negative parameter called the rectangularity index, and $\beta=1$ or $\beta=2$ for the real or complex induced Ginibre ensembles, respectively.
 Let $G$ denote a $N\times N$ complex induced Ginibre matrix. The joint probability density function of its entries is
-$P_{indG}^{(\beta)}(G)=\Gamma^{(\beta)}\left(\det G^{\dagger}G\right)^{\frac{\beta}{2}L}\exp\left(-\operatorname{Tr}G^{\dagger}G\right)\\$ (6)
-where $\beta=2$ for the complex induced Ginibre ensemble and the determinant $\left(\det G^{\dagger}G\right)^{a}=\prod_{k=1}^{N}|\lambda_{i}|^{2a}$, $a\geq 0$.
-The factor $\Gamma^{(\beta)}$ is the normalisation constant
+$P_{indG}^{(\beta)}(G)=\Gamma^{(\beta)}\left(\det G^{\dagger}G\right)^{\frac{\beta}{2}L}\exp\left(-\operatorname{Tr}G^{\dagger}G\right)$ (6)
+where $\beta=2$ and the determinant $\left(\det G^{\dagger}G\right)^{a}=\prod_{k=1}^{N}|\lambda_{i}|^{2a}$, $a\geq 0$. The factor $\Gamma^{(\beta)}$ is the normalisation constant
 $\Gamma^{(\beta)}=\pi^{-\frac{\beta}{2}N^{2}}\left(\frac{\beta}{2}\right)^{\frac{1}{2}N(N-L)}\prod_{j=1}^{N}\frac{\Gamma(\frac{\beta}{2}j)}{\Gamma(\frac{\beta}{2}(j+L)}$ (7)
-In the limit of large matrices dimensions, the eigenvalues are spread across an annulus in the complex plane which is distinguished from the phenomenon identified for complex Ginibre ensemble thus referenced as the circular law and presented in the work from V.
+In the limit of large matrix dimensions, the eigenvalues are spread across an annulus in the complex plane which is distinguished from the phenomenon identified for complex Ginibre ensemble, thus referenced as the circular law and presented in the work from V.
 L.
@@
 Krishnapur [25].
-The joint probability density function of the eigenvalues of complex induced Ginibre matrices is
-$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{\Gamma_{N}}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)\\$ (8)
+The joint probability density function of the eigenvalues of the complex induced Ginibre matrix $G$ is
+$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{\Gamma_{N}}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)$ (8)
 with normalisation constant
@@
 This phenomenon is explained from a mathematical perspective with the formulation of the mean eigenvalue density, i.e., $\langle\rho_{N}(\lambda)\rangle=R_{1}(\lambda)$ where $R_{1}(\lambda)$ is the one-point correlation function of the eigenvalues.
-More precisely taking its limit as $N$ goes to infinity and for a rectangularity index $L$, the distribution of the eigenvalues is uniform and supported by a ring centred at the origin of the complex plane with a outer circle of radius $r_{out}=\sqrt{L+N}$ and an inner circle of radius $r_{in}=\sqrt{L}$.
+More precisely taking its limit as $N$ goes to infinity and for a non-negative rectangularity index $L$, the distribution of the eigenvalues is uniform and supported by a ring centred at the origin of the complex plane with an outer circle of radius $r_{out}=\sqrt{L+N}$ and an inner circle of radius $r_{in}=\sqrt{L}$.
+This phenomenon refers to the ”Single Ring Theorem” stating that, in the limit of large matrix dimensions, the empirical distribution of the eigenvalues has for support a single ring [5] [15].
 For small rectangularity indexes $L$ (close to zero), the density of the eigenvalues is uniform over the disk of radius $\sqrt{N}$ with a mean eigenvalue density converging to the mean eigenvalue density of the complex Ginibre ensemble.
-In the present paper, an investigation is performed for the complex induced Ginibre ensemble from the work of B.
+A study is performed Section 3.2 for the complex induced Ginibre ensemble from the work of B.
 Rider [22].
 In this reference, the author details an analytical method to determine the limiting distribution of the scaled spectral radius for matrices from the complex and symplectic Ginibre ensembles.
-A similar investigation is undertaken here for the scaled spectral radius and scaled minimum modulus of eigenvalues of matrices from the complex induced Ginibre ensemble.
-The analysis is conducted at the outer and inner edges of the ring (the eigenvalues support on the complex plane) defined with a rectangularity index assumed to be proportional to $N$, i.e., $L=\alpha N$, $\forall\alpha>0$.
+A similar analysis is undertaken in Section 3.2.1 for the scaled spectral radius and scaled minimum modulus of eigenvalues of matrices from the complex induced Ginibre ensemble.
+The study is conducted at the outer and inner edges of the ring (the eigenvalues support on the complex plane) defined with a rectangularity index assumed to be proportional to $N$, i.e., $L=\alpha N$, $\forall\alpha>0$.
 An exact fit between the empirical distribution and analytical formulation of these extreme moduli is acknowledged numerically.
-Additionally, the independence of the spectral radius and minimum modulus is studied in the limit as $N$ goes to infinity for these two non-Hermitian ensembles.
-Results stating the right and left tail eigenvalues distribution functions asymptotics for the complex elliptic Ginibre ensemble in the limit of weak non-Hermiticity is presented in [4].
-The limiting tail distribution functions asymptotics of the eigenvalues modulus is also of interest.
-The similitude of the minimum modulus limiting (left and right) tail distributions between the two investigated non-Hermitian ensembles is explored in the present paper.
-This is the purpose of Section 4.
-Conclusions are set out in Section 5.
-2. Limit theorems for the complex Ginibre ensemble
-Let $A$ denote a $N\times N$ complex Ginibre matrix whose symmetrised joint probability density function of its eigenvalues, $P_{N}(z_{1},\cdots,z_{N})$, is defined as in Section 1.
-The probability that the minimum of the eigenvalue moduli, here denoted $r_{min}^{(N)}(A)$, is greater than the radius $r\in\mathbb{R}^{+}$ is expressed as follows
-$P\left(r_{min}^{(N)}(A)\geq r\right)=\int_{|z_{1}|\geq r}\cdots\int_{|z_{N}|\geq r}P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d^{2}z_{k}$
-This probability corresponds to the probability that no eigenvalue lies inside the disk of radius $r$.
-It is a function of the radius $r$ and is initially defined analytically in reference from P.
-J.
-Forrester [6] with the following formula,
-$P\left(r_{min}^{(N)}(A)\geq r\right)=\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ (10)
-This result is presented in the reference from [2] whose authors applied the Gram’s formula [18] to get the probability, that no eigenvalue lies in the disk of radius $r$ centred at the origin, as a finite product of regularised upper incomplete Gamma functions.
-Equation (10) is derived with another approach in the present paper as follows.
-More precisely, the Andreief’s integration formula [7] is used to express analytically the gap probabilities $P(r_{min}^{(N)}(A)\geq a)$, $\forall a\in\mathbb{R}^{+*}$.
-More precisely, from the definition of the eigenvalues joint probability density function defined in Section 1,
-$P\left(r_{min}^{(N)}(A)\geq a\right)=\frac{1}{\Gamma_{N}}\int_{|z_{1}|>a}\cdots\int_{|z_{N}|>a}\det\left[z_{k}^{N-j}\right]_{j,k=1}^{N}\det\left[\bar{z}_{k}^{N-j}\right]_{j,k=1}^{N}\prod_{k=1}^{N}dm(z_{k})$
+Additionally, the independence of the spectral radius and minimum modulus is studied in the limit as $N$ goes to infinity for these two non-Hermitian ensembles in Section 3.2.2.
+The limiting tail distribution functions of the eigenvalues modulus is also of interest in the field.
+Results stating the right and left tail eigenvalues distribution function asymptotics for the complex elliptic Ginibre ensemble in the limit of weak non-Hermiticity is presented in [4].
+The similitude of the minimum modulus limiting right tail distributions between the two non-Hermitian ensembles at scrutiny is explored in Section 3.3 of the present paper.
+The same analysis is performed for the left tail distribution of this extreme modulus.
+A discussion and a list of conclusions are set out in Section 4.
+2. Methods
+Let $P\left(r_{min}^{(N)}(A)\geq r\right)$ denote the survival distribution function of the minimum modulus $r_{min}^{(N)}(A)$ and
+$P\left(r_{max}^{(N)}(A)\leq r\right)$ the cumulative distribution function of the spectral radius $r_{max}^{(N)}(A)$, for a $N\times N$ complex matrix $A$.
+These distribution functions are derived analytically, in the present paper, with the use of Andreief’s integration formula presented in [8].
+A straightforward application is as follows.
+Let $A$ denote a $N\times N$ complex Ginibre matrix whose the symmetrised joint probability density function of its eigenvalues, denoted $P_{N}(z_{1},\cdots,z_{N})$, is defined as in Section 1 with Equation (3).
+The probability that the minimum of the eigenvalue moduli, here denoted $r_{min}^{(N)}(A)$, is greater than the radius $a\in\mathbb{R}^{+}$, is
+$P\left(r_{min}^{(N)}(A)\geq a\right)$ $=\int_{|z_{1}|\geq a}\cdots\int_{|z_{N}|\geq a}P_{N}(z_{1},...,z_{N})\prod_{k=1}^{N}d^{2}z_{k}$
+$=\frac{1}{\Gamma_{N}}\int_{|z_{1}|>a}\cdots\int_{|z_{N}|>a}\det\left[z_{k}^{N-j}\right]_{j,k=1}^{N}\det\left[\bar{z}_{k}^{N-j}\right]_{j,k=1}^{N}\prod_{k=1}^{N}dm(z_{k})$
 where $dm(z_{k})=e^{-|z_{k}|^{2}}d^{2}z_{k}$ and $\Gamma_{N}=N!\pi^{N}\prod_{k=0}^{N-1}k!$ is the normalisation constant.
-Applying the Andreief’s integration formula,
+Then, performing a change of variables in polar coordinates gives,
 $P\left(r_{min}^{(N)}(A)\geq a\right)=\frac{1}{\Gamma_{N}}N!\prod_{k=0}^{N-1}2\pi\int_{a}^{+\infty}e^{-r^{2}}r^{2k}rdr=\prod_{k=0}^{N-1}\frac{1}{\Gamma(k+1)}\int_{a^{2}}^{+\infty}e^{-t}t^{k}dt$
-It is elementary to derive the probability density function $p_{r_{min}^{(N)}(A)}$ of the smallest modulus $r_{min}^{(N)}(A)$ from Equation (10).
-It is convenient to take the logarithm of the probability $P(r_{min}^{(N)}(A)\geq r)$ and find its first derivative with respect to $r$ to finally get the probability density function of the smallest modulus $r_{min}^{(N)}(A)$ as
+This probability corresponds to the probability that no eigenvalue lies inside the disk of radius $a$ and is analytically formulated as
+$P\left(r_{min}^{(N)}(A)\geq a\right)=\prod_{k=0}^{N-1}\frac{\Gamma(k+1,a^{2})}{\Gamma(k+1)}$ (10)
+This result is initially stated without proof and without scaling applied to the minimum modulus in the reference [14], and for which asymptotic expansions are studied in [7].
+It is also presented in the reference [2] in which the Gram’s formula [20] is applied to get this probability (the probability that no eigenvalue lies in the disk of radius $a$ centred at the origin of the complex plane) as a finite product of regularised upper incomplete Gamma functions.
+It is straightforward to derive the probability density function $p_{r_{min}^{(N)}(A)}$ of the minimum modulus $r_{min}^{(N)}(A)$ from Equation (10).
+More precisely, it is convenient to take the logarithm of the probability $P(r_{min}^{(N)}(A)\geq r)$ and find its first derivative with respect to $r$.
+Let the survival distribution function $P(r_{min}^{(N)}(A)\geq r)$ denote the function $H(r)$ of the radius $r$.
+The corresponding probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is $p_{r_{min}^{(N)}(A)}(r)=-\frac{\text{d}H(r)}{\text{d}r}$.
+Furthermore,
+$\frac{\text{d}H(r)}{\text{d}r}$ $=H(r)\frac{\text{d}}{\text{d}r}\left(\log\left[H(r)\right]\right)=\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\left[\sum_{k=0}^{N-1}\frac{\partial}{\partial r}\left[\log\left(\Gamma(k+1,r^{2})\right)\right]\right]$
+with,
+$\frac{\partial\log\left[\Gamma(k+1,r^{2})\right]}{\partial r}$ $=\frac{1}{\Gamma(k+1,r^{2})}\frac{\partial\Gamma(k+1,r^{2})}{\partial r}$
+It is known that $\frac{\partial\Gamma(s,x)}{\partial x}=-x^{s-1}e^{-x}$.
+Thus,
+$\frac{\text{d}H(r)}{\text{d}r}$ $=\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}\frac{\partial\Gamma(j+1,r^{2})}{\partial r}\right]$
+where
+$\frac{\partial\Gamma(j+1,r^{2})}{\partial r}=-2r^{2j+1}e^{-r^{2}}$
+Finally, the probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is,
 $p_{r_{min}^{(N)}(A)}\left(r\right)$ $=2re^{-r^{2}}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]$ (11)
-Using a similar approach as for the case of no scaling, the survival distribution function of scaled minimum of eigenvalue moduli $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ for $N\times N$ complex Ginibre matrices $A$ has same analytical expression as for $r_{min}^{(N)}(A)$ replacing $r$ with $\sqrt{N}r$.
+Using a similar approach as for the case of no scaling, the survival distribution function of scaled minimum of eigenvalue moduli $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ for an $N\times N$ complex Ginibre matrix $A$ has same analytical expression as for $r_{min}^{(N)}(A)$ replacing $r$ with $\sqrt{N}r$.
 The corresponding probability density function is the probability density function of the minimum modulus $r_{min}^{(N)}(A)$ multiply by $\sqrt{N}$ at the point $\sqrt{N}r$, i.e.,
 $p_{\frac{r_{min}^{(N)}(A)}{\sqrt{N}}}\left(r\right)=\sqrt{N}p_{r_{min}^{(N)}(A)}\left(\sqrt{N}r\right)$ (12)
-The spectral radius of the real Ginibre matrices is investigated in reference [9].
+The spectral radius of matrices from the real Ginibre ensemble is investigated in reference [10].
 It is experimentally stated that its value converges almost surely to the standard deviation of the i.i.d.
 entries of scaled matrices from this random matrix ensemble.
-The distribution of the spectral radius for a fixed size of complex Ginibre matrices is explored in the following.
+The distribution function of the spectral radius for a fixed size of complex Ginibre matrices is explored in the following.
 Lemma 1.
+Let $A$ denote an $N\times N$ complex Ginibre matrix. Let $r_{max}(A)^{(N)}$ denote the spectral radius.
+Then,
 $P\left(r_{max}^{(N)}(A)\leq r\right)=\prod_{k=0}^{N-1}\frac{\gamma(k+1,r^{2})}{\Gamma(k+1)}$ (13)
@@
 Proof.
-The cumulative distribution function of the spectral radius of eigenvalues, denoted $r_{max}^{(N)}(A)$, for the complex Ginibre ensemble is also easily retrieved with the use of the Andreief’s integration formula applied to the following multiple integrals equation
+The cumulative distribution function of the spectral radius $r_{max}^{(N)}(A)$ for the complex Ginibre ensemble is also easily retrieved with the use of Andreief’s integration formula applied to the following multiple integrals equation
 $P\left(r_{max}^{(N)}(A)\leq r\right)$ $=\int_{|z_{1}|\leq r}\cdots\int_{|z_{N}|\leq r}P_{N}(z_{1},\cdots,z_{N})\prod_{k=1}^{N}d^{2}z_{k}=\prod_{k=0}^{N-1}\frac{1}{\Gamma(k+1)}\int_{0}^{r^{2}}e^{-t}t^{k}dt$
 Corollary 2.
-The $N$-th gap probability, i.e., the probability that all the eigenvalues of a $N\times N$ complex Ginibre matrix lies in the disk of radius $r$, is the joint distribution of independent random variables $\gamma_{k}$ each following a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ [1], with $\alpha_{k}=k$ and $\delta_{k}=1$.
+The $N$-th gap probability, i.e., the probability that all the eigenvalues of an $N\times N$ complex Ginibre matrix lies in the disk of radius $r$, is the joint distribution of independent random variables $\gamma_{k}$ each following a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ [1], with $\alpha_{k}=k$ and $\delta_{k}=1$.
 Proof.
 $P(r_{1}^{(N)}(A)\leq r,r_{2}^{(N)}(A)\leq r,...,r_{N}^{(N)}(A)\leq r)=\prod_{k=1}^{N}P(\gamma_{k}<r)$ (14)
-where each random variables $\gamma_{k}$ are independent and follows a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$ , with $\alpha_{k}=k$ and $\delta_{k}=1$.
-The Gamma-Rayleigh distribution is derived in the work of E.
+where the random variables $\gamma_{k}$ are independent and each follows a Gamma-Rayleigh distribution $GR(\alpha_{k},\delta_{k})$, with parameters $\alpha_{k}=k$ and $\delta_{k}=1$.
+The Gamma-Rayleigh distribution is derived in the work from E.
 Akarawak, I.
@@
 Okafor [1].
+∎
 The corresponding probability density function of the largest modulus $r_{max}^{(N)}(A)$ is derived from Equation (13).
-Taking the logarithm of the cumulative distribution function $P(r_{max}^{(N)}(A)\leq r)$ and then the first derivative with respect to $r$ of $\log\left(P(r_{max}^{(N)}(A)\leq r)\right)$, the probability density funvtion of the spectral radius $r_{max}^{(N)}(A)$ is,
+Taking the logarithm of the cumulative distribution function $P(r_{max}^{(N)}(A)\leq r)$ and then the first derivative with respect to $r$ of $\log\left(P(r_{max}^{(N)}(A)\leq r)\right)$, the probability density function of the spectral radius $r_{max}^{(N)}(A)$ is,
 $p_{r_{max}^{(N)}(A)}(r)=2re^{-r^{2}}\prod_{k=0}^{N-1}\frac{\gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2j}}{\gamma(j+1,r^{2})}\right]$ (15)
-The scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ is now considered in the following.
-The probability distribution function $P\left(\frac{r_{max}^{(N)}(A)}{\sqrt{N}}\leq r\right)$ of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ is also derived using the Andreief’s integration formula [7],
+A similar method of derivation (as the one applied for the scaled minimum modulus) leads to the formulation of the probability density function of the scaled spectral radius for the complex Ginibre ensemble as
+$p_{\frac{r_{max}^{(N)}(A)}{\sqrt{N}}}\left(r\right)=\sqrt{N}p_{r_{max}^{(N)}(A)}\left(\sqrt{N}r\right)$ (16)
+Asymptotic results are then derived from these analytical expressions in the limit as the size of matrices goes to infinity. They are presented in Section  3.
+The limiting distribution of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ for $N\times N$ complex Ginibre matrices is established in the research paper from B.
+Rider [22].
+His framework aims to derive the logarithm of the cumulative distribution function of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ for the complex Ginibre ensemble, in the limit as $N$ goes to infinity.
+The limiting distribution of the extreme moduli $r_{min}^{(N)}(G)$ and $r_{max}^{(N)}(G)$ and their independence for $N\times N$ matrices $G$ from the complex induced Ginibre ensemble, are derived at appropriate scaling with this framework whose main steps are detailed as follows.
+First, analytical formula of the survival (or cumulative) distribution function of these random variables are derived as a partial product of the cumulative distribution function of an averaged sum of independent and identically distributed random variables, each following a standard exponential distribution.
+As an illustration, B.
+Rider [22] considered the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ and established that
 $P\left(\frac{r_{max}^{(N)}(A)}{\sqrt{N}}\leq r\right)=\prod_{k=0}^{N-1}P\left(\frac{1}{N}\sum_{j=1}^{N-k}Z_{j}\leq r^{2}\right)$ (17)
@@
 Each of the $Z_{j}$ follows a standard exponential distribution.
-This is a known result presented in B.
-Rider’s work [20] from which he established the nature of the limiting distribution of the scaled spectral radius for matrices from the complex Ginibre ensemble in the limit as $N$ goes to infinity.
-The distribution of the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ for matrices from the complex Ginibre ensemble is the standard Gumbel distribution in the limit as $N$ goes to infinity.
-The limiting distribution of scaled $k$-th modulus has also been established for this random matrix ensemble as well as for the symplectic Ginibre ensemble in [21], in light of the framework presented in [20].
-A similar method of derivation (as the one applied for the scaled minimum modulus) leads to the formulation of the probability density function of the scaled spectral radius for the complex Ginibre ensemble as
-$p_{\frac{r_{max}^{(N)}(A)}{\sqrt{N}}}\left(r\right)=\sqrt{N}p_{r_{max}^{(N)}(A)}\left(\sqrt{N}r\right)$ (17)
-2.1. Limiting left and right tail distributions of the minimum modulus as $N$ goes to infinity
+Then, lower and upper bounds are derived for this analytical formula in the limit as $N$ goes to infinity.
+The limit, as $N$ goes to infinity, of the logarithm of this distribution function is then established by applying the classical Edgeworth expansion.
+Then, an approximation from this expression considering the leading order sum as the Riemann integral is derived.
+Finally, an appropriate definition of a random function is chosen such that the limiting distribution of the considered random variable, here the scaled spectral radius $\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$, is the identified statistical extreme distribution, here the Gumbel distribution.
+3. Results
+3.1. Limit theorems for the complex Ginibre ensemble
+Limiting distributions of extreme moduli of matrices from non-Hermitian ensembles are of interest in the field.
+The limiting distribution of the spectral radius is studied in references [22, 23, 24].
+Extreme moduli limiting left and right tail distribution as well as their mutual independence have not been investigated in the literature and are the themes of next Sections 3.1.1, 3.1.2, 3.2.1 and 3.2.2 for the complex Ginibre ensemble and its matrix ensemble generalisation, the complex induced Ginibre ensemble.
+Limiting stochastic expression of these extreme random variables are established for the complex induced Ginibre ensemble.
+3.1.1. Limiting left and right tail distributions of the minimum modulus for the complex Ginibre ensemble as $N$ goes to infinity
 Theorem 3.
-Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble.
-The left tail distribution of minimum modulus $r_{min}^{(N)}(A)$ (i.e., the smallest values of $r_{min}^{(N)}(A)$) converges to the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$ as $N$ goes to infinity
+Let $A$ denote an $N\times N$ matrix from the complex Ginibre ensemble.
+The left tail distribution of minimum modulus $r_{min}^{(N)}(A)$ converges to the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$, as $N$ goes to infinity.
 More precisely, for $0<r\ll 1$,
@@
 Proof.
-The statement of Theorem 3 is derived from the formulation of the gap probability $P(r_{min}^{(N)}(A)\geq r)$ which corresponds to the $N$-th partial product presented in Equation (10).
-$\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ $=e^{-r^{2}}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=e^{-r^{2}}\prod_{k=1}^{N-1}e^{-r^{2}}e_{k}(r^{2})$
-$=e^{-r^{2}}\prod_{k=1}^{N-1}\left[1-e^{-r^{2}}\sum_{t=n}^{+\infty}\frac{r^{2t}}{t!}\right]=e^{-r^{2}}H^{(N)}(r,0)$
-where $e_{n}(x)=\sum_{k=0}^{n}\frac{x^{k}}{k!}$ defines the $n$-th partial sum of the exponential function and $H^{(N)}(r,0)$ is the conditional probability (cf.
-Appendix B) that given one eigenvalue lies at the origin of the complex plane all the others are found outside the disk centred at zero with radius $r$.
-The limit
+The statement of Theorem 3 is derived from the formulation of the gap probability $P(r_{min}^{(N)}(A)\geq r)$ which corresponds to the $N$-th partial product presented in Equation (10),
+$\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ $=e^{-r^{2}}\prod_{k=1}^{N-1}e^{-r^{2}}e_{k}(r^{2})=e^{-r^{2}}H^{(N)}(r,0)$
+where $e_{n}(x)=\sum_{k=0}^{n}\frac{x^{k}}{k!}$ defines the $n$-th partial sum of the exponential function and $H^{(N)}(r,0)$ is the conditional probability that given one eigenvalue lies at the origin of the complex plane all the others are found outside the disk centred at zero with radius rr.
+[^1] The derivation of the conditional probability $H^{(N)}(r,0)$ is presented in the appendix.
+Thus,
 $\lim_{N\rightarrow+\infty}P(r_{min}^{(N)}(A)\geq r)=e^{-r^{2}}\lim_{N\rightarrow+\infty}H^{(N)}(r,0)$ (19)
-This conditional probability $H^{(N)}(r,0)$ is a result presented in [13] and [15], where, in the limit as $N$ goes to infinity and $r$ is small (or $r$ close to zero)
+The conditional probability $H^{(N)}(r,0)$ is a result presented in [14] and [17], where, in the limit as $N$ goes to infinity and $r$ is small (or $r$ close to zero)
 $\lim_{N\rightarrow+\infty}H^{(N)}(r,0)=1-\left[\frac{r^{4}}{2}+\frac{r^{6}}{6}+\frac{r^{8}}{24}+O(r^{10})\right]=1-O\left(r^{4}\right)$ (20)
-The limit presented in Equation (20) is derived as follows. For small $x$, the lower incomplete gamma function has asymptotic expansion
+The limit presented in Equation (20) is derived as follows. The lower incomplete gamma function has series expansion, for $x$ around zero,
 $\gamma(a,x)=x^{a}\sum_{n=0}^{+\infty}(-1)^{n}\frac{x^{n}}{n!(a+n)}$ (21)
@@
 Then, for small $r$, i.e., $0<r\ll 1$,
-$\lim_{N\rightarrow+\infty}H^{(N)}(r,0)=\prod_{k=1}^{+\infty}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{\gamma(k+1,r^{2})}{\Gamma(k+1)}\right]$
-$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{r^{2(k+1)}}{(k+1)\Gamma(k+1)}-\frac{r^{2(k+1)}}{\Gamma(k+1)}\sum_{n=1}^{+\infty}(-1)^{n}\frac{r^{2n}}{n!(k+n+1)}\right]$
-$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{r^{2(k+1)}}{(k+1)\Gamma(k+1)}+O\left(r^{2(k+2)}\right)\right]$
+$\lim_{N\rightarrow+\infty}H^{(N)}(r,0)=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\left[1-\frac{r^{2(k+1)}}{(k+1)\Gamma(k+1)}+O\left(r^{2(k+2)}\right)\right]$
 Thus,
-$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$
-$=\exp\left[\log\left(\prod_{k=1}^{+\infty}\left[1-\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right]\right)\right]=\exp\left[\sum_{k=1}^{+\infty}\log\left(1-\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)\right]$
-$=\exp\left[\sum_{k=1}^{+\infty}\left[-\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]\right]=\exp\left[-\sum_{k=1}^{+\infty}\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]$
-$=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]^{j}$
+$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$ $=\exp\left[\sum_{k=1}^{+\infty}\left[-\sum_{\gamma=1}^{+\infty}\frac{1}{\gamma}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{2(k+2)}\right)\right)^{\gamma}\right]\right]$
 $=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{4(k+1)}\right)\right)\right]^{j}$
 which implies for small $r$ (or $r$ close to zero),
-$\lim_{N\rightarrow+\infty}H^{(N)}(0,r)$ $=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$
-$=\left[\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{4(k+1)}\right)\right)\right]^{j}\right]$
-$=1-\left[\frac{r^{4}}{2}+\frac{r^{6}}{6}+\frac{r^{8}}{24}+O(r^{10})\right]=1-O\left(r^{4}\right)$
-This implies that
-$\lim_{N\rightarrow+\infty}H^{(N)}(r,0)$ $=\sum_{j=0}^{+\infty}\frac{(-1)^{j}}{j!}\left[\sum_{k=1}^{+\infty}\left(\frac{r^{2(k+1)}}{(k+1)!}+O\left(r^{4(k+1)}\right)\right)\right]^{j}$
+$\lim_{N\rightarrow+\infty}H^{(N)}(0,r)=1-O\left(r^{4}\right)$
 Consequently, for $0<r\ll 1$,
@@
 This does correspond to the cumulative distribution of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
-Remark 2.1.
-Using successive integrations by parts
-$\int_{r^{2}}^{+\infty}e^{-t}t^{k}dt$ $=e^{-r^{2}}\sum_{j=0}^{k}\frac{k!}{(k-j)!}r^{2(k-j)}+k!\int_{r^{2}}^{+\infty}e^{-t}dt$ (22)
-Consequently,
-$P(r_{min}^{(N)}(A)\geq r)$ $=\prod_{k=0}^{N-1}\left[\frac{1}{\Gamma(k+1)}e^{-r^{2}}\sum_{j=0}^{k}\frac{k!}{(k-j)!}r^{2(k-j)}+\int_{r^{2}}^{+\infty}e^{-t}dt\right]$ (23)
 Theorem 4.
-Let $A$ denote a $N\times N$ matrix from the complex Ginibre ensemble.
+Let $A$ denote an $N\times N$ matrix from the complex Ginibre ensemble.
 The right tail distribution of the minimum of moduli $r_{min}^{(N)}(A)$ converges to the Weibull distribution with shape parameter $\kappa=4$ and scale parameter $\lambda=k^{\frac{1}{k}}$ as $N$ goes to infinity, i.e., for large $r$,
@@
 This result is established from Equation $9$ of the reference [14].
-Let $r=\sqrt{N}|\lambda|$ such that $|\lambda|\leq 1$, where $\lambda$ is a complex number lying in the unit disk centred at the origin of the complex plane.
+Let $r=\sqrt{N}|\lambda|$ such that $|\lambda|\leq 1$, where $\lambda$ is a complex number lying in the unit disk centred at the origin of the complex plane. The limit
 $\lim_{N\rightarrow+\infty}P\left(\frac{r_{min}^{(N)}(A)}{\sqrt{N}}\geq|\lambda|\right)$ $=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)\geq\sqrt{N}|\lambda|\right)$
@@
 $=\exp\left(-\frac{N^{2}|\lambda|^{4}}{4}\left(1+O\left(\frac{1}{N|\lambda|^{2}}\right)\right)\right)$
-This result corresponds to the distribution of rare events, i.e., the distribution of extreme events corresponding to the scaled minimum of the moduli $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ as $N$ goes to infinity.
-This result refers to the large deviations theory.
-This limit at the logarithmic scale, as $N$ goes to infinity, is
+This result corresponds to the distribution of extreme events of the scaled minimum modulus $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ as $N$ goes to infinity.
+This result refers to the large deviations theory, where
 $\lim_{N\longrightarrow+\infty}\left(-\log\left[P\left(\frac{r_{min}^{(N)}(A)}{\sqrt{N}}\geq|\lambda|\right)\right]\right)=\frac{N^{2}}{4}|\lambda|^{4}=c_{N}\inf_{\{z\geq|\lambda|\}}{I(z)}$ (23)
 with speed $c_{N}=\frac{N^{2}}{4}$ and rate function $I(z)=|z|^{4}$.
-2.2. Independence of the scaled spectral radius and the scaled minimum modulus for the complex Ginibre ensemble
+3.1.2. Independence of the scaled spectral radius and the scaled minimum modulus for the complex Ginibre ensemble
 Theorem 5.
-Let A denote a $N\times N$ complex Ginibre matrix.
-The scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ are independent random variables.
+Let A denote an $N\times N$ complex Ginibre matrix.
+The scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(A)}{\sqrt{N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ are independent random variables, in the limit as $N$ goes to infinity.
 Proof.
-Applying the Andreief’s integration formula
+Applying Andreief’s integration formula [8],
 $P\left(r_{N}\geq r\text{ and }R_{N}\leq R\right)$ $=\prod_{k=0}^{N-1}\frac{\gamma(k+1,NR^{2})}{\Gamma(k+1)}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]$
@@
 The specification of $R^{(N,x)}$ is slightly different from the result presented in [22] as a scaling $\sqrt{N}$ is used here $\left(\text{and not }2\sqrt{N}\right)$ which corresponds to the radius of the disk defined as the support of the spectrum of $N\times N$ random matrices from the complex Ginibre ensemble.
-From the results presented in [2], for small $r$ of order $\frac{1}{\sqrt{N}}$ (i.e.
-where $Nr^{2}$ is kept fixed while $N$ goes to infinity), as $N$ goes infinity (or $N$ very large) also known as the origin limit
-$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}E_{0}^{(2)}\left(\sqrt{N}r\right)$
-Defined in [2], $E_{0}^{(\beta)}\left(\sqrt{N}r\right)$ is the probability that zero eigenvalue lies inside the disk of radius $\sqrt{N}r$ centred at the origin of the complex plane, and all eigenvalues lie outside.
+From the results presented in [2], for small $r$ of order $\frac{1}{\sqrt{N}}$ (i.e., where $Nr^{2}$ is kept fixed while $N$ goes to infinity), as $N$ goes infinity (or $N$ very large) also known as the origin limit
+$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}E_{0}^{(2,N)}\left(r\right)$
+Defined in [2], $E_{0}^{(\beta,N)}\left(r\right)$ is the probability that zero ($\sqrt{N}\text{ scaled}$-) eigenvalue lies inside the disk of radius $r$ centred at the origin of the complex plane, and all eigenvalues lie outside.
 The index $\beta=2$ for the complex Ginibre ensemble.
 This implies that
-$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}P\left(\frac{r_{min}^{(N)}(A)}{\sqrt{N}}\geq r\right)=\lim_{N\rightarrow+\infty}P\left(r_{N}\geq r\right)$
+$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\left[1-\frac{\gamma(k+1,Nr^{2})}{\gamma(k+1,NR^{2})}\right]=\lim_{N\rightarrow+\infty}P\left(r_{N}\geq r\right)$
 Consequently,
@@
 The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ for the complex Ginibre ensemble are then independent random variables in the limit as $N$ goes to infinity.
-3. Limit theorems for the complex induced Ginibre ensemble
-The limiting distribution of the spectral radius of non-Hermitian ensembles as well as its precise localisation near the edge of the unit disk is investigated in the work of B.
+3.2. Limit theorems for the complex induced Ginibre ensemble
+The limiting distribution of the spectral radius of non-Hermitian ensembles as well as its precise localisation near the edge of the unit disk is investigated in the work from B.
 Rider [22].
@@
 These statistical ensembles share similar features such as the universality conjecture known as the circular law.
-Their eigenvalues move towards the unit disk as the size of the matrices increases.
+Their eigenvalues (scaled with respect to $\sqrt{N}$) move towards the unit disk as the size of the matrices increases.
 The following results are derived for the complex induced Ginibre ensemble using the methodological approach exposed in Rider’s work.
-Let $G$ denote a $N\times N$ random matrix from the complex induced Ginibre ensemble with rectangularity index $L$ defined in [5].
-At appropriate scalings, the scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{L+N}}$ evolves near the outer radius $r_{out}$ while the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{L}}$ is a random variable fluctuating around the inner radius $r_{in}$.
-The outer and inner radii are equal to $\sqrt{L+N}$ and $\sqrt{L}$, respectively.
-3.1. Limit theorems of the outer and inner edges of the ring
+Let $G$ denote an $N\times N$ random matrix from the complex induced Ginibre ensemble with rectangularity index $L$ as defined in [6].
+The scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{L+N}}$ evolves near the outer radius $r_{out}$ while the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{L}}$ is a random variable fluctuating around the inner radius $r_{in}$.
+3.2.1. Limit theorems at the outer and inner edges of the ring
 The scaled spectral radius and scaled minimum modulus of eigenvalues for $N\times N$ matrices from the complex induced Ginibre ensemble, with proportional rectangularity index, are Gumbel distributed in the limit as $N$ goes to infinity.
@@
 Theorem 6.
-Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}$ denote the scaled spectral radius.
+Let $G$ denote an $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}$ denote the scaled spectral radius.
 Then,
@@
 where $\gamma_{\alpha,N}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$.
-This limiting probability distribution is the standard Gumbel distribution for maxima. The corresponding cumulative density function is the function $F_{X}^{(\text{Gumbel max})}(x)=\exp(-\exp(-x))$.
-The scaled spectral radius $R_{N}$ is approximated as
-$R_{N}\approx 1+T_{\alpha,N}+\xi_{\alpha,N}$ (28)
+The scaled spectral radius $R_{N}$ is approximated, as $N$ goes to infinity, by
+$R_{N}\simeq 1+T_{\alpha,N}+\xi_{\alpha,N}$ (26)
 where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$.
-The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ with $X$ following a standard Gumbel (maximum) distribution.
+The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ with $X$ following a standard Gumbel distribution for maxima.
 This result is similar to the one detailed in [22] for the complex Ginibre ensemble.
-The exact formula of the probability density function $p_{R_{N}}\left(r\right)$ of the scaled spectral radius $R_{N}$, for a proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$, is derived from equation (26) and is defined as follows for any value of NN
-$p_{R_{N}}\left(r\right)=2[(1+\alpha)N]^{\alpha N}r^{2(\alpha N-\frac{1}{2})}e^{-(1+\alpha)Nr^{2}}\prod_{k=1}^{N}\frac{\gamma\left(k+\alpha N,(1+\alpha)Nr^{2}\right)}{\Gamma(k+\alpha N)}\sum_{j=1}^{N}\left[\frac{[(1+\alpha)N]^{j}r^{2j}}{\gamma\left(j+\alpha N,(1+\alpha)Nr^{2}\right)}\right]$ (29)
-with $\gamma(.,.)$ is the lower incomplete Gamma function.
-The proof of Theorem 6 is presented in Appendix A.1.
-The analytical formulation of the probability density function $p_{R_{N}}$ of the scaled spectral radius $R_{N}$ does fit exactly its empirical distribution created from the generation of 1010 000000 complex induced Ginibre matrices with proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$.
-This is presented in Figure 1.
-As the size of these matrices increases to large numbers, the distribution of $R_{N}$ narrowing close to the outer radius $r_{out}=1$ (Figure 2).
-(a) $\alpha=\frac{1}{9}$
-(b) $\alpha=\frac{1}{90}(red),\alpha=\frac{1}{9}(blue)\text{ and }\alpha=\frac{4}{9}(black)$
+The exact formula of the probability density function $p_{R_{N}}\left(r\right)$ of the scaled spectral radius $R_{N}$, for a proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$, is derived from equation (24) and is defined as follows for any value of $N$,
+$p_{R_{N}}\left(r\right)=C^{(\alpha,N)}(r)\prod_{k=1}^{N}\frac{\gamma\left(k+\alpha N,(1+\alpha)Nr^{2}\right)}{\Gamma(k+\alpha N)}\sum_{j=1}^{N}\left[\frac{[(1+\alpha)N]^{j}r^{2j}}{\gamma\left(j+\alpha N,(1+\alpha)Nr^{2}\right)}\right]$ (27)
+where $\gamma(.,.)$ is the lower incomplete Gamma function and the function
+$C^{(\alpha,N)}(r)=2[(1+\alpha)N]^{\alpha N}r^{2(\alpha N-\frac{1}{2})}e^{-(1+\alpha)Nr^{2}}$ (28)
+Numerical illustration of these results are depicted in Figures 1, 2 and 3.
+The analytical formulation of the probability density function $p_{R_{N}}$ of the scaled spectral radius $R_{N}$ does fit exactly its empirical distribution created from the generation of 1010 000000 complex induced Ginibre matrices with proportional rectangularity index $L=\alpha N$, $\forall\alpha>0$, as it is presented in Figure 1.
+As the size of these matrices increases to large numbers, the distribution of $R_{N}$ does narrow close to the outer radius $r_{out}=1$ (Figure 2).
 Figure 1.
@@
 The corresponding exact (analytical) probability density function $p_{R_{N}}$ (solid curve).
-(right panel) The analytical (exact) probability density function of scaled spectral radii $R_{N}$ for $K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=90$.
-The rectangularity index $L$ is proportional to $N$ such that $L=\alpha N$ with $\alpha>0$.
+(right panel) The analytical (exact) probability density function of the scaled spectral radius $R_{N}$ for $K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=90$.
+The rectangularity index $L$ is proportional to NN such that $L=\alpha N$ with $\alpha>0$.
 The results are presented for different values of $\alpha=\{\frac{1}{90},\frac{1}{9},\frac{4}{9}\}$.
-Graphs generated with MATLAB.
-Copyright Olivia V. Auster for code and graphs.
 Figure 2.
@@
 Limiting probability distributions of the scaled spectral radius $R_{N}$ presented with the red and black curves for large $N=10e3$ and $N=2\times 10e3$, respectively.
-Graphs generated with MATLAB.
-Copyright Olivia V.
-Auster for code and graphs.
-The exact formula of the cumulative density function of $R_{N}$ converges towards the asymptotic distribution as $N$ goes to infinity (Figure 3).
 Figure 3.
 (left panel) The exact cdf of the scaled spectral radius $R_{N}$ for $N=10e3$ (blue), $N=10e4$ (blue dotted) and $N=10e7$ (black) and $\alpha=1$.
-The asymptotic cdf of $R_{N}$ with asymptotic location and scale parameters $\mu$ and $\sigma$, respectively (red dotted curve).
+The asymptotic cdf of $R_{N}$ (red dotted curve).
 (right panel) The exact formulation of the scaled spectral radius $R_{N}$ cumulative distribution (black curve) for $N=10e7$.
-The asymptotic cumulative distribution (red curve) parametrised with asymptotic location $\mu$ and scale $\sigma$ parameters.
+The asymptotic cumulative distribution (red curve).
 The parameter $\alpha=1$ for the two curves.
-Graphs generated with MATLAB.
-Copyright Olivia V.
-Auster for code and graphs.
-The scaled minimum modulus of eigenvalues, for matrices from the complex induced Ginibre ensemble with a proportional rectangularity index, follows a Gumbel (minimum) distribution at the edge of the inner circle of the ring.
-At the inner edge of the ring, setting $a=1-\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ is an increasing function in both $x$ and $N$, it is found that
+The scaled minimum modulus of eigenvalues, for matrices from the complex induced Ginibre ensemble with a proportional rectangularity index, follows a Gumbel (minimum) distribution at the edge of the inner circle of the ring (Theorem 7).
+Setting $a=1-\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ is an increasing function in both $x$ and $N$, it is found that
 $\mathcal{P}_{N}\left(a\right)=P\left(\frac{r_{min}^{(N)}(A)}{\sqrt{\alpha N}}\geq a\right)=\prod_{k=1}^{N}P\left(\frac{1}{\alpha N}\sum_{j=1}^{k+L}Z_{j}\geq a^{2}\right)$ (29)
-where the $Z_{j}$ for $j=\{1,\cdots,k+L\}$ are independent and identically distributed random variables following a standard exponential distribution.
+where the $Z_{j}$ for $j=\{1,\cdots,k+L\}$ are independent and identically distributed random variables, each following a standard exponential distribution.
 Theorem 7.
-Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $r_{N}$ denote the scaled minimum modulus such that $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$.
+Let $G$ denote an $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$ denote the scaled minimum modulus.
+Then,
 $\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]=e^{-e^{x}}$ (30)
-in which $\gamma_{\alpha,N}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$.
-This limit corresponds to the survival distribution function of the standard Gumbel distribution for minima. The scaled minimum of moduli $r_{N}$ is approximated with the following
-$r_{N}\approx 1-T_{\alpha,N}+\xi_{\alpha,N}$ (32)
+in which $\gamma_{\alpha,N}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$. This limit corresponds to the survival distribution function of the standard Gumbel distribution for minima.
+The scaled minimum of moduli $r_{N}$ is approximated as $N$ goes to infinity, by
+$r_{N}\simeq 1-T_{\alpha,N}+\xi_{\alpha,N}$ (31)
 where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$.
-The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ where $X$ follows a standard Gumbel (minimum) distribution.
-The exact formula of the probability density function $p_{r_{N}}\left(r\right)$ of the scaled minimum modulus $r_{N}$ for proportional rectangularity index $L=\alpha N$ is derived from equation (30) and is defined as follows for any value of NN
+The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$, where $X$ follows a standard Gumbel (minimum) distribution.
+The exact formula of the probability density function $p_{r_{N}}\left(r\right)$ of the scaled minimum modulus $r_{N}$ for proportional rectangularity index $L=\alpha N,\forall\alpha>0,$ is derived from equation (29) and is defined as follows for any value of NN
 $p_{r_{N}}\left(r\right)=2(\alpha N)^{\alpha N}r^{2(\alpha N-\frac{1}{2})}e^{-\alpha Nr^{2}}\prod_{k=1}^{N}\frac{\Gamma(k+\alpha N,\alpha Nr^{2})}{\Gamma(k+\alpha N)}\sum_{j=1}^{N}\left[\frac{(\alpha N)^{j}r^{2j}}{\Gamma(j+\alpha N,\alpha Nr^{2})}\right]$ (32)
-with $\Gamma(.,.)$ the upper incomplete Gamma function.
-A detailed proof of Theorem 7 is provided in Appendix A.2.
-Numerical results are acknowledged in Figures 4 and 5.
+where $\Gamma(.,.)$ is the upper incomplete Gamma function.
+Numerical results are acknowledged in Figures 4, 5 and 6.
 The analytical formulation of the probability density function of the scaled minimum modulus $r_{N}$ fits exactly its empirical distribution from the generation of 1010 000000 complex induced Ginibre matrices (left panel) with proportional rectangularity index $L=\alpha N$, $\alpha>0$.
 As the size of these matrices increases to large numbers (e.g., $N=10e3$), the distribution of $r_{N}$ narrows around the inner radius $r_{in}=1$.
-(a) $\alpha=\frac{1}{10}$
-(b) $\alpha=\frac{1}{100}(red)\text{, }\alpha=\frac{1}{10}(blue)\text{ and }\alpha=\frac{4}{10}(black)$
 Figure 4.
@@
 The exact (analytical) probability density function is presented with the solid curve.
-(right panel) The analytical (exact) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for
-$K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=100$.
-The rectangularity index $L$ is proportional to $N$ such that $L=\alpha N$ with $\alpha>0$.
+(right panel) The analytical (exact) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated $N\times N$ matrices from the complex induced Ginibre ensemble with $N=100$.
+The rectangularity index $L$ is proportional to $N$, i.e., $L=\alpha N$ with $\alpha>0$.
 The results are presented for different values of $\alpha=\{\frac{1}{100},\frac{1}{10},\frac{4}{10}\}$.
-Graphs generated with MATLAB.
-Copyright Olivia V.
-Auster for code and graphs.
 Figure 5.
-Empirical probability distribution (histogram) and the exact (analytical) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=100$, $\alpha=\frac{1}{10}$ and a proportional rectangularity index $L=\alpha N$.
+Empirical probability distribution (histogram) and the exact (analytical) probability density function (blue curve) of the scaled minimum modulus $r_{N}$ for $K=10$ 000000 generated matrices from the complex induced Ginibre ensemble with $N=100$, $\alpha=\frac{1}{10}$ and a proportional rectangularity index $L=\alpha N,\forall\alpha>0$.
 Limiting probability distributions of the scaled minimum modulus $r_{N}$ presented with the red and black curves for large $N=10e4$ and $N=2\times 10e4$, respectively.
-Graphs generated with MATLAB. Copyright Olivia V. Auster for code and graphs.
-The exact formula of the cumulative density function of $r_{N}$ converges towards the asymptotic distribution defined with the location parameter $\mu$ and the scale parameter $\sigma$ of the Gumbel distribution derived from Equation (32), as $N$ goes to infinity (Figure 6).
+The exact formula of the cumulative distribution function of $r_{N}$ converges towards the asymptotic distribution presented in Equation (31), as $N$ goes to infinity (Figure 6).
 Figure 6.
-(left panel) The exact cdf of the scaled minimum modulus $r_{N}$ for $N=10e5$ (blue), $N=10e6$ (blue dotted) and $N=10e7$ (black) and $\alpha=1$.
-The asymptotic cdf of $r_{N}$ with asymptotic location and scale parameters (red dotted curve).
-(right panel) The exact formulation of the scaled minimum modulus $r_{N}$ cumulative distribution (black curve) for $N=10e7$ and the asymptotic cumulative distribution (red curve) parametrised with asymptotic location and scale parameters.
+(left panel) The exact cdf of the scaled minimum modulus $r_{N}$ for $N=10e5$ (blue), $N=10e6$ (blue dotted), $N=10e7$ (black) and $\alpha=1$.
+The asymptotic cdf of $r_{N}$ (red dotted curve).
+(right panel) The exact formulation of the scaled minimum modulus $r_{N}$ cumulative distribution function (black curve) for $N=10e7$ and the asymptotic cumulative distribution function (red curve).
 The parameter $\alpha=1$ for each curve.
-Graphs generated with MATLAB.
-Copyright Olivia V.
-Auster for code and graphs.
-3.2. Independence of the spectral radius and the minimum of the eigenvalue moduli for the complex induced Ginibre ensemble
+3.2.2. Independence of the spectral radius and the minimum modulus for the complex induced Ginibre ensemble
 In this section, the independence of the spectral radius $r_{max}^{(N)}(G)$ and the minimum of moduli $r_{min}^{(N)}(G)$ is demonstrated as $N$ goes to infinity for the complex induced Ginibre ensemble for fixed and proportional rectangularity indexes.
@@
 Let $\mathcal{B}_{L+N}$ and $\mathcal{B}_{L}$ denote Borel sets in the neighbourhood of $\sqrt{L+N}$ and $\sqrt{L}$, respectively.
-The events $\{r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}\}$ and $\{r_{min}^{(N)}(G)\in\mathcal{B}_{L}\}$ are independent in the limit as $N$ goes to infinity for fixed rectangularity index $L$.
+Let $r$ and $R$ denote reals in $\mathcal{B}_{L}$ and $\mathcal{B}_{L+N}$, respectively.
+The events $\{r_{max}^{(N)}(G)\leq R\}$ and $\{r_{min}^{(N)}(G)\geq r\}$ are independent in the limit as $N$ goes to infinity for fixed rectangularity index $L$.
 Proof.
-Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of a $N\times N$ complex induced Ginibre matrix $G$, respectively.
+Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of an $N\times N$ complex induced Ginibre matrix $G$, respectively.
 The result, presented in Theorem 8, is derived with the formulation of the probability,
@@
 where $r$ and $R$ are in $\mathcal{B}_{L}$ and $\mathcal{B}_{L+N}$, respectively.
-The joint probability density function of the eigenvalues $\left(\lambda_{j}\right)_{j=1}^{N}$ for the complex induced Ginibre ensemble is
-$P_{N}(\lambda_{1},\cdots,\lambda_{N})=\frac{1}{N!\pi^{N}}\prod_{k=1}^{N}\frac{1}{\Gamma(k+L)}\prod_{j<k}^{N}|\lambda_{k}-\lambda_{j}|^{2}\prod_{j=1}^{N}|\lambda_{j}|^{2L}\exp\left(-\sum_{j=1}^{N}|\lambda_{j}|^{2}\right)$
-Let $r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}$ and $r_{min}^{(N)}(G)\in\mathcal{B}_{L}$ where $\mathcal{B}_{L+N}$ is a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$.
-Applying the Andreief’s integration formula, the probability
+The joint probability density function $P_{N}(\lambda_{1},\cdots,\lambda_{N})$ of the eigenvalues $\lambda_{j},j\in\{1,...,N\}$, for the complex induced Ginibre ensemble is the one presented Section 1 with Equation (8).
+Let $r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}$ and $r_{min}^{(N)}(G)\in\mathcal{B}_{L}$, where $\mathcal{B}_{L+N}$ is a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$.
+Applying Andreief’s integration formula, the probability
 $P\left(r_{min}^{(N)}(G)\geq r\text{ and }r_{max}^{(N)}(G)\leq R\right)=P\left(r_{max}^{(N)}(G)\leq R\right)\prod_{k=1}^{N}\left[1-\frac{\gamma(k+L,r^{2})}{\gamma(k+L,R^{2})}\right]$
@@
 Theorem 9.
-Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of a $N\times N$ matrices $G$ from the complex induced Ginibre ensemble, respectively.
+Let $r_{max}^{(N)}(G)$ and $r_{min}^{(N)}(G)$ denote the spectral radius and the minimum modulus of an $N\times N$ matrix $G$ from the complex induced Ginibre ensemble, respectively.
 Let $\mathcal{B}_{L+N}$ denote a Borel set in the neighbourhood of $\sqrt{L+N}$ and $\mathcal{B}_{L}$ a Borel set in the neighbourhood of $\sqrt{L}$, for proportional rectangularity index $L$.
-The events $\{r_{max}^{(N)}(G)\in\mathcal{B}_{L+N}\}$ and $\{r_{min}^{(N)}(G)\in\mathcal{B}_{L}\}$ are independent, where $L=\alpha N$ with $\alpha>1$, in the limit as $N$ goes to infinity.
+Let $r$ and $R$ denote reals in $\mathcal{B}_{L}$ and $\mathcal{B}_{L+N}$.
+The events $\{r_{max}^{(N)}(G)\leq R\}$ and $\{r_{min}^{(N)}(G)\geq r\}$ are independent, where $L=\alpha N$ with $\alpha>1$, in the limit as $N$ goes to infinity.
 Proof.
@@
 $\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\text{ and }r_{max}^{(N)}(G)\leq R\right)$ $=P\left(r_{max}^{(\infty)}(G)\leq R\right)\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$
-The asymptotics of the regularised lower incomplete Gamma functions is presented in [24]. It is, uniformly for $x\geq 0$, and $a\rightarrow+\infty$ and/or $x\rightarrow+\infty$,
-$P(a,x)=\frac{\gamma(a,x)}{\Gamma(a)}=\frac{1}{\Gamma(a)}\int_{0}^{x}t^{a-1}e^{-t}dt=\frac{1}{2}\operatorname{erfc}\left(-\eta(a/2)^{1/2}\right)-\mathcal{R}_{a}(\eta)$
-where $\eta=\left[2(\mu-\log(1+\mu))\right]^{1/2}$ with $\mu=\lambda-1$ and $\lambda=\frac{x}{a}$
-The remainder $\mathcal{R}_{a}(\eta)$ is represented as
-$\mathcal{R}_{a}(\eta)\sim(2\pi a)^{-1/2}e^{-\frac{1}{2}a\eta^{2}}\sum_{k=0}^{+\infty}c_{k}(\eta)^{-k}$
-with the coefficients $c_{k}$ defined as in [24] and [25]. These results are (uniformly) valid for $\eta\in\mathbb{R}$.
 Remark 3.1.
-$\gamma(k+\alpha N,R^{2})=\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)$
+For $R$ in $\mathcal{B}_{L+N}$,
+$\gamma(k+\alpha N,R^{2})=\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)$ (33)
 where
@@
 Then,
-$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$
-$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)}\right]=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}\right)\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$
+$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)-\varepsilon_{k,N}(R)}\right]$
+$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}\right)\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$
 where $E_{k,N}(R,r)=\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}(R)}{\Gamma(k+\alpha N)^{2}}+O\left(\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}^{2}(R)}{(\Gamma(k+\alpha N))^{3}}\right)$ where $\lim_{N\longrightarrow+\infty}\varepsilon_{k,N}(R)=0$.
-Furthermore, using the results in [24] and [25] where $\eta=\left[2(\mu-\log(1+\mu))\right]^{1/2}$ with $\mu=\frac{r^{2}}{k+\alpha N}-1$
-$\lim_{N\rightarrow+\infty}\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}$ $=\lim_{N\rightarrow+\infty}\left[\frac{1}{2}\operatorname{erfc}\left(-\eta((k+\alpha N)/2)^{1/2}\right)-\mathcal{R}_{t}(\eta)\right]$
-$=1-\lim_{N\rightarrow+\infty}\mathcal{R}_{t}(\eta)$
-where $t=k+\alpha N$.
-The limit of the partial product $\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ is established from the limit of its lower and upper bounds.
-Lower bound. The limit
-$\lim_{N\rightarrow+\infty}1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}=\lim_{N\rightarrow+\infty}\mathcal{R}_{t}(\eta)$
-And for an appropriate choice of $r$, i.e., $r=O\left(\sqrt{\alpha N}\right)$, the parameter $\eta$ is not infinite.
-This implies that the remainder $\mathcal{R}_{t}(\eta)$ is finite, i.e., $\frac{1}{\mathcal{R}_{t}(\eta)}>C$ where $C$ is a positive constant.
-This implies that, for $N$ large enough and finite,
-$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $>\prod_{k=1}^{N}\left(1+E_{k,N}(R,r)C\right)>\prod_{k=1}^{N}\left(1+\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}(R)C}{N\Gamma(k+\alpha N)^{2}}\right)=1+O\left(\varepsilon_{k,N}(R)\right)$
-Upper bound. Using the Stirling formula,
-$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $<1+O\left(\frac{\varepsilon_{k,N}(R)e^{-\alpha N}}{\left(N\right)^{\alpha N-\frac{1}{2}-1}}\right)$
+The limit of the partial product $\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ is established from the limit of its lower and upper bounds, as $N$ goes to infinity.
+Lower bound.
+For any fixed $k$ and $N$, $\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}<1$.
+Let $C$ denote a positive constant such that for any fixed $k$ and $N$, $\frac{1}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}>C$ .
+This implies that,
+$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $>\prod_{k=1}^{N}\left(1+E_{k,N}(R,r)C\right)$
+$>\prod_{k=1}^{N}\left(1+\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}(R)C}{N\Gamma(k+\alpha N)^{2}}\right)$
+$>\prod_{k=1}^{N}\left(1+\frac{\gamma(k+\alpha N,r^{2})\varepsilon_{k,N}(R)C}{N\Gamma(k+\alpha N)\Gamma((1+\alpha)N)}\right)=1+O\left(\frac{\varepsilon_{k,N}(R)}{\Gamma((1+\alpha)N)}\right)$
+Upper bound.
+$\prod_{k=1}^{N}\left(1+\frac{E_{k,N}(R,r)}{1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}}\right)$ $<\prod_{k=1}^{N}\left(1+E_{k,N}(R,r)\right)$
+$<\prod_{k=1}^{N}\left(1+\frac{\varepsilon_{k,N}(R)}{\Gamma(k+\alpha N)}\right)$
+$<\prod_{k=1}^{N}\left(1+\frac{\varepsilon_{k,N}(R)}{(N-1)!}\right)=1+O\left(\frac{\varepsilon_{k,N}(R)}{N^{-1}(N-1)!}\right)$
 where $\alpha>1$.
@@
 Consequently,
-$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$ $=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}\right]$
-$=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}$
-$=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)=P\left(r_{min}^{(\infty)}(G)\geq r\right)$
+$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\left[1-\frac{\gamma(k+\alpha N,r^{2})}{\gamma(k+\alpha N,R^{2})}\right]$ $=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+\alpha N,r^{2})}{\Gamma(k+\alpha N)}=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)$
 Finally,
 $\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\text{ and }r_{max}^{(N)}(G)\leq R\right)$ $=P\left(r_{min}^{(\infty)}(G)\geq r\right)P\left(r_{max}^{(\infty)}(G)\leq R\right)$
+Theorem 10 (Independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble).
+Let $G$ denote an $N\times N$ matrix from the complex induced Ginibre ensemble with proportional rectangularity index $L=\alpha N$, $\alpha=O\left(N^{m}\right),m>1$.
+The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ are independent random variables, under scaling $\sqrt{\alpha N}$, as $N$ goes to infinity.
+More precisely, let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$. Setting $\rho=\sqrt{\frac{1+\alpha}{\alpha}}$,
+$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\xi_{\alpha}^{(N)}(y)\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\rho^{2}\alpha N}}+\eta_{\alpha}^{(N)}(x)\right]=e^{-e^{y}}e^{-e^{-x}}$
+with $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$ and where
+$\xi_{\alpha}^{(N)}(y)=\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }\eta_{\alpha}^{(N)}(x)=\frac{x}{2\rho\sqrt{2\alpha N\gamma_{\alpha,N}}}$
+The scaled spectral radius $R_{N}$ and the scaled minimum of moduli $r_{N}$ are approximated as $N$ goes to infinity, by
+$R_{N}$ $\simeq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\rho^{2}\alpha N}}+\frac{X}{2\rho\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }r_{N}\simeq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{Y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
+where $X$ and $Y$ are standard Gumbel (maximum) and standard Gumbel (minimum) distributed random variables, respectively.
+The proof of Theorem 10 is detailed in the Appendix A.3.
+3.3. Limiting left and right tail distributions of the minimum modulus for matrices from the complex Ginibre and complex induced ensembles. A comparison.
+The limiting left and right tail probability distributions of the minimum modulus for matrices from the complex Ginibre and complex induced Ginibre ensembles are investigated in this section.
+Derived in Section 3.1, the survival probability distribution function of the minimum modulus $r_{min}^{(N)}(A)$ of an $N\times N$ complex Ginibre matrix $A$, is used to determine the corresponding analytical probability density function.
+As derived in Section 3.1,
+$p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]$ (34)
+It represents the exact formulation of the probability density function of the minimum modulus $r_{min}^{(N)}(A)$.
 Theorem 11.
-Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with proportional rectangularity index $L=\alpha N$, $\alpha>0$.
-The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ are independent random variables, under scaling $\sqrt{\alpha N}$, as $N$ goes to infinity.
-More precisely, let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$.
-Setting $\rho=\sqrt{\frac{1+\alpha}{\alpha}}$,
-$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\xi_{\alpha}^{(N)}(y)\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\eta_{\alpha}^{(N)}(x)\right]=e^{-e^{y}}e^{-e^{-x}}$
-in which $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$ and where
-$\xi_{\alpha}^{(N)}(y)=\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }\eta_{\alpha}^{(N)}(x)=\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
-The scaled spectral radius $R_{N}$ and the scaled minimum of moduli $r_{N}$ are approximated as
-$R_{N}$ $\simeq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }r_{N}\simeq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{Y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
-where $X$ and $Y$ are standard Gumbel (maximum) and standard Gumbel (minimum) distributed random variables, respectively.
-The proof of Theorem 10 is detailed in Appendix A.3.
-4. Limiting distribution of the minimum modulus of eigenvalues for matrices from the complex Ginibre and complex induced ensembles. A comparison
-The limiting probability distributions of the minimum modulus of eigenvalues for matrices from the complex Ginibre and complex induced Ginibre ensembles are investigated in this section.
-Derived in Section 2, the survival probability distribution function of the minimum modulus $r_{min}^{(N)}(A)$ of a $N\times N$ complex Ginibre matrix $A$, is used to determine the corresponding analytical probability density function, as $N$ goes infinity.
-As derived in Section 2,
-$p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]$
-It represents the exact formulation of the probability density function of the $r_{min}^{(N)}(A)$ as the product of a partial product of the regularised upper incomplete Gamma functions and a partial sum.
+Let $A$ denote an $N\times N$ random matrix from the complex Ginibre ensemble.
+The limiting left tail probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is the probability density function of the Rayleigh distribution with parameter $\frac{1}{\sqrt{2}}$, as $N$ goes to infinity.
+More precisely, for $0<r\ll 1$,
+$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\left(1+O(r^{2})\right)$ (35)
+Proof.
+The limit of partial product $\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$, as $N$ goes to infinity, has been investigated in Section 3.1 for the complex Ginibre ensemble.
+For $0<r\ll 1$, it is expressed as
+$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$
+The upper incomplete Gamma function $\Gamma(k,x)$, $k\in\mathbb{N}\setminus\{0\}$, has asymptotic expansion
+$\Gamma(k,x)\sim\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$, as $x\longrightarrow 0^{+}$. The limit of the partial sum is derived as follows.
+Knowing that, for $0<r\ll 1$,
+$\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2(j+1)}\right)$
+This implies,
+$\lim_{N\rightarrow+\infty}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$ $=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{r^{2j}}{j!+O(r^{2(j+1)})}=2r\left(1+O(r^{2})\right)$
+Thus, the limiting left tail probability density function of the minimum modulus for $N\times N$ matrices from the complex Ginibre ensemble is
+$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)$ $=2re^{-r^{2}}\left(1+O(r^{2})\right)$
+This function corresponds to the probability density function of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
+This result is in line with the result acknowledged in Theorem 3.
+Now, considering any rectangularity index $L\geq 0$, the limiting left tail distribution of the minimum modulus $r_{min}^{(N)}(G)$ for an $N\times N$ random matrix $G$ from the complex induced Ginibre ensemble, as $N$ goes to infinity, is derived as follows.
 Theorem 12.
-Let $A$ denote a $N\times N$ random matrix from the complex Ginibre ensemble.
-The limiting left tail probability density function of the minimum modulus $r_{min}^{(N)}(A)$ is the probability density function of the Rayleigh distribution with parameter $\frac{1}{\sqrt{2}}$, as $r$ goes to zero and $N$ goes to infinity.
-More precisely, for $0<r\ll 1$,
-$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (34)
-Proof.
-The limit of partial product $\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}$, as $N$ goes to infinity, has been investigated in Section 2 for the complex Ginibre ensemble, and for $0<r\ll 1$, it is expressed as
-$\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$
-The upper incomplete Gamma function $\Gamma(k,x)$, for small $x$ and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
-The limit of the partial sum is derived as follows. Knowing that, for $0<r\ll 1$,
-$\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$
-$\lim_{N\rightarrow+\infty}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$ $=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{r^{2j}}{\Gamma(j+1,r^{2})}=2e^{-r^{2}}r\sum_{j=0}^{+\infty}\frac{(r^{2})^{j}}{j!}=2e^{-r^{2}}re^{r^{2}}=2r$
-Thus, the limiting left tail probability density function of the minimum modulus for matrices from the complex Ginibre ensemble is
-$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)$ $=\lim_{N\rightarrow+\infty}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{1}{\Gamma(j+1,r^{2})}2r^{2j+1}e^{-r^{2}}\right]$
-$=2re^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$
-This function corresponds to the probability density function of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
-This result is in line with the result acknowledged with Theorem 3.
-The probability density function of the minimum modulus of eigenvalues for $N\times N$ matrices $G$ from the complex induced Ginibre ensemble is presented in Section 3.
-Its limit is established here for the particular case where the fixed rectangularity index $L=1$.
-Theorem 12.
-The minimum modulus of a $N\times N$ matrix $G$ from the complex induced Ginibre ensemble, with rectangularity index $L=1$, has a left tail probability density function of the Generalised Gamma distribution with scale parameter $\alpha=1$ and other parameters $\kappa=2$ and $\gamma=4$.
-The corresponding limiting probability density function, for $0<r\ll 1$, as $N$ goes to infinity, is
-$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2r^{3}e^{-r^{2}}\left(1-O\left(r^{4}\right)\right)$ (35)
-Proof.
-Let $G$ denote a $N\times N$ complex induced Ginibre matrix.
-The limiting probability density function of the minimum modulus of eigenvalues, denoted $r_{min}^{(N)}(G)$, as $r$ approaches zero and as $N$ goes to infinity, is
-$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=2re^{-r^{2}}\lim_{N\rightarrow+\infty}\left(\prod_{k=1}^{N}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=1}^{N}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]\right)$
-Furthermore, as $N$ goes to infinity and $r$ approaches zero,
-$\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}=1-O\left(r^{4}\right)$
-This is established with the results presented in Section 2. The partial product is extended from N1N-1 to $N$ which does not change the limit.
-Also, the upper incomplete Gamma function $\Gamma(k,x)$, as $x$ approaches zero and $k\in\mathbb{N}\setminus\{0\}$, is $\Gamma(k,x)=\Gamma(k)-\sum_{n=0}^{\infty}(-1)^{n}\frac{x^{k+n}}{n!(k+n)}$.
-For $r$ close to zero, $\Gamma(j+1,r^{2})=\Gamma(j+1)+O\left(r^{2}\right)$. This implies that the limit of the partial sum, as $r$ approaches zero and $N$ goes to infinity, is
-$\lim_{N\rightarrow+\infty}2re^{-r^{2}}\sum_{j=1}^{N}\left[\frac{r^{2j}}{\Gamma(j+1,r^{2})}\right]\\ =2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{\Gamma(j+1,r^{2})}\times\left(1-O\left(r^{4}\right)\right)=2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{j!}\times\left(1-O\left(r^{4}\right)\right)$
-Finally,
-$\lim_{N\rightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)$ $=2re^{-r^{2}}\sum_{j=1}^{+\infty}\frac{(r^{2})^{j}}{j!}\times\left(1-O\left(r^{4}\right)\right)$
-$=2r^{3}e^{-r^{2}}\times\left(1-O\left(r^{4}\right)\right)=\frac{\frac{\kappa}{\alpha^{\gamma}}}{\Gamma(\gamma/\kappa)}r^{\gamma-1}e^{-(r/\alpha)^{\kappa}}\times\left(1-O\left(r^{4}\right)\right)$
-This corresponds to the probability density function of the Generalised Gamma distribution with scale parameter $\alpha=1$ and other parameters $\kappa=2$ and $\gamma=4$, where $\Gamma(.)$ is the Gamma function.
-Now, considering any rectangularity index $L\geq 0$, the limiting left tail distribution of the minimum of eigenvalues moduli for $N\times N$ random matrices $G$ form the complex induced Ginibre ensemble, as $N$ goes to infinity, is presented as follows.
-Theorem 13.
-Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L\geq 0$.
-The left tail distribution of minimum of the moduli $r_{min}^{(N)}(G)$ is the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$ in the limit as $N$ goes to infinity and $r$ approaches zero,
+Let $G$ denote an $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L\geq 0$.
+The left tail distribution of minimum modulus $r_{min}^{(N)}(G)$ is the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$ in the limit as $N$ goes to infinity.
+More precisely, as $r$ approaches zero,
 $\lim_{N\longrightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-\exp\left[-\frac{r^{2(L+1)}}{(L+1)!}\right]+O\left(r^{4(L+1)}\right)$ (36)
 Proof.
-This result is derived using the same methodological approach presented in section 2.1.
+This result is derived using the same methodological approach as presented in section 3.1.1.
 For $0<r\ll 1$,
@@
 $\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-\exp\left[-\frac{r^{2(L+1)}}{(L+1)!}\right]+O\left(r^{4(L+1)}\right)$
-This corresponds to the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter $\lambda=\left((L+1)!\right)^{1/k}$. ∎
+This corresponds to the Weibull distribution with shape parameter $k=2(L+1)$ and scale parameter
+$\lambda=\left((L+1)!\right)^{1/k}$. ∎
 Corollary 13.
 Setting the rectangularity index $L$ to zero, the limiting distribution of the minimum modulus $r_{min}^{(N)}(G)$ for the complex induced Ginibre ensemble corresponds to the limiting distribution of the minimum modulus $r_{min}^{(N)}(A)$ for the complex Ginibre ensemble, i.e., for $0<r\ll 1$,
-$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-r^{2}}+O(r^{4})=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)<r\right)$
+$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-r^{2}}+O(r^{4})$ (37)
+Lemma 14.
+The probability density function of the minimum modulus $r_{min}^{(N)}(G)$ of matrices from the complex induced Ginibre ensemble, with rectangularity index $L\geq 0$, is
+$p_{r_{min}^{(N)}(G)}\left(r\right)=2\frac{e^{-r^{2}}}{r}\prod_{k=1}^{N}\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\sum_{j=1}^{N}\left[\frac{r^{2(j+L)}}{\Gamma(j+L,r^{2})}\right]$ (38)
 Corollary 15.
 For $0<r\ll 1$, as $N$ goes to infinity, and for a rectangularity index $L=0$,
-$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)-O\left(r^{5}\right)$
-with shape parameter $\kappa=2$ and scale parameter $\lambda=1$. This does correspond to the probability density function of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
+$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\frac{\kappa}{\lambda}\left(\frac{r}{\lambda}\right)^{\kappa-1}e^{-(r/\lambda)^{\kappa}}\left(1+O(r^{2})\right)$ (39)
+with Weibull distribution parameters $\kappa=2$ and $\lambda=1$. This corresponds to the probability density function of the Rayleigh distribution with parameter $\sigma=\frac{1}{\sqrt{2}}$.
 Proof.
-For $0<r\ll 1$,
-$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)=\frac{\kappa}{\lambda}\left(\frac{r}{\lambda}\right)^{\kappa-1}e^{-(r/\lambda)^{\kappa}}+O\left(r^{3\kappa+1}\right)=2re^{-r^{2}}+O\left(r^{7}\right)$
-where $\kappa=2$ and scale parameter $\lambda=1$. ∎
-The limiting distribution of the minimum modulus for the complex induced Ginibre ensemble, in the limit as $N$ goes to infinity for fixed rectangularity index $L\in\left[0,\varepsilon\right]$ (with $\varepsilon\ll 1$), is derived using the Euler-Maclaurin summation formula and the asymptotic of the upper incomplete Gamma function as $r$ goes to infinity [26].
+Let $G$ denote a $N\times N$ complex induced Ginibre matrix with rectangularity index $L=0$.
+For $0<r\ll 1$, from Lemma 14 and the methodological approach used to prove Theorem 11,
+$\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(G)}\left(r\right)$ $=\frac{2e^{-r^{2}}}{r}\prod_{k=0}^{N-1}\frac{\Gamma(k+1,r^{2})}{\Gamma(k+1)}\sum_{j=0}^{N-1}\left[\frac{r^{2(j+1)}}{\Gamma(j+1,r^{2})}\right]$
+$=\frac{\kappa}{\lambda}\left(\frac{r}{\lambda}\right)^{\kappa-1}e^{-(r/\lambda)^{\kappa}}\left(1+O(r^{2})\right)$
+$=\lim_{N\longrightarrow+\infty}p_{r_{min}^{(N)}(A)}\left(r\right)$
+with Weibull distribution parameters $\kappa=2$ and $\lambda=1$.
+The right tail limiting distribution of the minimum modulus for the complex induced Ginibre ensemble, in the limit as $N$ goes to infinity for fixed rectangularity index $L\in\left[0,\varepsilon\right]$ (with $\varepsilon\ll 1$), is derived using the Euler-Maclaurin summation formula and the asymptotic expansion of the upper incomplete Gamma function as $r$ goes to infinity.
 Thus,
-$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)=\lim_{N\rightarrow+\infty}\prod_{k=1}^{N}\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}=\exp{\left[\sum_{k=1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)\right]}$
-Remark 4.1.
-It is notified in [25], that
-$\lim_{k\longrightarrow+\infty}\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}=\frac{1}{2}\operatorname{erfc}\left(\eta\left(k+L\right)^{1/2}\right)+O\left((k+L)^{-1/2}e^{-\frac{1}{2}(k+L)\eta^{2}}\right)$ (37)
-where $\operatorname{erfc}$ is the complementary error function, $\lambda=\frac{r^{2}}{k+L}$, $\mu=\lambda-1$ and $\eta=\left(2\left[\mu-\log\left(1+\mu\right)\right]\right)^{1/2}$.
-From Remark 4.1,
-$\lim_{k\longrightarrow+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)=0$
-Now, the asymptotic expansion of the upper incomplete Gamma function (cf. [26]) as $z$ goes to infinity, is
-$\lim_{z\longrightarrow+\infty}\Gamma(a,z)=z^{a-1}e^{-z}+O(z^{a-2}e^{-z})$
+$\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)\geq r\right)=\exp{\left[\sum_{k=1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)\right]}$
+Now, the asymptotic expansion of the upper incomplete Gamma function (cf. [26]) is
+$\Gamma(a,z)=z^{a-1}e^{-z}+O(z^{a-2}e^{-z})\text{, as }z\text{ goes to infinity}$
 Thus, for large $r$,
-$\frac{1}{2}\log\left(\frac{\Gamma(L+1,r^{2})}{\Gamma(L+1)}\right)=\frac{1}{2}\left[-r^{2}+\log\left(\frac{r^{2L}\left(1+O\left(\frac{1}{r^{2}}\right)\right)}{\Gamma(L+1)}\right)\right]$ (38)
+$\frac{1}{2}\log\left(\frac{\Gamma(L+1,r^{2})}{\Gamma(L+1)}\right)=\frac{1}{2}\left[-r^{2}+\log\left(\frac{r^{2L}\left(1+O\left(\frac{1}{r^{2}}\right)\right)}{\Gamma(L+1)}\right)\right]$
 A precise evaluation of the integral $\int_{1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)dk$ is $\frac{r^{4}}{4}+O(r^{2})$.
@@
 $\sum_{k=1}^{+\infty}\log\left(\frac{\Gamma(k+L,r^{2})}{\Gamma(k+L)}\right)=-\frac{r^{4}}{4}+\frac{1}{2}\left[-r^{2}+\log\left(\frac{r^{2L}\left(1+O\left(\frac{1}{r^{2}}\right)\right)}{\Gamma(L+1)}\right)\right]+O\left(r^{2}\right)$
-This implies that for large $r$,
+This implies that, for large $r$,
 $\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-\frac{r^{4}}{4}}\exp\left(-\frac{r^{2}}{2}+\frac{1}{2}\log\left(\frac{r^{2L}+r^{2(L-1)}O\left(1\right)}{\Gamma(L+1)}\right)+O\left(r^{2}\right)\right)$ (40)
-For a rectangularity index $L=0$, for large $r$, and $A$ a $N\times N$ complex Ginibre matrix,
+For a rectangularity index $L=0$, for large $r$, and $A$ an $N\times N$ complex Ginibre matrix,
 $\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(G)<r\right)=1-e^{-\frac{r^{4}}{4}\left(1+O\left(\frac{1}{r^{2}}\right)\right)}=\lim_{N\rightarrow+\infty}P\left(r_{min}^{(N)}(A)<r\right)$ (41)
 4. Conclusions
-The distribution of the minimum modulus of matrices from the complex Ginibre ensemble is analytically derived with the use of the Andreief’s integration formula.
-It does correspond to a $N$-th partial product of regularised upper incomplete Gamma functions.
-The left and right tail asymptotic distributions of this extreme modulus have also been investigated for large sizes of complex Ginibre matrices.
+This research was conducted from pure academic interests in the field of Random Matrix Theory and Probability theory.
+The novelty is in the new derived results, providing a clear understanding of the dynamics of random variables for which little is present in the literature.
+The distribution of the minimum modulus of matrices from the complex Ginibre ensemble is analytically derived with the use of Andreief’s integration formula [8].
+It corresponds to an $N$-th partial product of regularised upper incomplete Gamma functions as in references [14, 7, 2].
+The scaled extreme moduli for the complex induced Ginibre ensemble are identified to follow a Gumbel distribution in the limit of large random matrices.
+More precisely, their asymptotic distribution is well approximated as a shifted and rescaled standard Gumbel distribution, as it is relevant for the study of physical phenomena in Nuclear Physics [31], where populations of levels of heavy nuclei are of order $10^{6}$.
+The stochastic modelling of these random variables might reveal more accurate results while taking into account their mean and second central moment within a similar methodological framework as the one established in reference [22].
+The analytical formulation of the latter statistical features might have its relevance.
+Derived for the non-Hermitian ensembles considered in this paper, the analytical formulations of the probability density functions of the extreme eigenvalue moduli (spectral radius and minimum modulus with and without scaling) exactly coincide with their empirical counterpart sampled from thousands of matrices.
+The independence of the minimum modulus and the spectral radius is a trivially accepted fact for which no formal proof was detailed for the complex Ginibre ensemble, as well as for its random matrix ensemble generalisation.
+The independence of these random variables has been proved in the present paper for these two random matrix ensembles at appropriate scalings.
+And it is acknowledged to hold as $N$ goes to infinity.
+The left and right tail asymptotic distributions of the minimum modulus have also been investigated for large sizes of complex Ginibre matrices.
 They are the Rayleigh distribution with parameter $\sigma=1/\sqrt{2}$ and the Weibull distribution with shape parameter $\kappa=4$ and scale parameter $\lambda=k^{1/k}$, respectively.
-Derived for the non-Hermitian ensembles considered in this paper, the analytical formulation of the probability density functions of the extreme eigenvalue moduli (spectral radius and minimum modulus with and without scaling) exactly coincides with its empirical counterpart sampled from thousands of matrices.
-The independence of these random variables has also been established for these two random matrix ensembles at appropriate scalings.
-And it is acknowledged to hold as $N$ goes to infinity.
-As final new results presented here, the minimum moduli of the complex Ginibre ensemble and the complex induced ensemble, for values of rectangularity indexes close to zero, have the same limiting left and right tail distributions, the Rayleigh and the Weibull distributions, respectively.
-The exact limiting distribution of the minimum modulus for the complex Ginibre ensemble, as $N$ goes to infinity, is absent from the literature.
+Limiting left and right tail distributions of the minimum modulus provide insight into the exact nature of its distribution for the complex Ginibre ensemble, in the limit as $N$ goes to infinity.
+They are in accordance with known features of the complex Ginibre ensemble and the complex induced Ginibre ensemble for a rectangularity index equal to zero.
+It is acknowledged that the minimum moduli of the complex Ginibre ensemble and the complex induced Ginibre ensemble have the same limiting left and right tail distributions, for a rectangularity index at zero, which are the Rayleigh and the Weibull distributions, respectively.
+The limiting distribution of the minimum modulus for the complex Ginibre ensemble, as $N$ goes to infinity, is absent from the literature.
+A mixture of distributions involving either the Rayleigh distribution or the Weibull distribution might define the right statistics of this minimum.
 This is left for forthcoming research.
-Appendix A Proof of the Limit theorems for the complex induced Ginibre ensemble
+Appendix A Proofs of limit theorems for the complex induced Ginibre ensemble
 A.1. Proof of Theorem 6
-For the convenience of the reader, Theorem 6 is restated here.
-Theorem (Limiting distribution of the scaled spectral radius $R_{N}$ for the complex induced Ginibre ensemble).
-Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}$ denote the scaled spectral radius.
-Then,
-$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[R_{N}\leq 1+\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}+\frac{x}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}\right]=e^{-e^{-x}}$ (41)
-where $\gamma_{\alpha,N}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$.
-This limiting probability distribution is the standard Gumbel distribution for maxima. The corresponding cumulative density function is the function $F_{X}^{(\text{Gumbel max})}(x)=\exp(-\exp(-x))$.
-The scaled spectral radius $R_{N}$ is approximated as
-$R_{N}\simeq 1+T_{\alpha,N}+\xi_{\alpha,N}$ (42)
-where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$.
-The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ with $X$ following a standard Gumbel (maximum) distribution.
-Proof.
-At the outer edge of the ring, i.e., $r_{out}=1$,
+At the outer edge of the ring,
 $P\left(\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}\leq a\right)$ $=\frac{1}{\pi^{N}\prod_{k=1}^{N}\Gamma(k+L)}\prod_{k=0}^{N-1}\pi\int_{0}^{(1+\alpha)Na^{2}}t^{k+L}e^{-t}dt$
@@
 $P\left(\frac{r_{max}^{(N)}(G)}{\sqrt{(1+\alpha)N}}\leq a\right)$ $=\prod_{k=1}^{N}P\left(\frac{1}{(1+\alpha)N}\sum_{j=1}^{N+L-k+1}Z_{j}\leq a^{2}\right)$
-where the $Z_{j}$ for $j=\{1,\cdots,N-k+L+1\}$ are independent and identically distributed random variables following an exponential distribution with parameter 11 $\left(i.e.,Z_{j}\sim\exp(1)\right)$.
-Near the edge of the outer circle of the ring, let $a=\frac{r_{out}}{\sqrt{(1+\alpha)N}}+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}$ denote the scaled radius close to $r_{out}$ with appropriate scaling $\sqrt{(1+\alpha)N}$.
+where the $Z_{j}$ for $j=\{1,\cdots,N-k+L+1\}$ are independent and identically distributed random variables, each following an exponential distribution with parameter 11 $\left(i.e.,Z_{j}\sim\exp(1)\right)$.
+Near the edge of the outer circle of the ring, let $r_{out}=\sqrt{(1+\alpha)N}$ and $a=\frac{r_{out}}{\sqrt{(1+\alpha)N}}+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}$.
 The probability
@@
 The function $f_{N}(x)=o(\sqrt{(1+\alpha)N})$, meaning that $\lim_{N\rightarrow+\infty}\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}=0$.
-The probability $\mathcal{P}_{N}(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}})$ is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf.
+The probability $\mathcal{P}_{N}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)$ is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf.
 [22]).
@@
 $\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}p_{k}\leq\mathcal{P}_{N}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)\leq\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$
-The following is derived using the same arguments as stated in [20] and applying the Markov inequality and the definition of the quantile of the exponential distribution.
+The following is derived using the same arguments as stated in [22] and applying the Markov inequality as well as the definition of the quantile of the Exponential distribution.
 The function $\phi_{N}(x)$ is a positive and increasing function.
@@
 The random variables $Z_{j}$, $i\in\{1,\cdots,(1+\alpha)N-k\}$ are independent and identically distributed.
-Applying the Markov inequality, with $0<\eta<1$ , as the exponential is an strictly increasing and convex function, this implies that
+Applying the Markov inequality, with $0<\eta<1$, and as the exponential is an strictly increasing and convex function, this implies that
 $\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}P\left(\frac{1}{\sqrt{(1+\alpha)N}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
@@
 $=e^{-\eta(1+\alpha)N(\frac{k}{(1+\alpha)N}-1)-\left((1+\alpha)N-k\right)\ln(1-\eta)}>1$
-And with $0<\eta<1$ and $\forall Y\in\mathbb{R}$
+And with $0<\eta<1$ and  $\forall\text{ }Y\in\mathbb{R}$,
 $e^{Y-\eta\sqrt{(1+\alpha)N}\phi_{N}(x)}<e^{Y}\Rightarrow 1-e^{Y-\eta\sqrt{(1+\alpha)N}\phi_{N}(x)}>1-e^{Y}$
@@
 $\prod_{k=(1+\alpha)N\delta_{N}}^{N-1}p_{k}\geq\left(1-e^{-(1+\alpha)N\delta_{N}^{2})}\right)^{N}$
-The parameter $\delta_{N}$ is chosen such that $\left(1-e^{-(1+\alpha)N\delta_{N}^{2})}\right)^{N}=\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$ and is defined as follows
-$\delta_{N}=\sqrt{\frac{2\log{N}}{(1+\alpha)N}}$
+The parameter $\delta_{N}$ is chosen such that $\left(1-e^{-(1+\alpha)N\delta_{N}^{2})}\right)^{N}=\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$,
+i. e., $\delta_{N}=\sqrt{\frac{2\log{N}}{(1+\alpha)N}}$.
 Consequently,
@@
 $=\prod_{k=0}^{(1+\alpha)N\delta_{N}}P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\left(\sqrt{\frac{(1+\alpha)N}{N+L-k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)\right)$
-$=\prod_{k=0}^{(1+\alpha)N\delta_{N}}P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
-$=\prod_{k=0}^{(1+\alpha)N\delta_{N}}P\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)<\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
 $=\prod_{k=0}^{(1+\alpha)N\delta_{N}}p_{k}$
 where
-$p_{k}=P\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)\leq\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)$
-As in the reference [20], it is assumed that $k=o((1+\alpha)N)$.
-The factor $\sqrt{\frac{(1+\alpha)N}{N+L-k}}$ is assumed equal (approximatively) to one for small and moderate values of $k$.
-For large values of $k$, this factor is rather large and the probability
-$P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\left(\sqrt{\frac{(1+\alpha)N}{N+L-k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)\right)$ is approximatively equal to one.
-This means that its logarithm is zero.
+$p_{k}=P\left(\frac{1}{\sqrt{N+L-k}}\sum_{j=1}^{N+L-k}(Z_{j}-1)\leq\left(\sqrt{\frac{(1+\alpha)N}{N+L-k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}\right)\right)$
+and, for fixed $k\in\left[0,(1+\alpha)N\delta_{N}\right]$, the factor $\sqrt{\frac{(1+\alpha)N}{N+L-k}}$ goes to one as $N$ goes to infinity.
 Finally,
@@
 $\frac{\sqrt{(1+\alpha)N-k}(\bar{Z}-\mu)}{\sigma}$ $=\frac{1}{\sqrt{(1+\alpha)N-k}}\sum_{j=1}^{(1+\alpha)N-k}(Z_{j}-1)$
-where each element of sequence $Z_{1},\cdots,Z_{(1+\alpha)N-k}$ is i.i.d. exponentially distributed with parameter equal to $1$.
+where each element of sequence $Z_{1},\cdots,Z_{(1+\alpha)N-k}$ is i.i.d.
+exponentially distributed with parameter equal to $1$.
 The random variable $\bar{Z}=\frac{1}{(1+\alpha)N-k}\sum_{j=1}^{(1+\alpha)N-k}Z_{j}$ is the empirical mean.
-The mean of $\bar{Z}$ is $\mu=1$ and its standard deviation is $\frac{\sigma}{\sqrt{(1+\alpha)N-k}}=\frac{1}{\sqrt{(1+\alpha)N-k}}$ with $\sigma=1$.
+The mean of $\bar{Z}$ is equal to $1$ and its standard deviation is $\frac{\sigma}{\sqrt{(1+\alpha)N-k}}=\frac{1}{\sqrt{(1+\alpha)N-k}}$ with $\sigma=1$.
 Applying the Edgeworth expansion, the logarithm of the probability $p_{k}$ is then
@@
 $+\log\left[1+\begin{pmatrix}O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)\\ +O\left(\frac{1}{(1+\alpha)N-k}\right)+O\left(K_{N}((1+\alpha)N-k)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)\end{pmatrix}\right$]
-$=\log\left[\int_{-K_{N}}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)$
-$+O\left(\frac{1}{(1+\alpha)N-k}\right)+O\left(K_{N}((1+\alpha)N-k)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)$
+$=\log\left[\int_{-K_{N}}^{\phi_{N}(x)+\frac{k}{\sqrt{(1+\alpha)N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{(1+\alpha)N}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)$
+$+O\left(\frac{1}{(1+\alpha)N}\right)+O\left(K_{N}((1+\alpha)N)^{-3/2}\right)+O\left(e^{-\frac{K_{N}^{2}}{2}}\right)$
 As mentioned in [22], the lower limit of integration in the leading term of the logarithm of the probability $p_{k}$ is extended from $K_{N}$ down to -$\infty$. Defining the function $f(k)$ as
@@
 $\sum_{k=0}^{\sqrt{2(1+\alpha)N\log N}}f(k)=\sqrt{(1+\alpha)N}\int_{\phi_{N}(x)}^{\phi_{N}(x)+\sqrt{2\log{N}}}\log\left[\int_{-\infty}^{u}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right]du+E_{N}$
-where
+where, with respect to the error terms $O\left(\frac{1}{\sqrt{(1+\alpha)N-k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(x)}{2}}\right)$ and $O\left(\frac{1}{(1+\alpha)N}\right)$ of the Edgeworth expansion of the logarithm of the probability $p_{k}$,,
 $E_{N}=O\left(\left(\frac{\sqrt{\log N}}{\sqrt{(1+\alpha)N}}\right)\vee\left(\sqrt{\log{N}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)\right)$
@@
 This completes the remark.
-The term $\frac{1}{\sqrt{(1+\alpha)N}}$ appearing in the lower limit of integration is negligible and the function $\phi_{N}(x)=o\left(\sqrt{\log\left((1+\alpha)N\right)}\right)$.
+The term $\frac{1}{\sqrt{(1+\alpha)N}}$ appearing in the lower limit of integration is negligible.
+Extending the upper limit to $\infty$, the function $\phi_{N}(x)=o\left(\sqrt{\log\left((1+\alpha)N\right)}\right)$.
 This implies the following result
-$\log{\mathcal{P}_{N}}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)\approx\sqrt{(1+\alpha)N}\int_{\phi_{N}(x)}^{+\infty}\log\left[\int_{-\infty}^{t}\frac{e^{-\frac{s^{2}}{2}}}{2\pi}ds\right]dt+E_{N}$
-Let $F_{\infty}(x)$ denote the limiting distribution of the scaled spectral radius of any $N\times N$ complex induced Ginibre matrix as $N$ goes to infinity. The limit of its logarithm is
+$\log{\mathcal{P}_{N}}\left(1+\frac{f_{N}(x)}{\sqrt{(1+\alpha)N}}\right)=\sqrt{(1+\alpha)N}\int_{\phi_{N}(x)}^{+\infty}\log\left[\int_{-\infty}^{t}\frac{e^{-\frac{s^{2}}{2}}}{2\pi}ds\right]dt+E_{N}$
+Let $F_{\infty}(x)$ denote the limiting distribution of the scaled spectral radius of a $N\times N$ complex induced Ginibre matrix as $N$ goes to infinity. The limit of its logarithm is
 $\log{F_{\infty}}(x)=\lim_{N\rightarrow+\infty}\sqrt{(1+\alpha)N}\int_{\phi_{N}(x)}^{+\infty}\log\left[1-\int_{t}^{+\infty}e^{-\frac{s^{2}}{2}}\frac{ds}{\sqrt{2\pi}}\right]dt$
 where $\lim_{N\rightarrow+\infty}E_{N}=0$.
-The integral $\int_{t}^{+\infty}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds$ is less than $1$ and is close to zero such that, applying the Taylor expansion of the function $\log(1-x)\approx-x+O(x^{2})$, the logarithm of the limiting distribution $F_{\infty}(x)$ is
+As the function $f_{N}$ goes to infinity, the logarithm of the limiting distribution $F_{\infty}(x)$ is
 $\log{F_{\infty}}(x)$ $=-\lim_{N\rightarrow+\infty}\sqrt{\frac{(1+\alpha)N}{2\pi}}\int_{\phi_{N}(x)}^{+\infty}\left[\int_{t}^{+\infty}e^{-\frac{s^{2}}{2}}ds+O\left(\left(\int_{t}^{+\infty}e^{-\frac{s^{2}}{2}}ds\right)^{2}\right)\right]dt$
@@
 $\int_{\phi_{N}(x)}^{+\infty}\frac{1}{t}e^{-t^{2}/2}dt=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}-2\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt$
-and in the limit as $N$ goes to infinity
-$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt\approx\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times O\left(\frac{1}{\phi_{N}^{2}(x)}\right)$
+and in the limit as $N$ goes to infinity,
+$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}+O\left(\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{5}}dt\right)$
 This implies that
@@
 $=-\lim_{N\rightarrow+\infty}\sqrt{\frac{(1+\alpha)N}{2\pi}}\left[\frac{1}{4f_{N}^{2}(x)}e^{-2f_{N}^{2}(x)}\left(1+O\left(\frac{1}{4f_{N}^{2}(x)}\right)\right)\right]$
-The function $f_{N}(x)$ is chosen with the use of the definition of a class of limiting distributions exposed in the reference [12].
-For a convenient choice of the function $f_{N}(x)$, the limiting distribution $F_{\infty}(x)$ is from the class of Extreme Value Distributions composed of the three types of extreme value distributions for maxima.
+Then, a convenient choice of the function $f_{N}(x)$ is performed, such that the limiting distribution $F_{\infty}(x)$ is from the class of Extreme Value Distributions composed of the three types of extreme value distributions for maxima.
+More precisely, choosing
 $f_{N}^{2}(x)=\frac{1}{2}\log\left(\frac{e^{x}\sqrt{(1+\alpha)N/2\pi}}{\log N}\right)$
+The limit, uniformly on compact sets in $x$, as in [22],
 $\lim_{N\rightarrow+\infty}\log\mathcal{P}_{N}\left[R_{N}\leq 1+\sqrt{\frac{1}{2(1+\alpha)N}}\left(\log\frac{\sqrt{(1+\alpha)N/2\pi}}{\log{N}}+x\right)^{1/2}\right]$
@@
 $=-\lim_{N\rightarrow+\infty}\exp(-x)\frac{\log{N}}{2\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}$
-And, for any $\alpha>0$ and fixed $|x|\ll N$
-$\frac{\log{N}}{2\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}$
-$=\frac{1}{2}\frac{\log{N}}{\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}$
-$=\frac{1}{2}\frac{\log{N}}{\left[x+\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N)\right]}$
-$\sim\frac{1}{2}\frac{\log{N}}{\left[\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N\right]}$
-$\sim\frac{\log{N}}{\log((1+\alpha)N/(2\pi))}$
-$=\frac{1}{1+\frac{\log\left(\frac{1+\alpha}{2\pi}\right)}{\log N}}$
+Also,
+$\frac{\log{N}}{2\log(e^{x}\sqrt{(1+\alpha)N/(2\pi)}\times\frac{1}{\log{N}})}\\ =\frac{1}{2}\frac{\log{N}}{\left[x+\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N)\right]}$
+And, for any $\alpha>0$ and fixed $|x|\ll N$,
+$\lim_{N\rightarrow+\infty}\frac{1}{2}\frac{\log{N}}{\left[x+\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N)\right]}$
+$=\lim_{N\rightarrow+\infty}\frac{1}{2}\frac{\log{N}}{\left[\frac{1}{2}\log((1+\alpha)N/(2\pi))-\log\log N\right]}$
+$=\lim_{N\rightarrow+\infty}\frac{\log{N}}{\log((1+\alpha)N/(2\pi))}$
+$=\lim_{N\rightarrow+\infty}\frac{1}{1+\frac{\log\left(\frac{1+\alpha}{2\pi}\right)}{\log N}}$
 This implies
 $\lim_{N\rightarrow+\infty}\frac{\log{N}}{2\log(e^{x}\sqrt{(1+\alpha)N/2\pi}\times\frac{1}{\log{N}})}=1$
-Finally, let $R_{N}$ denote the scaled spectral radius.
+Finally,
 $\lim_{N\rightarrow+\infty}\log\mathcal{P}_{N}\left[R_{N}\leq 1+\sqrt{\frac{1}{2(1+\alpha)N}}\left(\log\frac{\sqrt{(1+\alpha)N/2\pi}}{\log{N}}+x\right)^{1/2}\right]=-\exp(-x)$
@@
 $\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[R_{N}\leq 1+\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}+\frac{x}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}\right]=\exp(-\exp(-x))$
-This limiting distribution is the standard Gumbel distribution of maxima. The corresponding cumulative density function is the function $F_{Gumbel}^{(max)}(x)=\exp(-\exp(-x))$.
-The scaled spectral radius $R_{N}$ is a random variable approximated as
+This limiting distribution is the standard Gumbel distribution of maxima.
+The scaled spectral radius $R_{N}$ is a random variable approximated by
 $R_{N}$ $\simeq 1+\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}+\frac{X}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$
@@
 where the term $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2(1+\alpha)N}}$ with $\gamma_{\alpha N}=\log{\sqrt{(1+\alpha)N/2\pi}}-\log\log{N}$.
-The random variable $Z$ denote a random variable following a standard exponential distribution which implies that $X=-\log(Z)$ is a standard Gumbel(maximum)-distributed random variable.
+The random variable $Z$ denote a random variable following a standard Exponential distribution.
+The random variable $X=-\log(Z)$ is a standard Gumbel(maximum)-distributed random variable.
 The random variable $\xi_{\alpha,N}=-\frac{\log(Z)}{2\sqrt{2(1+\alpha)N\gamma_{\alpha,N}}}$ is Gumbel-distributed.
@@
 A.2. Proof of Theorem 7
-It is demonstrated in Section 2 that the scaled minimum radius $\frac{r_{min}^{(N)}(A)}{\sqrt{N}}$ of eigenvalues for matrices from the complex Ginibre ensemble, close to zero, follows a Rayleigh distribution.
-The complex Ginibre ensemble does correspond to the complex induced Ginibre ensemble when the rectangularity index $L$ is equal to zero.
-Similarly stated in Section 3, Theorem 7 is as follows.
-Theorem (Limiting distribution of the scaled minimum modulus $r_{N}$ for the complex induced Ginibre ensemble).
-Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with rectangularity index $L$ proportional to $N$ (i.e., $L=\alpha N$, $\alpha>0$) and let $r_{N}$ denote the scaled minimum modulus such that $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$.
-Then,
-$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]=e^{-e^{x}}$ (43)
-where $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$.
-This limit corresponds to the survival distribution function of the standard Gumbel distribution for minima. The scaled minimum modulus $r_{N}$ is approximated with the following
-$r_{N}\simeq 1-T_{\alpha,N}+\xi_{\alpha,N}$ (44)
-where $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$.
-The random variable $\xi_{\alpha,N}=\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ where $X$ follows a standard Gumbel (minimum) distribution.
-Proof.
 The limiting distribution of the scaled minimum radius at the inner edge of the ring, in the limit as $N$ goes to infinity and with the rectangularity index $L$ proportional to $N$, i.e., $L=\alpha N$, $\alpha>0$, is studied in this section.
@@
 The function $f_{Gamma(k+L,1)}(t)$ is the probability density function of the Gamma distribution with shape parameter $k+L$ and rate parameter $1$.
-which implies,
+This implies,
 $P\left(\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}\geq a\right)=\prod_{k=1}^{N}P\left(\sum_{j=1}^{k+L}Z_{j}\geq\alpha Na^{2}\right)=\prod_{k=1}^{N}P\left(\frac{1}{\alpha N}\sum_{j=1}^{k+L}Z_{j}\geq a^{2}\right)$
@@
 $\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
-$\approx\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
-$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-\frac{1}{2}\left(1+\operatorname{erf}\left(-\frac{\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)}{\sqrt{2}}\right)\right)\right]$
+$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\geq-\frac{\sqrt{\alpha N}}{\sqrt{\alpha N+k}}\phi_{N}(x)-\frac{k}{\sqrt{\alpha N+k}}\right)$
+$>\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\geq-\frac{k}{\sqrt{\alpha N+k}}\right)$
+$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-\frac{1}{2}\left(1+\operatorname{erf}\left(-\frac{\left(\frac{k}{\sqrt{\alpha N+k}}\right)}{\sqrt{2}}\right)\right)\right]$
 $>\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\frac{k^{2}}{\alpha N}}\right)>\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\alpha N\delta_{N}^{2}}\right)>\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$
@@
 $=\prod_{k=1}^{\alpha N\delta_{N}}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\left(\sqrt{\frac{\alpha N}{\alpha N+k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)\right)$
-$=\prod_{k=1}^{\alpha N\delta_{N}}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
-$=\prod_{k=1}^{\alpha N\delta_{N}}P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
 $=\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$
 where
-$p_{k}=P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
-As in the reference [20], it is assumed that $k=o(\alpha N)$.
-The factor $\sqrt{\frac{\alpha N}{\alpha N+k}}$ is assumed equal (approximatively) to one for small and moderate values of $k$.
-For large values of $k$, this factor is rather large and the probability
-$P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\left(\sqrt{\frac{\alpha N}{\alpha N+k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)\right)$ is approximatively equal to one.
-This means that its logarithm is zero.
+$p_{k}=P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\left(\sqrt{\frac{\alpha N}{\alpha N+k}}\right)\left(\phi_{N}(x)+\frac{k}{\sqrt{\alpha N}}\right)\right)$
+and, for fixed $k\in\left[0,\sqrt{2\alpha N\log(N)}\right]$, the factor $\sqrt{\frac{\alpha N}{\alpha N+k}}$ goes to one as $N$ goes to infinity.
 Finally,
@@
 where each element of sequence $Z_{1},\cdots,Z_{\alpha N+k}$ is i.i.d. exponentially distributed with parameter equal to $1$.
-The variable $\bar{Z}=\frac{1}{\alpha N+k}\sum_{j=1}^{\alpha N+k}Z_{j}$ is the empirical mean. The mean of $\bar{Z}$ is $\mu=1$ and its standard deviation is
-$\frac{\sigma}{\sqrt{\alpha N+k}}=\frac{1}{\sqrt{\alpha N+k}}$ with $\sigma=1$.
-Applying the Edgeworth expansion, the logarithm of the probability $p_{k}$ is then
+The variable $\bar{Z}=\frac{1}{\alpha N+k}\sum_{j=1}^{\alpha N+k}Z_{j}$ is the empirical mean.
+The mean of $\bar{Z}$ is equal to $1$ and its standard deviation is $\frac{\sigma}{\sqrt{\alpha N+k}}=\frac{1}{\sqrt{\alpha N+k}}$ with $\sigma=1$.
+Applying the Edgeworth expansion, as in [22], the logarithm of the probability $p_{k}$ is then
 $\log\left[P\left(\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\geq-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}\right)\right]$
-$=\log\left[\int_{-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}}^{+K_{N}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N+k}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)+O\left(\frac{1}{\alpha N+k}\right)$
-$+O\left(K_{N}(\alpha N+k)^{-3/2}\right)+O\left(e^{-\frac{1}{2}K_{N}^{2}}\right)$
-where, similarly stated in [20], for $x$ restricted as in $|x|<Y$ for some large positive $Y$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq Y}\phi_{N}(x)$.
+$=\log\left[\int_{-\phi_{N}(x)-\frac{k}{\sqrt{\alpha N}}}^{+K_{N}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N}}\sup_{|c|\leq H}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)+O\left(\frac{1}{\alpha N}\right)$
+$+O\left(K_{N}(\alpha N)^{-3/2}\right)+O\left(e^{-\frac{1}{2}K_{N}^{2}}\right)$
+where, similarly stated in [22], for $x$ restricted as in $|x|<H$ for some large positive $H$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq Y}\phi_{N}(x)$.
 Defining the function $f(k)$ as
@@
 $\sum_{k=1}^{\sqrt{2\alpha N\log N}}f(k)=\sqrt{\alpha N}\int_{\phi_{N}(x)+\frac{1}{\sqrt{\alpha N}}}^{\phi_{N}(x)+\sqrt{2\log N}}\log\left[\int_{-u}^{+\infty}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right]du+E_{N}$
-where $E_{N}$ is the error term
+where, with respect to the error terms $O\left(\frac{1}{\sqrt{\alpha N}}\sup_{|c|\leq H}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)$ and $O\left(\frac{1}{\alpha N}\right)$ of the Edgeworth expansion of the logarithm of the probability $p_{k}$,
 $E_{N}=O\left(\left(\frac{\sqrt{\log N}}{\sqrt{\alpha N}}\right)\vee\left(\sqrt{\log{N}}\sup_{|c|\leq Y}\phi_{N}^{2}(c)e^{-\frac{\phi_{N}^{2}(c)}{2}}\right)\right)$
-The term $\frac{1}{\sqrt{\alpha N}}$ appearing in the lower limit of integration is negligible and the function $\phi_{N}(x)=o\left(\sqrt{\log\left(\alpha N\right)}\right)$.
+The term $\frac{1}{\sqrt{\alpha N}}$ appearing in the lower limit of integration is negligible.
+Extending the upper limit to $\infty$, the function $\phi_{N}(x)=o\left(\sqrt{\log\left(\alpha N\right)}\right)$.
 This implies the following result
@@
 and $\lim_{N\rightarrow+\infty}E_{N}=0$.
-The integral $\int_{-\infty}^{-t}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds$ is less than $1$ and is close to zero such that, applying the Taylor expansion of the function $\log(1-x)\approx-x+O(x^{2})$.
-Thus, the logarithm of the limiting survival distribution function $1-F_{\infty}(x)$ is
+Thus, as the function $f_{N}$ goes to infinity, the logarithm of the limiting survival distribution function $1-F_{\infty}(x)$ is
 $\log\left(1-F_{\infty}(x)\right)$ $=-\lim_{N\rightarrow+\infty}\sqrt{\frac{\alpha N}{2\pi}}\int_{\phi_{N}(x)}^{+\infty}\left[\int_{-\infty}^{-t}e^{-\frac{s^{2}}{2}}ds+O\left(\left(\int_{-\infty}^{-t}e^{-\frac{s^{2}}{2}}ds\right)^{2}\right)\right]dt$
@@
 $\int_{\phi_{N}(x)}^{+\infty}\frac{1}{t}e^{-t^{2}/2}dt=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}-\int_{\phi_{N}(x)}^{+\infty}\frac{2e^{-t^{2}/2}}{t^{3}}dt$
-and in the limit as $N$ goes to infinity
-$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt\approx\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times O\left(\frac{1}{\phi_{N}^{2}(x)}\right)$
+and in the limit as $N$ goes to infinity,
+$\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{3}}dt=\frac{1}{\phi_{N}^{2}(x)}e^{-\frac{\phi_{N}^{2}(x)}{2}}\times\frac{1}{\phi_{N}^{2}(x)}+O\left(\int_{\phi_{N}(x)}^{+\infty}\frac{e^{-t^{2}/2}}{t^{5}}dt\right)$
 This implies that
@@
 For a convenient choice of the function $f_{N}(x)$, the limiting distribution $F_{\infty}(x)$ is from the class of Extreme Value Distributions composed of the three types of extreme value distributions for minima [13].
+More precisely, choosing
 $f_{N}^{2}(x)=\frac{1}{2}\log\left(\frac{e^{-x}\sqrt{\alpha N/2\pi}}{\log N}\right)$
@@
 $\left(\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}-x\right)^{1/2}=\gamma_{\alpha,N}^{1/2}\left(1-\frac{x}{\gamma_{\alpha,N}}\right)^{1/2}=\gamma_{\alpha,N}^{1/2}-\frac{x}{2\gamma_{\alpha,N}^{1/2}}+O\left(\left(\frac{x}{\gamma_{\alpha,N}}\right)^{2}\right)$
-which implies that
-$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\left(\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}-\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right)\right]$
-$=\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]$
-$=\exp(-\exp(x))$
+This implies that
+$\lim_{N\rightarrow+\infty}\mathcal{P}_{N}\left[r_{N}\geq 1-\left(\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}-\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right)\right]=\exp(-\exp(x))$
 The limiting distribution function of the scaled minimum modulus of eigenvalues for matrices from the complex induced Ginibre ensemble is
 $F_{\infty}(x)=1-\exp\left(-\exp\left(x\right)\right)$
-and corresponds to distribution function of the standard Gumbel distribution for minima.
+and corresponds to the distribution function of the standard Gumbel distribution for minima.
 The scaled minimum modulus denoted $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$ is a random variable well approximated as
@@
 where the term $T_{\alpha,N}=\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}$ with $\gamma_{\alpha N}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$.
-The random variable $Z$ denotes a random variable following a standard exponential distribution which implies that $X=-\log(Z)$ is a standard Gumbel(minimum)-distributed random variable.
+The random variable $Z$ denotes a random variable following a standard Exponential distribution which implies that $X=-\log(Z)$ is a standard Gumbel-distributed random variable.
 The random variable $\xi_{\alpha,N}=-\frac{\log(Z)}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$ is Gumbel-distributed.
@@
 A.3. Proof of Theorem 10
-The independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble is established from Theorem 10 presented here again for more convenience.
-Theorem (Independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble).
-Let $G$ denote a $N\times N$ matrix from the complex induced Ginibre ensemble with proportional rectangularity index $L=\alpha N$, $\alpha>0$.
-The scaled spectral radius $R_{N}$ and the scaled minimum modulus $r_{N}$ are independent random variables, under scaling $\sqrt{\alpha N}$, as $N$ goes to infinity.
-More precisely, let $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$.
-Setting $\rho=\sqrt{\frac{1+\alpha}{\alpha}}$,
-$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\xi_{\alpha}^{(N)}(y)\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\eta_{\alpha}^{(N)}(x)\right]=e^{-e^{y}}e^{-e^{-x}}$
-with $\gamma_{\alpha,N}=\log\frac{\sqrt{\alpha N/2\pi}}{\log{N}}=\log{\sqrt{\alpha N/2\pi}}-\log\log{N}$ and where
-$\xi_{\alpha}^{(N)}(y)=\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }\eta_{\alpha}^{(N)}(x)=\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
-The scaled spectral radius $R_{N}$ and the scaled minimum of moduli $r_{N}$ are approximated as
-$R_{N}$ $\simeq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{X}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{ }\text{ and }\text{ }r_{N}\simeq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{Y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}$
-where $X$ and $Y$ are standard Gumbel (maximum) and standard Gumbel (minimum) distributed random variables, respectively.
-Proof.
+The independence of the scaled spectral radius and the scaled minimum modulus for the complex induced Ginibre ensemble is established as follows.
 The independence of the scaled spectral radius and the scaled minimum modulus is derived from their joint cumulative distribution function
@@
 The rectangularity index $L$ is assumed proportional to $N$, i.e., $L=\alpha N$, $\alpha>0$.
-Considering the scaling $\sqrt{\alpha N}$, the inner radius $r_{in}=1$ and the outer radius $r_{out}=\sqrt{\frac{1+\alpha}{\alpha}}=\gamma$.
-Let $r=1-\frac{f_{N}(y)}{\sqrt{\alpha N}}$ and $R=\sqrt{\frac{1+\alpha}{\alpha}}+\frac{f_{N}(x)}{\sqrt{\alpha N}}=\rho+\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ and $f_{N}(y)$ are $o\left(\alpha N\right)$.
+Considering the scaling $\sqrt{\alpha N}$, the inner radius $r_{in}=1$ and the outer radius $r_{out}=\sqrt{\frac{1+\alpha}{\alpha}}=\rho$.
+Let $r=1-\frac{f_{N}(y)}{\sqrt{\alpha N}}$ and $R=\sqrt{\frac{1+\alpha}{\alpha}}+\frac{f_{N}(x)}{\sqrt{\alpha N}}=\rho+\frac{f_{N}(x)}{\sqrt{\alpha N}}$ where $f_{N}(x)$ and $f_{N}(y)$ are $o\left(\sqrt{\alpha N}\right)$.
 This implies
@@
 $=\prod_{k=1}^{N}P\left(r^{2}\leq\frac{1}{\alpha N}\sum_{j=1}^{k+L}Z_{j}\leq R^{2}\right)$
-$=\prod_{k=1}^{N}\left[P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\right]$
-where $C=\rho^{2}-1>0$. The functions $\phi_{N}^{(r)}(y)=2f_{N}(y)\left(1-\frac{f_{N}(y)}{2\sqrt{\alpha N}}\right)$ and
-$\phi_{N}^{(R)}(x)=2\rho f_{N}(x)\left(1+\frac{f_{N}(x)}{2\rho\sqrt{\alpha N}}\right)$.
-Furthermore, with proportional rectangularity index $L=\alpha N$, $\alpha>0$,
-$\prod_{k=1}^{N}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
-$=\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\times$
-$\prod_{k=1}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$
-$=\prod_{k=1}^{N}p_{k}\prod_{k=1}^{N}h_{k}$
+$=\prod_{k=1}^{N}\left[P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)\right]$
+where $C_{\alpha,N}=\sqrt{\alpha N}\left(\rho^{2}-1\right)>0$. The functions $\phi_{N}^{(r)}(y)=2f_{N}(y)\left(1-\frac{f_{N}(y)}{2\sqrt{\alpha N}}\right)$ and
+$\phi_{N}^{(R)}(x)=2\rho f_{N}(x)\left(1-\frac{f_{N}(x)}{2\rho\sqrt{\alpha N}}\right)$.
+Furthermore, with proportional rectangularity index $L=\alpha N$, $\alpha>0$, the partial product
+$\prod_{k=1}^{N}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$
+is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf. [22]), such as
+$\prod_{k=1}^{\alpha N\delta_{N}}p_{k}\prod_{k=\alpha N\delta_{N}}^{N}p_{k}\leq P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)\leq\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$
 where
-$p_{k}=P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)<\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
-and
-$h_{k}=\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$
-The partial product $\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$ is bounded by any partial product of $p_{k}$, i.e., $\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$, for whatever positive $\delta_{N}$ less than 1 (cf.
-[20]), such as
-$\prod_{k=1}^{\alpha N\delta_{N}}p_{k}\prod_{k=\alpha N\delta_{N}}^{N}p_{k}\leq\prod_{k=1}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\leq\prod_{k=1}^{\alpha N\delta_{N}}p_{k}$
-The following is derived using the same arguments as stated in [20] and using the Markov inequality and the definition of the quantile of the exponential distribution.
-The function $\phi_{N}^{(R)}(x)$ is a positive and increasing function.
-$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
-$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\sqrt{\alpha N}\left(\phi_{N}^{(R)}(x)+C\right)-k\right)$
-The random variables $Z_{j}$, $i\in\{1,\cdots,\alpha N+k\}$ are independent and identically distributed.
-Applying the Markov inequality, with $0<\eta<1$ , as the exponential is an strictly increasing and convex function, this implies that
-$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{(\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
-$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\sqrt{\alpha N}\left(\phi_{N}^{(R)}(x)+C\right)-k\right)$
-$=\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}Z_{j}\leq\sqrt{\alpha N}\left(\phi_{N}^{(R)}(x)+C\right)+\alpha N\right)$
-$\geq\prod_{k=\alpha N\delta_{N}}^{N}P\left(\sum_{j=1}^{\alpha N+k}Z_{j}\leq\sqrt{\alpha N}\phi_{N}^{(R)}(x)\right)$
-$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{\alpha N+k}Z_{j}>\sqrt{\alpha N}\phi_{N}^{(R)}(x)\right)\right]$
-$\geq\prod_{k=\alpha N\delta_{N}}^{N}\left[1-e^{-\eta\sqrt{\alpha N}\phi_{N}(x)}E\left[e^{\eta Z_{1}}\right]^{\alpha N+k}\right]$
+$p_{k}=P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$
 Furthermore,
-$e^{-\eta\alpha N\left(\frac{k}{\alpha N}-1\right)-\alpha N\ln(1-\eta)}$ $\geq e^{(\alpha N+k)(\eta-\ln(1-\eta))}$
-$=e^{-\eta\alpha N(\frac{k}{\alpha N}-1)-\left(\alpha N+k\right)\ln(1-\eta)}>1$
-And with $0<\eta<1$ and $\forall Y\in\mathbb{R}$
-$e^{Y-\eta\sqrt{\alpha N}\phi_{N}^{(R)}(x)}<e^{Y}\Rightarrow 1-e^{Y-\eta\sqrt{\alpha N}\phi_{N}^{(R)}(x)}>1-e^{Y}$
-Now, setting $\eta=1-\frac{1}{\frac{k}{\alpha N}-1}$ which does maximise the remaining exponent (similarly presented in [20]) and applying the the quantile formula of the exponential distribution, this implies
-$\prod_{k=\alpha N\delta_{N}}^{N}P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}(Z_{j}-1)\leq\phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}\right)$
-$\geq\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\alpha N\left[\eta\left(\frac{k}{\alpha N}-1\right)+\ln(1-\eta)\right]}\right)$
-$\geq\prod_{k=\alpha N\delta_{N}}^{N}\left(1-e^{-\alpha N\left[\frac{k}{\alpha N}-\ln\left(\frac{k}{\alpha N}-1\right)\right]}\right)$
-$\geq\left(1-e^{-\alpha N\delta_{N}^{2}}\right)^{N}$
-Finally,
-$\prod_{k=\alpha N\delta_{N}}^{N}p_{k}\geq\left(1-e^{-\alpha N\delta_{N}^{2})}\right)^{N}$
-The parameter $\delta_{N}$ is chosen such that $\left(1-e^{-\alpha N\delta_{N}^{2})}\right)^{N}=\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$, i.e., $\delta_{N}=\sqrt{\frac{2\log{N}}{\alpha N}}$.
-The partial product $\prod_{k=1}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$ is bounded.
-It is established from the following that the limit of the partial product $\prod_{k=1}^{\alpha N\delta_{N}}h_{k}$ is equal to one as $N$ goes infinity.
-More precisely,
-$\prod_{k=1}^{\alpha N\delta_{N}}h_{k}\prod_{k=\alpha N\delta_{N}}^{N}h_{k}\leq\prod_{k=1}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]\leq\prod_{k=1}^{\alpha N\delta_{N}}h_{k}$
-Lower bound of the partial product $\prod_{k=\alpha N\delta_{N}}^{N}h_{k}$.
-For large $N$, as the random variable $\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)$ follows a standard Gaussian distribution, and taking into account a similar remark as in the paper of B.
-Rider [20] saying that the upper limit of integration can be extended to $\infty$, the probability
-$P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)\approx 1$
-Now, considering the probability
-$P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)$
-$=1-P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)>-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)$
-$=1-P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)<\phi_{N}^{(r)}(y)+\frac{k}{\sqrt{\alpha N}}\right)<\frac{1}{N^{2}}$
-where the random variable $\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)$ follows a standard Gaussian distribution.
-This implies that
-$P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)>-\frac{1}{N^{2}}$
-Consequently,
-$1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}>1-\frac{1}{N^{2}}$
-Thus,
-$\prod_{k=\alpha N\delta_{N}}^{N}\left[1-\frac{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)}{P\left(\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{\alpha N+k}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)}\right]$
-$>\prod_{k=\alpha N\delta_{N}}^{N}\left(1-\frac{1}{N^{2}}\right)>\left(1-\frac{1}{N^{2}}\right)^{N}=1-O\left(\frac{1}{N}\right)$
-The lower bound is $1-O\left(\frac{1}{N}\right)$ and goes to $1$ as $N$ goes to infinity.
+$\prod_{k=\alpha N\delta_{N}}^{N}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$
+$=\prod_{k=\alpha N\delta_{N}}^{N}\left[P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)-P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)\right]$
+$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\geq\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}\right)-P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)\right]$
+$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\geq\sqrt{\alpha N}\Phi_{N}^{(R)}(x)-k\right)-P\left(\frac{\sum_{j=1}^{k+L}\left(Z_{j}-1\right)}{\sqrt{\alpha N}}\leq-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)\right]$
+$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{k+L}Z_{j}\geq\sqrt{\alpha N}\Phi_{N}^{(R)}(x)+\alpha N\right)-P\left(\frac{\sum_{j=1}^{k+L}Z_{j}}{\sqrt{\alpha N+k}}\leq-\left(\sqrt{\frac{\alpha N}{\alpha N+k}}\right)\phi_{N}^{(r)}(y)+\frac{k}{\alpha N+k}\right)\right]$
+$>\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{k+L}Z_{j}\geq\sqrt{\alpha N}\Phi_{N}^{(R)}(x)+\alpha N\right)-P\left(\frac{\sum_{j=1}^{k+L}Z_{j}}{\sqrt{\alpha N+k}}\leq\frac{k}{\alpha N+k}\right)\right]$
+$=\prod_{k=\alpha N\delta_{N}}^{N}\left[1-P\left(\sum_{j=1}^{k+L}Z_{j}\geq\sqrt{\alpha N}\Phi_{N}^{(R)}(x)+\alpha N\right)-P\left(\frac{\sum_{j=1}^{k+L}Z_{j}}{\sqrt{\alpha N+k}}\leq\frac{k}{\alpha N+k}\right)\right]$
+$>\prod_{k=\alpha N\delta_{N}}^{N}\left[1-O\left(\frac{1}{N}\right)\right]$
 Now,
-$\log\prod_{k=1}^{\alpha N\delta_{N}}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
-$=\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
-Furthermore, using the classical Edgeworth expansion
-$\log P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}\right)$
-$=\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N-k}}\sup_{|c|\leq Y}(\phi_{N}^{(R)}(c))^{2}e^{-\frac{(\phi_{N}^{(R)}(c))^{2}}{2}}\right)$
-$+O\left(\frac{1}{\alpha N-k}\right)+O\left(\phi_{N}^{(r)}(y)(\alpha N-k)^{-3/2}\right)+O\left(e^{-\frac{(\phi_{N}^{(r)}(y))^{2}}{2}}\right)$
-where, as in [20], for $x$ restricted as in $|x|\leq Y$ for some large positive $Y$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq Y}\phi_{N}(x)$.
-The leading order sum as the Riemann integral
-$\sum_{k=1}^{\sqrt{2N\log N}}\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$ $\simeq\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\log\left[\int_{-\infty}^{u}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du$
-$+\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{+\infty}\log\left[\int_{-u}^{+\infty}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du+E_{N}$
+$\log\prod_{k=1}^{\alpha N\delta_{N}}P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$
+$=\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log P\left(\alpha_{k}^{(N)}\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\right)\leq\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{k+\alpha N}\left(Z_{j}-1\right)\leq\alpha_{k}^{(N)}\left(\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)\right)$
+where, for fixed $k\in\left[0,\alpha N\delta_{N}\right]$, the factor $\alpha_{k}^{(N)}=\sqrt{\frac{\alpha N}{\alpha N+k}}$ goes to one as $N$ goes to infinity.
+Furthermore, using the classical Edgeworth expansion, in line with equation $(9)$ in [22], this implies that
+$\log P\left(-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}\leq\frac{1}{\sqrt{\alpha N+k}}\sum_{j=1}^{k+L}\left(Z_{j}-1\right)\leq\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}\right)$
+$=\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]+O\left(\frac{1}{\sqrt{\alpha N}}\sup_{|c|\leq H}(\phi_{N}^{(R)}(c))^{2}e^{-\frac{(\phi_{N}^{(R)}(c))^{2}}{2}}\right)$
+$+O\left(\frac{1}{\alpha N}\right)+O\left(\phi_{N}^{(r)}(y)(\alpha N)^{-3/2}\right)+O\left(e^{-\frac{(\phi_{N}^{(r)}(y))^{2}}{2}}\right)$
+where, as in [22], for $x$ restricted as in $|x|\leq H$ for some large positive $H$ and any $K_{N}$ goes to $\infty$ faster than $\sup_{|x|\leq H}\phi_{N}^{(R)}(x)$.
+Then,
+$\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log\left[\int_{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}^{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
+$=\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log\left[1-\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt-\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
+In the limit as $N$ goes to infinity,
+$\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log\left[1-\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt-\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
+$=\sum_{k=1}^{\sqrt{2\alpha N\log N}}\left[-\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt-\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
+$+\sum_{k=1}^{\sqrt{2\alpha N\log N}}O\left(\left(\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}+\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt+\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)^{2}\right)$
+$=-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{s^{2}}{2}}}{\sqrt{2\pi}}ds\right)\right)$
+$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
+$=-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)+O\left(\frac{e^{-\frac{\left(\phi_{N}^{(r)}(y)\right)^{2}}{2}}}{\phi_{N}^{(r)}(y)}\right)\right)$
+$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
+This implies that,
+$\sum_{k=1}^{\sqrt{2\alpha N\log N}}\log\left[1-\int_{\phi_{N}^{(R)}(x)+C_{\alpha,N}-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt-\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right]$
+$=-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
+$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
+where $\Phi_{N}^{(R)}(x)=\phi_{N}^{(R)}(x)+C_{\alpha,N}$.
+The leading order sum as the Riemann integral of last equation is,
+$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{\Phi_{N}^{(R)}(x)-\frac{k}{\sqrt{\alpha N}}}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
+$-\sum_{k=1}^{\sqrt{2\alpha N\log N}}\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-\phi_{N}^{(r)}(y)-\frac{k}{\sqrt{\alpha N}}}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)$
+$=-\sqrt{\alpha N}\int_{\Phi_{N}^{(R)}(x)}^{\Phi_{N}^{(R)}(x)-\sqrt{2\log(N)}}\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)\right)du$
+$-\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{\phi_{N}^{(r)}(y)+\sqrt{2\log(N)}}\left(\int_{-\infty}^{-u}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-u}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)\right)du+E_{N}$
 where
-$E_{N}=O\left(\left(\sqrt{\frac{\log N}{\alpha N}}\right)\vee\left(\sqrt{\log{N}}\sup_{|c|\leq Y}(\phi_{N}^{(R)}(c))^{2}e^{-\frac{(\phi_{N}^{(R)}(c))^{2}}{2}}\right)\right)$
-Following the framework presented in [20] and detailed in Appendices A.1 and A.2,
-$\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\log\left[\int_{-\infty}^{u}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du=\log F_{X}^{(\infty)}(x)=-\exp(-x)$
-and
-$\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{+\infty}\log\left[\int_{-u}^{+\infty}\frac{e^{-s^{2}/2}}{\sqrt{2\pi}}ds\right]du=\log\left(1-F_{Y}^{(\infty)}(y)\right)=-\exp(y)$
+$E_{N}=O\left(\left(\sqrt{\frac{\log N}{\alpha N}}\right)\vee\left(\sqrt{\log{N}}\sup_{|c|\leq H}(\Phi_{N}^{(R)}(c))^{2}e^{-\frac{(\Phi_{N}^{(R)}(c))^{2}}{2}}\right)\right)$
+Remark.
+For $\alpha=N^{m},(\rho^{2}-1)\sqrt{\alpha N}=\frac{1}{\alpha}\sqrt{\alpha N}=\frac{1}{N^{m-\frac{m}{2}-\frac{1}{2}}}=\frac{1}{N^{\frac{m-1}{2}}}$.
+In the limit as $N$ goes to infinity, with $\alpha>0$, under the condition that
+    $\exists m_{1}>1$, $\forall m>m_{1}$ and $\alpha=O\left(N^{m}\right)$, as $N$ goes to infinity, $\phi_{N}^{(R)}(x)=o\left(\sqrt{\alpha N}\right)>\sqrt{2\log(N)}$.
+This assumption is used such that the term $C_{\alpha,N}$, goes to zero as $N$ goes to infinity.
+Following the framework presented in [22],
+$\log F_{X}^{(\infty)}(x)$ $=-\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)\right)du$
+$=-\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(R)}(x)}^{+\infty}\left(\int_{u}^{+\infty}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)du\times\left(1+O\left(\frac{1}{\left(\phi_{N}^{(R)}(x)\right)^{2}}\right)\right)$
+$=-\lim_{N\rightarrow+\infty}\sqrt{\frac{\alpha N}{2\pi}}\frac{e^{-\frac{\left(\phi_{N}^{(R)}(x)\right)^{2}}{2}}}{\left(\phi_{N}^{(R)}(x)\right)^{2}}\times\left(1+O\left(\frac{1}{\left(\phi_{N}^{(R)}(x)\right)^{2}}\right)\right)$
+$=-\exp(-x)$
+with, $x\in\mathbb{R}$,
+$f_{N}^{2}(x)=\frac{1}{2\rho^{2}}\left(\log\left(\frac{e^{x}\sqrt{N/(2\pi)}}{\log(N)}\right)\right)$
+And,
+$\log\left(1-F_{Y}^{(\infty)}(y)\right)$ $=-\lim_{N\rightarrow+\infty}\sqrt{\alpha N}\int_{\phi_{N}^{(r)}(y)}^{+\infty}\left(\int_{-\infty}^{-u}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\times\left(1+O\left(\int_{-\infty}^{-u}\frac{e^{-\frac{t^{2}}{2}}}{\sqrt{2\pi}}dt\right)\right)\right)du$
+$=-\exp(y)$
+with, $y\in\mathbb{R}$,
+$f_{N}^{2}(y)=\frac{1}{2}\left(\log\left(\frac{e^{-y}\sqrt{N/(2\pi)}}{\log(N)}\right)\right)$
 which implies
-$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{x}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\right]$
+$\lim_{N\rightarrow+\infty}P\left[r_{N}\geq 1-\sqrt{\frac{\gamma_{\alpha,N}}{2\alpha N}}+\frac{y}{2\sqrt{2\alpha N\gamma_{\alpha,N}}}\text{}\text{ and }\text{}R_{N}\leq\rho+\sqrt{\frac{\gamma_{\alpha,N}}{2\rho^{2}\alpha N}}+\frac{x}{2\sqrt{2\rho^{2}\alpha N\gamma_{\alpha,N}}}\right]$
 $=F_{X}^{(\infty)}(x)\left(1-F_{Y}^{(\infty)}(y)\right)$
@@
 These limiting probability distributions are the standard Gumbel distribution for maxima and the standard Gumbel distribution for minima with cumulative density function $F_{X}(x)=\exp(-\exp(-x))$ and $F_{Y}(y)=1-\exp(-\exp(y))$, respectively.
-This proves the independence of the scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$ for $N\times N$ matrices from the complex induced Ginibre ensemble with proportional rectangularity index as $N$ goes to infinity.
+This proves the independence of the scaled spectral radius $R_{N}=\frac{r_{max}^{(N)}(G)}{\sqrt{\alpha N}}$ and the scaled minimum modulus $r_{N}=\frac{r_{min}^{(N)}(G)}{\sqrt{\alpha N}}$ of a $N\times N$ matrices from the complex induced Ginibre ensemble with proportional rectangularity index as $N$ goes to infinity.
 Appendix B Conditional probability for the complex Ginibre ensemble
-Let $\mathcal{D}(z_{0},s)$ denote the disk of radius $s$ centred at $z_{0}$ and $H^{(N)}(s,z_{0})$ the conditional probability that given one eigenvalue lies at the point $z_{0}$, the others are found outside $\mathcal{D}(z_{0},s)$, i.e.,
+Let $\mathcal{D}(z_{0},s)$ denote the disk of radius $s$ centred at $z_{0}$.
+The conditional probability $H^{(N)}(s,z_{0})$, that given one eigenvalue lies at the point $z_{0}$ and the others are found outside $\mathcal{D}(z_{0},s)$, is a result initially presented in [14] and then referenced in [17].
+This result is derived as follows.
+Let $z_{1},...,z_{N}$ denote the $N$ eigenvalues of an $N\times N$ complex Ginibre matrix.
+The conditional probability $H^{(N)}(s,z_{0})$ is then,
 $H^{(N)}(s,z_{0})$ $=P(\left\{z_{2},\cdots,z_{N}\right\}\notin\mathcal{D}(z_{0},s)|z_{1}=z_{0})$
@@
 $P(0,z_{2},\cdots,z_{N})=\frac{1}{N!\pi^{N}\prod_{j=0}^{N-1}j!}e^{-\sum_{j=2}^{N}|z_{j}|^{2}}\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}$
-Let $z_{1}=z_{0}=0$, the Vandermonde determinant
+Let $z_{1}=z_{0}=0$, the Vandermonde determinant is
 $\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{2}=\prod_{j=2}^{N}|z_{j}|^{2}\prod_{2\leq i<j\leq N}|z_{i}-z_{j}|^{2}$
@@
 $R_{1}^{(N)}(0)=\frac{1}{\pi}\frac{\Gamma(N,0)}{\Gamma(N)}=\frac{1}{\pi}$
-Applying the Andreief’s integration formula
+Applying Andreief’s integration formula
 $H^{(N)}(s,0)$ $=\frac{N(N-1)!}{N!\pi^{N-1}\prod_{j=0}^{N-1}j!}\det\left[\int_{|z|>s}\left(|z|^{2}\right)^{N+1-j}e^{-|z|^{2}}d^{2}z\right]_{j=2}^{N}$